### v1

# On the Linear Speedup of the Push-Pull Method for Decentralized Optimization over Digraphs

Abstract
The linear speedup property is essential for demonstrating the advantage of distributed algorithms over their single-node counterparts. In this paper, we study the stochastic Push-Pull method, a widely adopted decentralized optimization algorithm over directed graphs (digraphs). Unlike methods that rely solely on row-stochastic or column-stochastic mixing matrices, Push-Pull avoids nonlinear correction and has shown superior empirical performance across a variety of settings. However, its theoretical analysis remains challenging, and the linear speedup property has not been generally established—revealing a significant gap between empirical success and limited theoretical understanding. To bridge this gap, we propose a novel analysis framework and prove that Push-Pull achieves linear speedup over arbitrary strongly connected digraphs. Our results provide the comprehensive theoretical understanding for stochastic Push-Pull, aligning its theory with empirical performance. Code: https://github.com/pkumelon/PushPull.
1 Introduction
Scaling machine learning tasks to large datasets and models necessitates efficient distributed computation across multiple nodes. In distributed optimization, a network of $n$ nodes cooperates to solve
$\displaystyle\min_{x\in\mathbb{R}^{d}}\quad f(x):=\frac{1}{n}\sum_{i=1}^{n}f_{i}(x)\quad\mbox{where}\quad f_{i}(x)=\mathbb{E}_{\xi_{i}\sim\mathcal{D}_{i}}[F(x;\xi_{i})].$ (1)
Here, $\xi_{i}$ denotes a random data sample drawn from the local distribution $\mathcal{D}_{i}$, and $f_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is the local loss function, which may be nonconvex. During optimization, each node $i$ can access its own local gradient $\nabla F_{i}(x;\xi_{i})$, but must communicate to obtain gradient information from the other nodes. Data heterogeneity is typically present in (1), meaning the local distributions $\{\mathcal{D}_{i}\}_{i=1}^{n}$ differ across nodes. Consequently, the loss functions $f_{i}(x)$ and $f_{j}(x)$ are generally not equal, i.e., $f_{i}(x)\neq f_{j}(x)$.
Parallel stochastic gradient descent (SGD) is a widely adopted approach in distributed optimization, where local gradients are globally averaged using either a central server [1] or Ring All-Reduce communication primitives [2]. However, such global synchronization incurs substantial bandwidth overhead and high latency as the number of nodes increases, thereby hurting the scalability of the optimization process. To address these limitations, decentralized SGD has recently been proposed to solve problem (1) without requiring global gradient aggregation. In this setting, each node independently maintains and updates its model by communicating only with immediate neighbors, significantly reducing communication overhead. Beyond efficiency, decentralized algorithms offer improved robustness to node and link failures, maintaining convergence as long as the underlying communication graph remains connected. In contrast, centralized SGD breaks down if the central server crashes.
In this paper, we consider a general decentralized setting where the network topology is modeled as a strongly connected directed graph (digraph), which naturally encompasses undirected networks as a special case. Compared to undirected graphs, digraphs are particularly well-suited for ad hoc decentralized systems, such as heterogeneous robotic networks that operate under asynchronous, one-sided communication and random node activation. More recently, digraphs have shown great potential in distributed deep learning, where certain classes of directed networks exhibit both topological sparsity and strong algebraic connectivity. These properties make them especially attractive for the distributed training of large-scale models. Notable examples include the exponential graph [3], EquiTopo [4], and B-ary tree [5].
1.1 Decentralized algorithms over digraphs
In any connected network, the topology can be characterized by a mixing matrix. This matrix reflects how nodes are connected and is important for understanding how the network affects algorithm performance. For undirected networks, it is straightforward to construct symmetric and doubly stochastic mixing matrices. For directed networks, achieving a doubly stochastic matrix is generally infeasible. Consequently, mixing matrices in such cases are typically either column-stochastic [6, 7] or row-stochastic [8, 9], but not both.
Building on the connectivity pattern of the underlying network topology, decentralized optimization algorithms over directed graphs can be broadly categorized into three families in the literature. The first family relies on the Push-Sum technique [10, 11], which employs column-stochastic mixing matrices for all communications. Accordingly, we refer to this class as Push-Only algorithms in this paper. When exact gradients $\nabla f_{i}$ are available, the well-known Subgradient-Push algorithm [12, 11] converges to the desired solution, albeit with relatively slow sublinear rates even under strong convexity. More advanced methods such as EXTRA-Push [13], DEXTRA [14], ADD-OPT [15], and Push-DIGing [7] improve convergence speed by mitigating the effects of data heterogeneity. Stochastic variants, including SGD-Push [16] and S-ADDOPT [17], extend these approaches to settings with stochastic gradients.
The second family consists of algorithms that use only row-stochastic matrices for communication, referred to as the Pull-Only setting. Just as Push-Sum underpins the design of Col-Only algorithms, the Pull-Diag gossip protocol [9] serves as the foundation for Row-Only methods. Building on Pull-Diag, reference [9] adapted the distributed gradient descent algorithm to the Row-Only setting. Subsequently, [18, 19] extended gradient tracking techniques to this framework, while momentum-based Pull-Only gradient tracking methods were developed by [20, 21].
The third family comprises algorithms that alternate between row-stochastic and column-stochastic mixing matrices, commonly referred to as the Push-Pull method [22] or the AB method [23]. In this framework, a row-stochastic matrix is used to exchange model parameters, while a column-stochastic matrix is employed to share gradient information. As demonstrated in [23, 22, 24], the Push-Pull family typically achieves significantly faster convergence rates than algorithms relying solely on column- or row-stochastic matrices. [25] extended the Push-Pull method to stochastic optimization settings, while [5] improved its convergence rate on B-ary tree network topologies. Accelerated or constrained variant of the Push-Pull method have also been explored in [26, 27].
1.2 Linear speedup in iteration complexity
Linear speedup refers to the property of a distributed algorithm whereby the number of iterations required to reach an $\epsilon$-accurate solution decreases linearly with the number of nodes. For example, centralized SGD achieves an $\epsilon$-accurate solution in $\mathcal{O}\left(\textstyle\frac{1}{n\epsilon^{2}}\right)$ iterations , where $n$ denotes the number of nodes. This scaling behavior demonstrates that centralized SGD achieves linear speedup with respect to the number of nodes. The linear speedup property is crucial for establishing the advantage of distributed algorithms over their single-node counterparts. Without this property, the practical utility of a distributed algorithm may be limited, as it would offer little benefit from using multiple computing nodes.
The linear speedup property of decentralized algorithms over undirected graphs is well established. The first such analysis for decentralized SGD was presented in [28], demonstrating that decentralized SGD can achieve the same asymptotic convergence rate as its centralized counterpart. Building on this foundation, a variety of advanced decentralized algorithms—such as Exact Diffusion/NIDS [29, 30] and Gradient Tracking methods [31, 32, 33]—are proven to achieve linear speedup [34, 29, 35]. In addition, [36] provides a linear speedup analysis for decentralized SGD under time-varying undirected network topologies.
Recent literature has increasingly focused on establishing the linear speedup property of decentralized algorithms over digraphs. For Push-Only algorithms that employ column-stochastic mixing matrices, linear speedup has been formally proven in [16, 37]. Moreover, a recent study [38] proposes methods that achieve the optimal iteration complexity in this setting. For Pull-Only algorithms that employ row-stochastic mixing matrices, convergence has been extensively studied under various settings [9, 18, 19, 20, 21]; however, these analyses do not establish the linear speedup property. The main challenge lies in the fact that using a row-stochastic matrix alone introduces a deviation in the descent direction from the globally averaged gradient. With a novel analysis framework addressing this issue, the linear speedup property of Pull-Only algorithms was established only recently in [39]. These results reveal the effectiveness of Push-Only and Pull-Only algorithms over digraphs.
However, although Push-Pull algorithms typically achieve significantly better empirical performance than Push-Only and Pull-Only methods, their linear speedup property has not been generally established. Existing theoretical analyses of Push-Pull remain limited, primarily focusing on strongly convex loss functions in deterministic settings [23, 40, 22, 5]. In the nonconvex and stochastic setting, neither linear speedup nor even basic convergence guarantees have been formally established for Push-Pull. The most relevant theoretical result to date is presented in [5], which introduces a variant called B-ary Tree Push-Pull (BTPP). While BTPP is shown to achieve linear speedup, the result relies on the specialized B-ary tree topology, and the analysis does not generalize to general network topologies.
1.3 Motivation and main results
Building on the above discussion, the existing theoretical convergence analyses for Push-Pull methods remain rather limited, underscoring a clear discrepancy between the weak theoretical guarantees and their strong empirical performance. This gap gives rise to the following two open questions:
  • Q1.
    What is the intrinsic difference between the Push-Pull method and the other two families, i.e., Push-Only and Pull-Only? Why is the theoretical analysis of Push-Pull more challenging?
  • Q2.
    Can Push-Pull achieve linear speedup over a broader class of network topologies—for instance, all strongly connected digraphs?
Addressing these two questions is crucial for establishing a theoretical foundation that supports the effectiveness of Push-Pull methods.
Regarding Question Q1, we demonstrate that both Push-Only and Pull-Only algorithms closely align with centralized SGD; specifically, after a careful numerical transformation, their variable updates closely resemble those of centralized SGD. This structural similarity explains why the linear speedup property can be readily established for Push-Only and Pull-Only methods. In contrast, Push-Pull is not designed to align with centralized SGD and consistently exhibits a non-vanishing deviation from the centralized trajectory. Consequently, the linear speedup analysis for Push-Pull is substantially more challenging and fundamentally distinct from that of Push-Only and Pull-Only algorithms.
For Question Q2, we establish a linear speedup convergence result for the Push-Pull method. To the best of our knowledge, this is the first such result for Push-Pull over all strongly connected digraphs under standard assumptions, whereas existing analyses of stochastic Push-Pull [5, 41] are restricted to tree-structured topologies. By introducing a novel analysis framework that addresses the misalignment between Push-Pull and centralized SGD, this paper bridges the gap between the convergence theory and the strong empirical performance of the Push-Pull method.
The remainder of the paper is organized as follows. Section 2 introduces necessary notations, assumptions, and measures. Section 3 reviews the Push-Only, Pull-Only, and Push-Pull methods, and offers insights into the derivation of linear speedup convergence. In Section 4, we present the key components of our analysis and state the main convergence theorem. Section 5 provides extensive numerical experiments that demonstrate the ability of the Push-Pull method to achieve linear speedup. Finally, Section 6 concludes the paper.
2 Notations, Assumptions and Measures
This section introduces notations, assumptions, and measures that will be used throughout the paper.
Notations. The directed communication network is denotes as ${\mathcal{G}}=({\mathcal{N}},{\mathcal{E}})$, where ${\mathcal{N}}=\{1,2,\ldots,n\}$ is the set of nodes, and ${\mathcal{E}}$ is the set of directed edges. An edge $(j,i)\in{\mathcal{E}}$ means that node $i$ can send information to node $i$. We say a directed graph is strongly connected if there is a directed path from any node to all other nodes in the graph. We say that ${\mathcal{G}}=({\mathcal{N}},{\mathcal{E}})$ is the underlying topology of weight matrix $A$ if $A$ is entrywise nonnegative, and $a_{ij}>0$ if and only if $(j,i)\in{\mathcal{E}}$.
Let $\mathds{1}\in\mathbb{R}^{n}$ denote the vector of all-ones. The set $[n]$ denotes the indices $\{1,2,\ldots,n\}$. For a vector $v$, $\mathrm{Diag}(v)$ signifies the diagonal matrix whose diagonal elements are comprised of $v$.
We use matrix $A$ to denote the row-stochastic matrix ($A\mathds{1}=\mathds{1}$) and matrix $B$ to denote the column-stochastic matrix ($B^{\top}\mathds{1}=\mathds{1}$). The Perron vectors [42] for matrices $A$ and $B$, if they exist, are denoted as $\pi_{A}$ and $\pi_{B}$, respectively. We introduce $A_{\infty}:=\mathds{1}\pi_{A}^{\top}$ and $B_{\infty}:=\pi_{B}\mathds{1}^{\top}$. Additionally, we define $n\times d$ matrices
$\displaystyle{\mathbf{x}}^{(k)}$ $\displaystyle:=[({x}_{1}^{(k)})^{\top};({x}_{2}^{(k)})^{\top};\cdots;({x}_{n}^{(k)})^{\top}]$
$\displaystyle{\mathbf{g}}^{(k)}$ $\displaystyle:=[\nabla F_{1}({x}_{1}^{(k)};\xi_{1}^{(k)})^{\top};\cdots;\nabla F_{n}({x}_{n}^{(k)};\xi_{n}^{(k)})^{\top}]$
$\displaystyle\nabla f({\mathbf{x}}^{(k)})$ $\displaystyle:=[\nabla f_{1}({x}^{(k)}_{1})^{\top};\cdots;\nabla f_{n}({x}^{(k)}_{n})^{\top}]$
[…]
Assumptions. We first specify the underlying graphs (${\mathcal{G}}_{A}$, ${\mathcal{G}}_{B}$) and weight matrices ($A$, $B$) used in stochastic Push-Pull.
Assumption 1 (Communication Graph).
Graphs ${\mathcal{G}}_{A}$ and ${\mathcal{G}}_{B}$ are strongly connected.
Assumption 2 (Weight Matrix).
Weight matrix $A$ is row-stochastic, with an underlying topology of ${\mathcal{G}}_{A}$. Weight matrix $B$ is column-stochastic, with an underlying topology of ${\mathcal{G}}_{B}$. The traces of $A$ and $B$ are positive.
Assumptions 1 and 2 are commonly used in the analysis of Push-Pull [24, 40, 26]. In our paper, the primary role of these assumptions is to derive Proposition 1. Strictly speaking, replacing these assumptions while retaining Proposition 1 would not affect our main convergence results.
Proposition 1 (Perron Vector, Exponential Decay).
For $A$ and $B$ satisfying Assumption 1 and 2, there exist entrywise positive Perron vectors $\pi_{A},\pi_{B}\in\mathbb{R}^{n}$ such that
$\pi_{A}^{\top}A=\pi_{A}^{\top},\quad B\pi_{B}=\pi_{B},\quad\|\pi_{A}\|_{1}=\|\pi_{B}\|_{1}=1,\quad\pi_{A}^{\top}\pi_{B}>0.$
Besides, there exists constants $p_{A},p_{B}>0$, $\lambda_{A},\lambda_{B}\in[0,1)$ satisfying
$\|A^{k}-A_{\infty}\|_{2}\leq p_{A}\lambda_{A}^{k},\quad\|B^{k}-B_{\infty}\|_{2}\leq p_{B}\lambda_{B}^{k},\quad\forall k\geq 0.$
Proof.
The proof follows directly from the Perron-Frobenius theorem [42]. ∎
Finally, we state standard assumptions regarding the stochastic gradients and function smoothness.
Assumption 3 (Gradient Oracle).
There exists a constant $\sigma\geq 0$ such that
$\displaystyle\mathbb{E}_{k}[\nabla F_{i}(x_{i}^{(k)};\xi_{i}^{(k)})]$ $\displaystyle=\nabla f_{i}(x_{i}^{(k)}),\quad\mathbb{E}_{k}[\|\nabla F_{i}(x_{i}^{(k)};\xi_{i}^{(k)})-\nabla f_{i}(x_{i}^{(k)})\|_{F}^{2}]\leq\sigma^{2},\quad\forall i\in[n],\,k\geq 0.$
Assumption 4 (smoothness).
There exist positive constants $L,\Delta$ such that
$\displaystyle\|\nabla f_{i}(x)-\nabla f_{i}(y)\|\leq L\|x-y\|\quad\mbox{and}\quad f_{i}(x)-\inf_{x}f_{i}(x)\leq\Delta,\quad\forall i\in[n].$
Measures. When Proposition 1 holds, we can define the following constants
$\displaystyle s_{A}:=\sum_{i=0}^{\infty}\|A^{i}-A_{\infty}\|_{2},\quad s_{B}:=\sum_{i=0}^{\infty}\|B^{i}-B_{\infty}\|_{2},\quad c:=n\pi_{A}^{\top}\pi_{B}.$ (2)
In this paper, we use only $s_{A}$, $s_{B}$, $c$, and $n$ to characterize the influence of the mixing matrices. Existing analyses on directed graphs typically employ measures related to the induced matrix norm [25, 38, 39]. In contrast, our measures only require the $\ell_{2}$ norm. An interesting relationship between our measures and other measures is presented in Proposition 2.
Proposition 2 (Relation with other measures).
Suppose Assumption 1 and 2 hold. We have the following estimate on $s_{A}$, $s_{B}$ and $c$:
$\displaystyle 1\leq s_{A}\leq\frac{1+\frac{1}{2}\ln(\kappa_{A})}{1-\beta_{A}},$ $\displaystyle\quad 1\leq s_{B}\leq\frac{1+\frac{1}{2}\ln(\kappa_{B})}{1-\beta_{B}},$
$\displaystyle n\max\{\min\pi_{A},\min\pi_{B}\}$ $\displaystyle\leq\,c\leq n\min\{\max\pi_{A},\max\pi_{B}\},$
where $\beta_{A}:=\|\Pi_{A}^{\frac{1}{2}}(A-A_{\infty})\Pi_{A}^{-\frac{1}{2}}\|_{2}\in[0,1)$, $\beta_{B}:=\|\Pi_{B}^{-\frac{1}{2}}(B-B_{\infty})\Pi_{B}^{\frac{1}{2}}\|_{2}\in[0,1),$ $\kappa_{A}:=\max\pi_{A}/\min\pi_{A}$, $\kappa_{B}:=\max\pi_{B}/\min\pi_{B}$ are measures from literature [25, 38, 39], $\Pi_{A}=\mathrm{Diag}(\pi_{A})$ and $\Pi_{B}=\mathrm{Diag}(\pi_{B})$.
Proof.
See Appendix A. ∎
3 Preliminary
In this section, we provide a concise overview of the linear speedup analysis for the Push-Only and Pull-Only algorithms. Furthermore, we discuss the limitations of existing analytical approaches and explain why they cannot be directly extended to the Push-Pull algorithm.
Push-Only and Pull-Only Algorithms Align with Centralized SGD. Most decentralized algorithms, such as decentralized SGD [36], decentralized gradient tracking [43, 14], (sub)gradient-Push [44], Push-Diging [7], (sub)gradient-Pull [9], and FROST [19], are designed to mimic centralized SGD. These algorithms can generally be formulated as:
$\displaystyle\hat{x}^{(k+1)}=\hat{x}^{(k)}-\alpha\bar{g}^{(k)}+\,\epsilon^{(k)}$ (3)
where $\hat{x}^{(k)}=\sum_{i=1}^{n}v_{i}^{(k)}x_{i}^{(k)}$ is a convex combination of local decision variables, and $\bar{g}^{(k)}=\frac{1}{n}\sum_{i=1}^{n}\nabla F_{i}(x_{i}^{(k)},\xi_{i}^{(k)})$ represents the average gradient. Here, $\epsilon^{(k)}$ denotes an additional error term.
For Push-Only and Pull-Only optimization algorithms, the extra error term is elaborately designed to be zero or to approach zero exponentially fast. Such design ensures that the trajectories of decision variables closely resemble those of centralized SGD. Given the sufficiently small extra error, the linear speedup result naturally follows within the same analytical framework used for centralized SGD. The following is a brief example of how to obtain the formulation (3).
Example 1.
Push-Diging [7]. The algorithm can be written as:
$\displaystyle{\mathbf{x}}^{(k+1)}$ $\displaystyle=A_{k}{\mathbf{x}}^{(k)}-\alpha V_{k+1}^{-1}{\mathbf{y}}^{(k)},$ (4a)
$\displaystyle{\mathbf{y}}^{(k+1)}$ $\displaystyle=B{\mathbf{y}}^{(k)}+{\mathbf{g}}^{(k+1)}-{\mathbf{g}}^{(k)},$ (4b)
where $V_{k}=\mathrm{Diag}(v_{k})=\mathrm{Diag}(B^{k}\mathds{1})$ is a diagonal matrix and $A_{k}=V_{k+1}^{-1}BV_{k}$. By left-multiplying $n^{-1}v_{k+1}^{\top}$ on both sides of (4a) and omitting some calculation details, we will obtain
$n^{-1}v_{k+1}^{\top}{\mathbf{x}}^{(k+1)}=n^{-1}v_{k}^{\top}{\mathbf{x}}^{(k)}-\alpha\bar{g}^{(k)}.$
By taking $\hat{x}^{(k)}=n^{-1}v_{k}^{\top}{\mathbf{x}}^{(k+1)}$, this is exactly the form of (3).
Figure 1: Distributed logistic regression on synthetic dataset. We use stochastic gradients and a constant learning rate 0.0050.005 0.0050.005. The top three plots illustrate the convergence behavior across different network topologies, while the bottom three plots demonstrate that the size of extra error term rapidly stabilizes and persists as a significant constant. For detailed settings, see 36, Appendix C.
Push-Pull Analysis Suffers from Extra Error Terms. Unlike the aforementioned algorithms, it is generally impossible to reformulate Push-Pull to obtain a centralized SGD-like update. The recursion of Push-Pull is
$\displaystyle{\mathbf{x}}^{(k+1)}$ $\displaystyle=A{\mathbf{x}}^{(k)}-\alpha{\mathbf{y}}^{(k)},$ (5a)
$\displaystyle{\mathbf{y}}^{(k+1)}$ $\displaystyle=B{\mathbf{y}}^{(k)}+{\mathbf{g}}^{(k+1)}-{\mathbf{g}}^{(k)},$ (5b)
[…]
$\hat{x}^{(k+1)}=\hat{x}^{(k)}-c\alpha\bar{g}^{(k)}-\alpha\pi_{A}^{\top}(I-B_{\infty}){\mathbf{y}}^{(k)},$
where $\hat{x}^{(k)}=\pi_{A}^{\top}{\mathbf{x}}^{(k)}$ is the projected decision variable, $c=n\pi_{A}^{\top}\pi_{B}$ is a constant. We observe that the error term $\pi_{A}^{\top}(I-B_{\infty}){\mathbf{y}}^{(k)}$ does not vanish but remains a large constant, see Figure 1.
This non-vanishing error term makes achieving linear speedup in Push-Pull highly nontrivial, because in traditional linear speedup analysis, only the noise from $\bar{g}^{(k)}$ is the dominating factor. But here, the error term $\|\pi_{A}^{\top}(I-B_{\infty}){\mathbf{y}}^{(k)}\|$ introduces additional noise that may dominate the noise in $\bar{g}^{(k)}$.
For certain specific topologies, the noise in the error term can either be ignored or absorbed. For instance, the work in [5] explores a special case known as the $B$-ary tree Push-Pull and demonstrates the linear speedup result on this particular topology. In their setting, $\pi_{A}=\pi_{B}=[1,0,\ldots,0]^{\top}$, which leads to the expression $\pi_{A}^{\top}(I-B_{\infty})=(0,-1,-1,\ldots,-1)$. By setting the loss function at the first node to zero, they obtain the relationship $\pi_{A}^{\top}(I-B_{\infty}){\mathbf{y}}^{(k)}=-n\bar{g}^{(k)}$, which simplifies the process of achieving linear speedup.
Unfortunately, for general topologies, the error term lacks exploitable structure, making direct analysis overly conservative. This leads to a significant overestimation of the noise term, which in turn obstructs the derivation of a linear speedup result. Consequently, the central challenge becomes: How can we accurately characterize the noise within this error term? We tackle this key question in the following section.
4 Achieving Linear Speedup for Push-Pull
For simplicity, we denote $\Delta_{y}^{(k)}:=(I-B_{\infty}){\mathbf{y}}^{(k)}$. Using the gradient tracking recursion (5b), we discover the following relationship about extra error terms: for any integer $k\geq 0,m\geq 1$:
$\displaystyle\sum_{i=1}^{m}\Delta_{y}^{(k+i)}=(B-B_{\infty})\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}+(I-B_{\infty})({\mathbf{g}}^{(k+m)}-{\mathbf{g}}^{(k)}).$ (6)
Equation (6) reveals a contraction-type recursion governing the sum of consecutive error terms. Informally, if we restrict our analysis to terms involving $\sigma^{2}$, we observe that $\|{\mathbf{g}}^{(k+m)}-{\mathbf{g}}^{(k)}\|F^{2}={\mathcal{O}}(\sigma^{2})$. This contraction property further implies that the aggregated term $\|\sum{i=1}^{m}\Delta_{y}^{(k+i)}\|_{F}^{2}$ maintains the same ${\mathcal{O}}(\sigma^{2})$ scaling. Consequently, we can deduce that individual terms satisfy $\|\Delta_{y}^{(k)}\|_{F}^{2}={\mathcal{O}}(\frac{\sigma^{2}}{m})$.
This scaling relationship becomes particularly significant when $m$ is sufficiently large: in this regime, the $\sigma^{2}$-dependent component of the extra error term becomes negligible. As a result, our analysis preserves the crucial linear speedup property. To formalize this logic, we start from the following contraction relationship.
Lemma 3.
For any integers $k\geq 0$ and $m\geq 1$, we have
$\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}=\pi_{A}^{\top}\left(\sum_{j=0}^{m-1}(B^{j}-B_{\infty})\right){\mathbf{y}}^{(k)}+\pi_{A}^{\top}\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})({\mathbf{g}}^{(k+j)}-{\mathbf{g}}^{(k)}).$ (7)
Furthermore, we can define a ”noiseless” version for the right-hand side, i.e.
$H_{m,k}:=\pi_{A}^{\top}\left(\sum_{j=0}^{m-1}(B^{j}-B_{\infty})\right){\mathbf{y}}^{(k)}+\pi_{A}^{\top}\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)})),$
and the distance therein can be bounded by:
$\mathbb{E}_{k}[\|\sum_{i=0}^{m-1}\varepsilon^{(k+i)}-H_{m,k}\|_{F}^{2}]\leq 2ns_{B}^{2}\sigma^{2}.$ (8)
The proof of Lemma 3 is provided in Appendix B.1. This lemma implies that the summation of the $m$ error terms can be represented as $H_{m,k}$ plus a bounded noise term. Moreover, the term $H_{m,k}$ exclusively involves difference of deterministic gradients and a single instance of ${\mathbf{y}}^{(k)}$, making it straightforward to estimate its magnitude.
To effectively leverage the result of Lemma 3, we introduce an $m$-step descent lemma. This new lemma follows by an estimation on $H_{m,k}$ and applying the standard $L$-smoothness descent lemma to the following recursion:
$\hat{x}^{(k+m)}=\hat{x}^{(k)}-c\alpha\sum_{i=0}^{m-1}\bar{g}^{(k+i)}-\alpha\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}.$ (9)
Lemma 4 (m-step Descent Lemma).
Suppose Assumption 34 and Proposition 1 hold. For any integers $k\geq 0$, we define $\hat{x}^{(k)}=\pi_{A}^{\top}{\mathbf{x}}^{(k)}$. The following inequality holds for any integers $k\geq 0,m\geq 1$:
$\displaystyle\quad\frac{\alpha cm}{4}\|\nabla f(\hat{x}^{(k)})\|^{2}+\frac{\alpha c(1-4\alpha mcL)}{2}\sum_{i=0}^{m-1}\mathbb{E}_{k}[\|\overline{\nabla f}^{(k+i)}\|^{2}]$ (10)
$\displaystyle\leq f(\hat{x}^{(k)})-\mathbb{E}_{k}[f(\hat{x}^{(k+m)})]+\frac{2\alpha^{2}c^{2}mL}{n}\sigma^{2}+2\alpha^{2}Lns_{B}^{2}\sigma^{2}$
$\displaystyle\quad+\frac{L}{16m}\Delta_{1}^{(k)}+\frac{L}{16m}\Delta_{2}^{(k)}+{3\alpha^{2}Ls_{B}^{2}}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}],$
where
$\displaystyle\Delta_{1}^{(k)}=\sum_{i=0}^{m-1}\mathbb{E}_{k}[\|{\mathbf{x}}^{(k+i)}-\hat{{\mathbf{x}}}^{(k)}\|_{F}^{2}],\quad\Delta_{2}^{(k)}=\sum_{j=1}^{m-1}\mathbb{E}_{k}[\|{\mathbf{x}}^{(k+j)}-{\mathbf{x}}^{(k)}\|_{F}^{2}].$
Remarkably, if we divide both sides of (10) by $\frac{c\alpha m}{4}$, the first noise term simplifies to $\frac{8\alpha cL}{n}\sigma^{2}$, which corresponds to the standard linear speedup term. The second noise term simplifies to $\frac{8\alpha ns_{B}^{2}}{cm}\sigma^{2}$. When we select some $m$ larger than $c^{-2}n^{2}s_{B}^{2}$, the second noise term should have a smaller order than the first term, thereby ensuring the achievement of the standard linear speedup, which verifies our logic before. Next, we try to estimate $\Delta_{1}^{(k)}$ and $\Delta_{2}^{(k)}$.
Lemma 5 (Consensus Error).
Suppose Assumption 34 and Proposition 1 hold. With $\Delta_{1}^{(k)}$ and $\Delta_{2}^{(k)}$ defined above, when $\alpha\leq\min\{\frac{3c}{16mL},\frac{1}{6s_{B}L}\}$, we have
$\displaystyle\quad\frac{L}{16m}\mathbb{E}_{k}[\Delta_{1}^{(k)}+\Delta_{2}^{(k)}]\leq\frac{s_{A}^{2}L}{m}\|\Delta_{x}^{(k)}\|_{F}^{2}+\alpha^{2}Ls_{B}^{2}\|\Delta_{y}^{(k)}\|_{F}^{2}+2\alpha^{2}Lns_{B}^{2}\sigma^{2}.$
where $\Delta_{x}:={\mathbf{x}}^{(k)}-A_{\infty}{\mathbf{x}}^{(k)}$, $\Delta_{y}^{(k)}:={\mathbf{y}}^{(k)}-B_{\infty}{\mathbf{y}}^{(k)}$.
Lemma 5 transfers the complicated error terms to commonly-seen consensus error terms. Now, we can apply techniques similar to [38] to bound $\Delta_{x}$ and $\Delta_{y}$.
Lemma 6.
Suppose Assumption 34 and Proposition 1 hold. For any $T\geq 1$, when $\alpha\leq\frac{1}{6\sqrt{n}s_{A}s_{B}L}$, we have
$\displaystyle\sum_{k=0}^{T}\mathbb{E}[\|\Delta_{x}^{(k+1)}\|^{2}]$ $\displaystyle\leq 2\alpha^{2}s_{A}^{2}\sum_{k=0}^{T}\mathbb{E}[\|\Delta_{y}^{(k)}\|^{2}]+2\alpha^{2}n^{2}s_{A}^{2}\sum_{k=0}^{T}\mathbb{E}[\|\overline{\nabla f}^{(k)}\|^{2}]+2n\alpha^{2}s_{A}^{2}(T+1)\sigma^{2},$
and
$\displaystyle\sum_{k=0}^{T}\|\Delta_{y}^{(k+1)}\|_{F}^{2}\leq 30ns_{B}^{2}(T+1)\sigma^{2}+8ns_{B}^{2}L\Delta+108\alpha^{2}L^{2}n^{3}s_{A}^{2}s_{B}^{2}\sum_{k=0}^{T}\mathbb{E}[\|\overline{\nabla f}^{(k)}\|_{F}^{2}].$
Now we are ready to state the main result of this paper.
Theorem 1.
Suppose Assumption 34 and Proposition 1 hold. For given positive integers $m$ and $K$, we define $[m^{*}]:=\{0,m,\ldots,(K-1)m\}$ and $T:=mK$. Let $m=\lceil 20n^{2}c^{-2}s_{B}^{2}(s_{A}^{4}+s_{B}^{2})\rceil$. When the learning rate $\alpha$ satisfies $\alpha\leq\min\{\frac{1}{8cL},\frac{c}{21n^{2}s_{B}^{2}(s_{A}^{4}+s_{B}^{2})L}\}$, we have
$\displaystyle\quad\frac{1}{K}\sum_{k\in[m^{*}]}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]\leq\frac{4\Delta}{c\alpha T}+\frac{30c\alpha L}{n}\sigma^{2}+\frac{64n\alpha L^{2}(s_{A}^{4}+s_{B}^{2})\Delta}{cmT}.$
where $C_{1}$ is a constant determined by $A$ and $B$. Furthermore, if we choose
$\alpha=\min\left\{\frac{c}{21n^{2}s_{B}^{2}(s_{A}^{4}+s_{B}^{2})L},\frac{1}{8cL},\frac{\sqrt{n\Delta}\sigma}{3c\sqrt{LT}}\right\},$
we will obtain that
$\displaystyle\min_{t\in\{0,1,\ldots,T-1\}}\mathbb{E}[\|\nabla f(\hat{x}^{(t)}\|^{2}]\leq\frac{1}{K}\sum_{k\in[m^{*}]}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]\leq\frac{20\sigma\sqrt{L\Delta}}{\sqrt{nT}}+\frac{C_{\Delta}L\Delta}{T},$ (11)
where
$C_{\Delta}=32+32ncs_{A}+84n^{2}c^{-2}s_{B}^{2}(s_{A}^{4}+s_{B}^{2})$
is a constant decided by $A$ and $B$.
Comparison with existing analyses. Theorem 1 provides the first finite-time non-convex convergence result for Push-Pull, while previous analyses focus either on the strongly convex case [45, 40] or on variants of Push-Pull relying on specific topologies [5]. The first term $\frac{20\sigma\sqrt{L\Delta}}{\sqrt{nT}}$ explicitly shows that Push-Pull enjoys a linear speedup on arbitrary strongly connected topologies.
Impact of $m$. In our analysis, we consider a $m$-step gradient descent. The number $m$ effectively controls the $\sigma^{2}$-related term in the extra error. A common question is can we choose an arbitrarily large $m$ in this analysis. We can do that, but choosing a larger $m$ will force us to select a smaller learning rate, which finally leads to a larger transient time. To sum up, number $m$ provides a chance for tradeoff between the noise level and the transient time.
Transient time. Our theorem provides a coarse estimate for the transient time, which is of ${\mathcal{O}}(nC_{\Delta}^{2})$. We do not recommend using this to estimate the real transient time, but it still reflects the impact of topology. For example, if we use a directed ring graph where $s_{A}=s_{B}={\mathcal{O}}(n)$, $c={\mathcal{O}}(1)$, the transient time can be ${\mathcal{O}}(n^{17})$. If we use a directed exponential graph [3] where $s_{A}=s_{B}={\mathcal{O}}(1)$, $c={\mathcal{O}}(1)$, the transient time will be ${\mathcal{O}}(n^{5})$.
5 Experiment
In this section, we empirically investigate the linear speedup properties of the Push-Pull algorithm across diverse network topologies. For the stochastic gradient oracle, we ensure that each node operates on a local dataset, computing gradients from randomly sampled data points at each iteration. The experimental validation proceeds in two phases: first, we examine a nonconvex regression problem using synthetic data, followed by training a fully connected neural network on the heterogeneously distributed MNIST dataset.
To further assess practical performance on established benchmarks, we extend our evaluation to ResNet-18 training on CIFAR-10. The distributed training experiments are conducted across multiple nodes ($n>1$) using three distinct network topologies: directed exponential graphs, grid graphs, and undirected random graphs. All results are systematically compared against the single-node baseline ($n=1$) to quantify the speedup achieved through distributed optimization.
Complete implementation specifications are provided in Appendix C. The source code and experimental scripts are publicly available in our GitHub repository.
5.1 Nonconvex Logistic Regression on Synthetic Dataset
In the first experiment, the loss function at each node is a logistic regression function with nonconvex regularization terms (see (36) in Appendix C). We studied six different network topologies: the exponential, grid, and ring graphs are directed, while the random, geometric, and nearest neighbor graphs are undirected. The mixing matrices $A$ and $B$ are generated by a special procedure (see Appendix C.1), which produces row-stochastic or column-stochastic matrices, but not doubly-stochastic ones.
We can observe clear linear speedup curves in Figure 2. For each topology, the noise level in gradients decreases proportionally to the square root of the number of nodes after the same number of communication rounds. All learning rates are set to $lr=0.005$.
Figure 2: Linear speedup performance over various topologies on synthetic data, the learning rates are all 0.0050.005 0.0050.005.
5.2 Training Neural Network for MNIST Classification
In the second experiment, we trained a three-layer fully connected neural network for MNIST classification. We used the same topologies as before: directed exponential, grid, and ring graphs, and undirected random, geometric, and nearest neighbor graphs. Data is distributed heterogeneously using a Directlet distribution, see Appendix C.3. Figure 3 shows that Push-Pull is able to achieve linear speedup on both directed and undirected topologies.
Figure 3: Linear speedup performance over various topologies, the learning rates are the same over the same topology.
5.3 Training Neural Network for CIFAR10 Image Classification
In the third experiment, we trained Resnet18 for CIFAR10 classification across different numbers of nodes and various topologies. We used three topologies mentioned before: directed exponential, grid graphs, and undirected random graphs. Here $n=1$ means training a single Resnet18 on a single-node without distributed optimization.
Figure 4 demonstrates the performance of our Push-Pull method. We observe that Push-Pull’s accuracy is initially lower compared to single-node optimization, which can be attributed to the overhead of the influence of topology. Nevertheless, after a sufficient number of epochs, Push-Pull achieves superior accuracy, leveraging its inherent speed advantages.
Figure 4: The test accuracy across different numbers of nodes and various topologies. All experiments were conducted using a learning rate of 0.0050.005 0.0050.005. The lower three figures specifically depict the accuracy after 100100 100100 epochs.
6 Conclusion
In this paper, we introduce a novel analysis technique and demonstrate linear speedup convergence for the Push-Pull method across arbitrary strongly connected topologies. This discovery enhances the community’s comprehension of the Push-Pull method and validates its effectiveness. Numerical experiments have corroborated our theoretical findings.
Appendix A Rolling-Sum Lemma
Lemma 7 (Rolling Sum Lemma).
For a row-stochastic weight matrix satisfying Proposition 1, the following inequality holds for any $T\geq 0$:
$\displaystyle\sum_{k=0}^{T}\|\sum_{i=0}^{k}(A^{k-i}-A_{\infty})\Delta^{(i)}\|_{F}^{2}\leq s_{A}^{2}\sum_{i=0}^{T}\|\Delta^{(i)}\|_{F}^{2},$ (12)
where $\Delta^{(i)}\in\mathbb{R}^{n\times d}$ are arbitrary matrices, $s_{A}$ is defined as in (2). Furthermore, if Assumption 1 and 2 hold, we have
$\displaystyle s_{A}:=\sum_{i=0}^{\infty}\|A^{i}-A_{\infty}\|_{2}\leq\frac{1+\frac{1}{2}\ln(\kappa_{A})}{1-\beta_{A}}.$
where $\beta_{A}$, $\kappa_{A}$ are defined in Proposition 2. The row-stochastic matrix $A$ and the subscript $A$ can be replaced by a column-stochastic matrix $B$ without changing the lemma.
Proof.
First, we prove that
$\displaystyle\|A^{i}-A_{\infty}\|_{2}\leq\sqrt{\kappa_{A}}\beta_{A}^{i},\forall i\geq 0.$ (13)
Notice that $\beta_{A}:=\|A-A_{\infty}\|_{\pi_{A}}$ and
$\|A^{i}-A_{\infty}\|_{\pi_{A}}=\|(A-A_{\infty})^{i}\|_{\pi_{A}}\leq\|A-A_{\infty}\|^{i}_{\pi_{A}}=\beta_{A}^{i},$
we have
$\|(A^{i}-A_{\infty})v\|=\|\Pi_{A}^{-1/2}(A^{i}-A_{\infty})v\|_{\pi_{A}}\leq\sqrt{\underline{\pi_{A}}}\beta_{A}^{i}\|v\|_{\pi_{A}}\leq\sqrt{\kappa_{A}}\beta_{A}^{i}\|v\|,$
where $\underline{\pi_{A}}:=\min\pi_{A}$, $\overline{\pi_{A}}:=\max\pi_{A}$. The last inequality comes from $\|v\|_{\pi_{A}}\leq\overline{\pi_{A}}\|v\|$. Therefore, (13) holds.
Second, we want to prove that
$\displaystyle s_{A}\leq\frac{1+\frac{1}{2}\ln(\kappa_{A})}{1-\beta_{A}}.$ (14)
Towards this end, we define $M_{A}:=\max\{1,\max_{k\geq 0}\|A^{k}-A_{\infty}\|_{2}\}$. According to (13), $M_{A}$ is well-defined. We also define $p=\max\left\{\frac{\ln(\sqrt{\kappa_{A}})-\ln(M_{A})}{-\ln(\beta_{A})},0\right\}$, then we can verify that $\|A^{i}-A_{\infty}\|_{2}\leq\min\{M_{A},M_{A}\beta_{A}^{i-p}\},\forall i\geq 0.$ With this inequality, we can bound $\sum_{i=0}^{k}\|A^{k+1-i}-A_{\infty}\|_{2}$ as follows:
$\displaystyle\sum_{i=0}^{k}\|A^{k-i}-A_{\infty}\|_{2}=\sum_{i=0}^{\min\{\lfloor p\rfloor,k\}}\|A^{i}-A_{\infty}\|_{2}+\sum_{i=\min\{\lfloor p\rfloor,k\}+1}^{k}\|A^{i}-A_{\infty}\|_{2}$ (15)
[…]
If $p=0$, (15) is simplified to $\sum_{i=0}^{k}\|A^{k-i}-A_{\infty}\|_{2}\leq M_{A}\cdot\frac{1}{1-\beta_{A}}$ and (14) is naturally satisfied. If $p>0$, let $x=\min\{\lfloor p\rfloor,k\}+1-p\in[0,1)$, (14) is simplified to
$\sum_{i=0}^{k}\|A^{k-i}-A_{\infty}\|_{2}\leq M_{A}(x+p+\frac{\beta_{A}^{x}}{1-\beta_{A}})\leq M_{A}(p+\frac{1}{1-\beta_{A}}).$
By the definition of $p$, we have $p\leq\frac{\frac{1}{2}\ln(\kappa_{A})}{1-\beta_{A}}$. This finishes the proof of (14).
Finally, to obtain (12), we use Jensen’s inequality. For any positive numbers $a_{0},a_{1},\ldots,a_{k}$ satisfying $\sum_{i=0}^{k}a_{i}=1$, we have
$\displaystyle\|\sum_{i=0}^{k}(A^{k-i}-A_{\infty})\Delta^{(i)}\|_{F}^{2}=\|\sum_{i=0}^{k}a_{k-i}\cdot a_{k-i}^{-1}(A^{k-i}-A_{\infty})\Delta^{(i)}\|_{F}^{2}$ (16)
[…]
By choosing $a_{k-i}=(\sum_{i=0}^{k}\|A^{k-i}-A_{\infty}\|_{2})^{-1}\|A^{k-i}-A_{\infty}\|_{2}$ in (16), we obtain that
$\displaystyle\|\sum_{i=0}^{k}(A^{k-i}-A_{\infty})\Delta^{(i)}\|_{F}^{2}\leq\sum_{i=0}^{k}\|A^{k-i}-A_{\infty}\|_{2}\cdot\sum_{i=0}^{k}\|A^{k-i}-A_{\infty}\|_{2}\|\Delta^{(i)}\|_{F}^{2}.$ (17)
By summing up (17) from $k=0$ to $T$, we obtain that
$\displaystyle\sum_{k=0}^{T}\|\sum_{i=0}^{k}(A^{k-i}-A_{\infty})\Delta^{(i)}\|_{F}^{2}\leq s_{A}\sum_{k=0}^{T}\sum_{i=0}^{k}\|A^{k-i}-A_{\infty}\|_{2}\|\Delta^{(i)}\|_{F}^{2}$
[…]
which finishes the proof of this lemma. ∎
Appendix B Proofs of Main Theorem
B.1 Proof of Lemma 3
For $1\displaystyle 1\leq i\leq m-$ 1, consider
$\displaystyle(I-B_{\infty}){\mathbf{y}}^{(k+i)}$ $\displaystyle=(I-B_{\infty})(B{\mathbf{y}}^{(k+i-1)}+{\mathbf{g}}^{(k+i)}-{\mathbf{g}}^{(k+i-1)})$
$\displaystyle=(B-B_{\infty}){\mathbf{y}}^{(k+i-1)}+(I-B_{\infty}){\mathbf{g}}^{(k+i)}-(I-B_{\infty}){\mathbf{g}}^{(k+i-1)}$
Sum it up for $1\displaystyle 1\leq i\leq m-$ 1, we obtain that
$(I-B_{\infty})\sum_{i=1}^{m-1}{\mathbf{y}}^{(k+i)}=(B-B_{\infty})\sum_{i=0}^{m-2}{\mathbf{y}}^{(k+i)}+(I-B_{\infty}){\mathbf{g}}^{(k+m-1)}-(I-B_{\infty}){\mathbf{g}}^{(k)}.$
Similarly, consider $(B-B_{\infty}){\mathbf{y}}^{(k+i)}$, $2\displaystyle 1\leq i\leq m-$ 2, we have
$(B-B_{\infty})\sum_{i=1}^{m-2}{\mathbf{y}}^{(k+i)}=(B^{2}-B_{\infty})\sum_{i=0}^{m-3}{\mathbf{y}}^{(k+i)}+(B-B_{\infty}){\mathbf{g}}^{(k+m-2)}-(B-B_{\infty}){\mathbf{g}}^{(k)}.$
Repeat this procedure for $m-3,m-4,\ldots,1$ and sum them up, we finally arrive at
$\sum_{i=0}^{m-1}(I-B_{\infty}){\mathbf{y}}^{(k+i)}=\sum_{i=0}^{m-1}(B^{i}-B_{\infty}){\mathbf{y}}^{(k)}+\sum_{i=1}^{m-1}(B^{m-i-1}-B_{\infty})({\mathbf{g}}^{(k+i)}-{\mathbf{g}}^{(k)}),$
which is just the first statement. Next, note that the difference between $\sum_{i=0}^{m-1}\varepsilon^{(k+i)}$ and $H_{m,k}$ equals to:
$\pi_{A}^{\top}\sum_{i=1}^{m-1}(B^{m-i-1}-B_{\infty})({\mathbf{g}}^{(k+i)}-\nabla f({\mathbf{x}}^{(k+i)}))-\pi_{A}^{\top}\sum_{i=1}^{m-1}(B^{m-i-1}-B_{\infty})({\mathbf{g}}^{(k)}-\nabla f({\mathbf{x}}^{(k)})).$
By taking the squared norm, we obtain that:
$\displaystyle\quad\sum_{i=1}^{m-1}\|\pi_{A}^{\top}(B^{m-i-1}-B_{\infty})({\mathbf{g}}^{(k+i)}-\nabla f({\mathbf{x}}^{(k+i)}))\|^{2}$
$\displaystyle+\|\pi_{A}^{\top}(\sum_{i=1}^{m-1}(B^{m-i-1}-B_{\infty}))({\mathbf{g}}^{(k)}-\nabla f({\mathbf{x}}^{(k)}))\|^{2}+\text{Inner Product Terms}.$
When we take the expectation conditioned on ${\mathcal{F}}_{k}$, all the inner product terms become zero. For example, for any $1\displaystyle 0\leq i<j\leq m-$ 1,
$\displaystyle\mathbb{E}_{k}\left[\left\langle(B^{m-i-1}-B_{\infty})({\mathbf{g}}^{(k+i)}-\nabla f({\mathbf{x}}^{(k+i)})),(B^{m-j-1}-B_{\infty})({\mathbf{g}}^{(k+j)}-\nabla f({\mathbf{x}}^{(k+j)}))\right\rangle\right]$
$\displaystyle=\mathbb{E}_{k}[\mathbb{E}_{k+j}[\text{This inner product}]]$
$\displaystyle=\mathbb{E}_{k}\left[\left\langle(B^{m-i-1}-B_{\infty})({\mathbf{g}}^{(k+i)}-\nabla f({\mathbf{x}}^{(k+i)})),0\right\rangle\right]=\mathbb{E}[0]=0.$
Therefore, we only need to estimate the squared terms. By the Cauchy-Schwarz inequality, and $\|\pi_{A}\|\leq 1$, we have
$\displaystyle\quad\sum_{i=1}^{m-1}\mathbb{E}_{k}[\|\pi_{A}^{\top}(B^{m-i-1}-B_{\infty})({\mathbf{g}}^{(k+i)}-\nabla f({\mathbf{x}}^{(k+i)}))\|^{2}]$
$\displaystyle\leq\sum_{i=1}^{m-1}\mathbb{E}_{k}[\|\pi_{A}\|^{2}\|(B^{m-i-1}-B_{\infty})({\mathbf{g}}^{(k+i)}-\nabla f({\mathbf{x}}^{(k+i)}))\|_{F}^{2}]$
$\displaystyle\leq\sum_{i=1}^{m-1}\|B^{m-i-1}-B_{\infty}\|_{2}^{2}\mathbb{E}_{k}[\|{\mathbf{g}}^{(k+i)}-\nabla f({\mathbf{x}}^{(k+i)})\|_{F}^{2}]$
$\displaystyle\leq n\sigma^{2}(\sum_{i=1}^{m-1}\|B^{m-i-1}-B_{\infty}\|_{2})^{2}\leq ns_{B}^{2}\sigma^{2},$
and
$\displaystyle\quad\mathbb{E}_{k}[\|\pi_{A}^{\top}(\sum_{i=1}^{m-1}(B^{m-i-1}-B_{\infty}))({\mathbf{g}}^{(k)}-\nabla f({\mathbf{x}}^{(k)}))\|^{2}]$
$\displaystyle\leq\mathbb{E}_{k}[\|\pi_{A}\|^{2}\|(\sum_{i=1}^{m-1}(B^{m-i-1}-B_{\infty}))({\mathbf{g}}^{(k)}-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
$\displaystyle\leq\|\sum_{i=1}^{m-1}(B^{m-i-1}-B_{\infty})\|_{2}^{2}\mathbb{E}_{k}[\|{\mathbf{g}}^{(k)}-\nabla f({\mathbf{x}}^{(k)})\|_{F}^{2}]$
$\displaystyle\leq(\sum_{i=1}^{m-1}\|B^{m-i-1}-B_{\infty}\|_{2})^{2}n\sigma^{2}\leq ns_{B}^{2}\sigma^{2}.$
By combining the two parts, we finish the proof. $\square$
B.2 Proof of Lemma 4
For simplicity, we define $\Delta_{y}^{(k)}:=(I-B_{\infty}){\mathbf{y}}^{(k)}$ and $\overline{\nabla f}^{(k)}:=\frac{1}{n}\sum_{i=1}^{n}\nabla f_{i}(x^{(k)}_{i})$. According to Assumption 3, we have
$\mathbb{E}[\bar{g}^{(k+i)}]=\mathbb{E}[\mathbb{E}_{k}[\bar{g}^{(k+i)}]]=\mathbb{E}[\overline{\nabla f}^{(k+i)}],\quad\forall i\geq 0.$
Note that
$\pi_{A}^{\top}\sum_{i=0}^{m-1}{\mathbf{y}}^{(k+i)}=c\sum_{i=0}^{m-1}\bar{g}^{(k+i)}+\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)},$
the descent lemma on $\hat{x}^{(k+m)}=\hat{x}^{(k)}-\alpha\pi_{A}^{\top}\sum_{i=k}^{k+m-1}{\mathbf{y}}^{(i)}$ can be written as follows:
$\displaystyle f(\hat{x}^{(k+m)})-f(\hat{x}^{(k)})$ $\displaystyle\leq-c\alpha\left\langle\sum_{i=0}^{m-1}\bar{g}^{(k+i)},\nabla f(\hat{x}^{(k)})\right\rangle-\alpha\left\langle\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)},\nabla f(\hat{x}^{(k)})\right\rangle$
$\displaystyle\quad+c^{2}\alpha^{2}L\|\sum_{i=0}^{m-1}\bar{g}^{(k+i)}\|^{2}+\alpha^{2}L\|\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}\|^{2}.$
By taking expectation (from time 00 0), we obtain that
$\displaystyle\mathbb{E}[f(\hat{x}^{(k+m)})-f(\hat{x}^{(k)})]\leq-c\alpha\sum_{i=0}^{m-1}\mathbb{E}\left[\left\langle\overline{\nabla f}^{(k+i)},\nabla f(\hat{x}^{(k)})\right\rangle\right]$ (18)
$\displaystyle\quad-\alpha\mathbb{E}\left[\left\langle\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)},\nabla f(\hat{x}^{(k)})\right\rangle\right]+c^{2}\alpha^{2}L\mathbb{E}[\|\sum_{i=0}^{m-1}\bar{g}^{(k+i)}\|^{2}]$
$\displaystyle\quad+\alpha^{2}L\mathbb{E}[\|\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}\|^{2}].$
The first inner product term in (18) can be bounded by:
$\displaystyle-\left\langle\overline{\nabla f}^{(k+i)},\nabla f(\hat{x}^{(k)})\right\rangle$ $\displaystyle=-\frac{1}{2}\|\overline{\nabla f}^{(k+i)}\|^{2}-\frac{1}{2}\|\nabla f(\hat{x}^{(k)})\|^{2}+\frac{1}{2}\|\overline{\nabla f}^{(k+i)}-\nabla f(\hat{x}^{(k)})\|^{2}$
$\displaystyle\leq-\frac{1}{2}\|\overline{\nabla f}^{(k+i)}\|^{2}-\frac{1}{2}\|\nabla f(\hat{x}^{(k)})\|^{2}+\frac{L^{2}}{2}\|{\mathbf{x}}^{(k+i)}-\hat{{\mathbf{x}}}^{(k)}\|_{F}^{2},$
where the last inequality comes from Cauchy-Schwarz inequality and $L$-smoothness. The first quartic term in (18) can be bounded by:
$\mathbb{E}[\|\sum_{i=0}^{m-1}\bar{g}^{(k+i)}\|^{2}]\leq n^{-1}m\sigma^{2}+\mathbb{E}[\|\sum_{i=0}^{m-1}\overline{\nabla f}^{(k+i)}\|^{2}],$
where we use the fact that $\bar{g}^{(k+i)}$ are conditional unbiased estimate and
$\mathbb{E}\left[\left\langle\bar{g}^{(k+i)}-\overline{\nabla f}^{(k+i)},\bar{g}^{(k+j)}-\overline{\nabla f}^{(k+j)}\right\rangle\right]=0.$
Plug these estimations into (18), we obtain that
$\displaystyle\frac{\alpha cm}{2}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]+\frac{\alpha c}{2}\sum_{i=0}^{m-1}\mathbb{E}[\|\overline{\nabla f}^{(k+i)}\|^{2}]\leq\mathbb{E}[f(\hat{x}^{(k)})-f(\hat{x}^{(k+m)})]$ (19)
$\displaystyle+\frac{c\alpha L^{2}}{2}\sum_{i=0}^{m-1}\mathbb{E}[\|{\mathbf{x}}^{(k+i)}-\hat{{\mathbf{x}}}^{(k)}\|_{F}^{2}]+\frac{\alpha^{2}c^{2}mL}{n}\sigma^{2}+c^{2}\alpha^{2}L\mathbb{E}[\|\sum_{i=0}^{m-1}\overline{\nabla f}^{(k+i)}\|^{2}]$
$\displaystyle-\alpha\mathbb{E}\left[\left\langle\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)},\nabla f(\hat{x}^{(k)})\right\rangle\right]+\alpha^{2}L\mathbb{E}[\|\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}\|^{2}].$
Next, we try to simplify the last quadratic term in (19) . Notice that
$\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}=\left(\sum_{j=0}^{m-1}(B^{j}-B_{\infty})\right)\Delta_{y}^{(k)}+\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})({\mathbf{g}}^{(k+j)}-{\mathbf{g}}^{(k)}),$
the quadratic term can be bounded as:
$\displaystyle\mathbb{E}[\|\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}\|^{2}]\leq\mathbb{E}[\|\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}\|_{F}^{2}]$ (20)
$\displaystyle\leq 2\|\sum_{j=0}^{m-1}(B^{j}-B_{\infty})\|_{2}^{2}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]+2\mathbb{E}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})({\mathbf{g}}^{(k+j)}-{\mathbf{g}}^{(k)})\|_{F}^{2}].$
The first term here can be estimated using the fact that $\|\sum_{j=0}^{m-1}(B^{j}-B_{\infty})\|_{2}^{2}\leq s_{B}^{2}$. For the second term in (20), according to Assumption 3, we have
$\displaystyle\quad\mathbb{E}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})({\mathbf{g}}^{(k+j)}-{\mathbf{g}}^{(k)})\|_{F}^{2}]$ (21)
$\displaystyle=\sum_{j=1}^{m-1}\mathbb{E}[\|(B^{m-j-1}-B_{\infty})({\mathbf{g}}^{(k+j)}-\nabla f({\mathbf{x}}^{(k+j)}))\|_{F}^{2}]$
$\displaystyle\quad+\mathbb{E}[\|\left(\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})\right)({\mathbf{g}}^{(k)}-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
$\displaystyle\quad+\mathbb{E}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
$\displaystyle\leq n\sigma^{2}\sum_{j=1}^{m-1}\|B^{m-j-1}-B_{\infty}\|_{2}^{2}+n\sigma^{2}\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})\|_{2}^{2}$
$\displaystyle\quad+\mathbb{E}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
$\displaystyle\leq 2ns_{B}^{2}\sigma^{2}+\mathbb{E}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
The inner product can produce a similar term. According to Lemma 3,
$\mathbb{E}[\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}]=\mathbb{E}[H_{m}({\mathbf{y}}^{(k)},\nabla f({\mathbf{x}}^{(k)},\ldots,\nabla f({\mathbf{x}}^{(k+m-1)}))].$
Therefore, we have
$\displaystyle\quad-\alpha\mathbb{E}\left[\left\langle\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)},\nabla f(\hat{x}^{(k)})\right\rangle\right]$ (22)
$\displaystyle=-\alpha\mathbb{E}\left[\left\langle\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)})),\nabla f(\hat{x}^{(k)})\right\rangle\right]$
$\displaystyle\quad+\mathbb{E}\left[\left\langle\pi_{A}^{\top}\left(\sum_{i=0}^{m-1}(B^{m-i-1}-B_{\infty})\right){\mathbf{y}}^{(k)},\nabla f(\hat{x}^{(k)})\right\rangle\right]$
$\displaystyle\leq\frac{1}{cmL}\mathbb{E}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
$\displaystyle\quad+\alpha^{2}Ls_{B}^{2}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]+\frac{\alpha cm}{2}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}].$
Plug (20) and (22) into (19), we obtain that
$\displaystyle\quad\frac{\alpha cm}{4}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]+\frac{\alpha c}{2}\sum_{i=0}^{m-1}\mathbb{E}[\|\overline{\nabla f}^{(k+i)}\|^{2}]-2c^{2}\alpha^{2}L\mathbb{E}[\|\sum_{i=0}^{m-1}\overline{\nabla f}^{(k+i)}\|^{2}]$ (23)
$\displaystyle\leq\mathbb{E}[f(\hat{x}^{(k)})-f(\hat{x}^{(k+m)})]+\frac{c\alpha L^{2}}{2}\sum_{i=0}^{m-1}\mathbb{E}[\|{\mathbf{x}}^{(k+i)}-\hat{{\mathbf{x}}}^{(k)}\|_{F}^{2}]$
$\displaystyle+\frac{2\alpha^{2}c^{2}mL}{n}\sigma^{2}+2\alpha^{2}Lns_{B}^{2}\sigma^{2}+3\alpha^{2}Ls_{B}^{2}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]$
$\displaystyle+(2\alpha^{2}L+\frac{1}{cmL})\mathbb{E}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
For $2c^{2}\alpha^{2}L\mathbb{E}[\|\sum_{i=0}^{m-1}\overline{\nabla f}^{(k+i)}\|^{2}]$, we have the following estimate:
$\displaystyle\quad\mathbb{E}[\|\sum_{i=0}^{m-1}\overline{\nabla f}^{(k+i)}\|^{2}]\leq m\sum_{i=0}^{m-1}\mathbb{E}[\|\overline{\nabla f}^{(k+i)}\|^{2}].$ (24)
Take (24) into (23), we obtain that
$\displaystyle\quad\frac{\alpha cm}{4}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]+\frac{\alpha c(1-4cm\alpha L)}{2}\sum_{i=0}^{m-1}\mathbb{E}[\|\overline{\nabla f}^{(k+i)}\|^{2}]$ (25)
$\displaystyle\leq\mathbb{E}[f(\hat{x}^{(k)})-f(\hat{x}^{(k+m)})]+\frac{c\alpha L^{2}}{2}\sum_{i=0}^{m-1}\mathbb{E}[\|{\mathbf{x}}^{(k+i)}-\hat{{\mathbf{x}}}^{(k)}\|_{F}^{2}]$
$\displaystyle+\frac{2\alpha^{2}c^{2}mL}{n}\sigma^{2}+2\alpha^{2}Lns_{B}^{2}\sigma^{2}+3\alpha^{2}Ls_{B}^{2}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]$
$\displaystyle+(2\alpha^{2}L+\frac{1}{cmL})\mathbb{E}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
We can further apply Jensen inequality on the last term in (25), i.e.
$\displaystyle\quad\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}$
$\displaystyle\leq(\sum_{j=1}^{m-1}\|B^{m-j-1}-B_{\infty}\|_{2})^{2}\sum_{j=1}^{m-1}\|\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)})\|_{F}^{2}$
$\displaystyle\leq s_{B}^{2}\sum_{j=1}^{m-1}\|\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)})\|_{F}^{2}\leq s_{B}^{2}L^{2}\sum_{j=1}^{m-1}\|{\mathbf{x}}^{(k+j)}-{\mathbf{x}}^{(k)}\|_{F}^{2}.$
Therefore, (25) can be further simplified to
$\displaystyle\quad\frac{\alpha cm}{4}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]+\frac{\alpha c(1-4cm\alpha L)}{2}\sum_{i=0}^{m-1}\mathbb{E}[\|\overline{\nabla f}^{(k+i)}\|^{2}]$
$\displaystyle\leq\mathbb{E}[f(\hat{x}^{(k)})-f(\hat{x}^{(k+m)})]+\frac{c\alpha L^{2}}{2}\sum_{i=0}^{m-1}\mathbb{E}[\|{\mathbf{x}}^{(k+i)}-\hat{{\mathbf{x}}}^{(k)}\|_{F}^{2}]$
$\displaystyle+\frac{2\alpha^{2}c^{2}mL}{n}\sigma^{2}+2\alpha^{2}Lns_{B}^{2}\sigma^{2}+3\alpha^{2}Ls_{B}^{2}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]$
$\displaystyle+(2\alpha^{2}L+\frac{1}{cmL})s_{B}^{2}L^{2}\sum_{j=1}^{m-1}\|{\mathbf{x}}^{(k+j)}-{\mathbf{x}}^{(k)}\|_{F}^{2}.$
By taking $\alpha\leq\min\{\frac{1}{8cmL},\frac{c}{8n^{2}s_{B}^{2}L}\}$, this can be further simplified to
$\displaystyle\quad\frac{\alpha cm}{4}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]+\frac{\alpha c}{4}\sum_{i=0}^{m-1}\mathbb{E}[\|\overline{\nabla f}^{(k+i)}\|^{2}]$ (26)
$\displaystyle\leq\mathbb{E}[f(\hat{x}^{(k)})-f(\hat{x}^{(k+m)})]+\frac{L}{16m}\sum_{i=0}^{m-1}\mathbb{E}[\|{\mathbf{x}}^{(k+i)}-\hat{{\mathbf{x}}}^{(k)}\|_{F}^{2}]$
$\displaystyle+\frac{2\alpha^{2}c^{2}mL}{n}\sigma^{2}+2\alpha^{2}Lns_{B}^{2}\sigma^{2}+3\alpha^{2}Ls_{B}^{2}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]$
$\displaystyle+\frac{L}{16m}\sum_{j=1}^{m-1}\mathbb{E}[\|{\mathbf{x}}^{(k+j)}-{\mathbf{x}}^{(k)}\|_{F}^{2}],$
which finishes the proof.
B.3 Proof of Lemma 5
We consider how to measure the difference of ${\mathbf{x}}$. Use the fact that
$\displaystyle{\mathbf{x}}^{(k+i)}-\hat{{\mathbf{x}}}^{(k)}$ $\displaystyle=(A^{i}-A_{\infty})\Delta_{x}^{(k)}-\alpha\left(\sum_{j=0}^{i-1}(B^{j}-B_{\infty})\right)\Delta_{y}^{(k)}$
$\displaystyle\quad-\alpha\sum_{j=1}^{i-1}(B^{i-j-1}-B_{\infty})({\mathbf{g}}^{(k+i)}-{\mathbf{g}}^{(k)}),$
We decompose ${\mathbf{x}}^{(k+i)}-\hat{{\mathbf{x}}}^{(k)}$ and apply the Cauchy-Schwarz inequality:
$\displaystyle\quad\sum_{i=1}^{m-1}\mathbb{E}_{k}[\|{\mathbf{x}}^{(k+i)}-\hat{{\mathbf{x}}}^{(k)}\|_{F}^{2}]\leq 3\left(\sum_{i=1}^{m-1}\|A^{i}-A_{\infty}\|_{2}^{2}\right)\|\Delta_{x}^{(k)}\|_{F}^{2}$ (27)
$\displaystyle\quad+3\alpha^{2}\sum_{i=1}^{m-1}\|\sum_{j=0}^{i-1}(B^{j}-B_{\infty})\|_{2}^{2}\|\Delta_{y}^{(k)}\|_{F}^{2}$
$\displaystyle\quad+3\alpha^{2}\sum_{i=1}^{m-1}\mathbb{E}_{k}[\|\sum_{j=1}^{i-1}(B^{i-j-1}-B_{\infty})({\mathbf{g}}^{(k+i)}-{\mathbf{g}}^{(k)})\|_{F}^{2}]$
$\displaystyle\leq 3s_{A}^{2}\|\Delta_{x}^{(k)}\|_{F}^{2}+3\alpha^{2}ms_{B}^{2}\|\Delta_{y}^{(k)}\|_{F}^{2}+6\alpha^{2}mns_{B}^{2}\sigma^{2}$
$\displaystyle\quad+3\alpha^{2}\sum_{i=1}^{m-1}\mathbb{E}_{k}[\|\sum_{j=1}^{i-1}(B^{i-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
$\displaystyle\leq 3s_{A}^{2}\|\Delta_{x}^{(k)}\|_{F}^{2}+3\alpha^{2}ms_{B}^{2}\|\Delta_{y}^{(k)}\|_{F}^{2}+6\alpha^{2}mns_{B}^{2}\sigma^{2}+3\alpha^{2}s_{B}^{2}L^{2}\Delta_{2}^{(k)},$
where the last inequality comes from Cauchy-Schwarz inequality and $L$-smoothness.
For ${\mathbf{x}}^{(k+i)}-{\mathbf{x}}^{(k)}$, similarly, we decompose it into
$\displaystyle{\mathbf{x}}^{(k+i)}-{\mathbf{x}}^{(k)}$ $\displaystyle=(A^{i}-A_{\infty})\Delta_{x}^{(k)}-\alpha\left(\sum_{j=0}^{i-1}(B^{j}-B_{\infty})\right)\Delta_{y}^{(k)}$
$\displaystyle\quad-\alpha\sum_{j=1}^{i-1}(B^{i-j-1}-B_{\infty})({\mathbf{g}}^{(k+i)}-{\mathbf{g}}^{(k)})+\Delta_{x}^{(k)},$
which finally produces
$\displaystyle\quad\sum_{i=1}^{m-1}\mathbb{E}_{k}[\|{\mathbf{x}}^{(k+i)}-{\mathbf{x}}^{(k)}\|_{F}^{2}]\leq 4s_{A}^{2}\|\Delta_{x}^{(k)}\|_{F}^{2}+4\alpha^{2}ms_{B}^{2}\|\Delta_{y}^{(k)}\|_{F}^{2}$ (28)
$\displaystyle\quad+8\alpha^{2}mns_{B}^{2}\sigma^{2}+4\alpha^{2}s_{B}^{2}L^{2}\sum_{i=1}^{m-1}\mathbb{E}_{k}[\|{\mathbf{x}}^{(k+i)}-{\mathbf{x}}^{(k)}\|_{F}^{2}]+4m\|\Delta_{x}^{(k)}\|_{F}^{2}.$
Inequalities (27) and (28) tell us
$\displaystyle\begin{pmatrix}1&-3\alpha^{2}s_{B}^{2}L^{2}\\ 0&1-4\alpha^{2}s_{B}^{2}L^{2}\end{pmatrix}\begin{pmatrix}\mathbb{E}_{k}[\Delta_{1}^{(k)}]\\ \mathbb{E}_{k}[\Delta_{2}^{(k)}]\end{pmatrix}\leq\begin{pmatrix}v_{1}\\ v_{2}\end{pmatrix}$
where $v_{1}=3s_{A}^{2}\|\Delta_{x}^{(k)}\|_{F}^{2}+3\alpha^{2}ms_{B}^{2}\|\Delta_{y}^{(k)}\|_{F}^{2}+6\alpha^{2}mns_{B}^{2}\sigma^{2}$, $v_{2}=4s_{A}^{2}\|\Delta_{x}^{(k)}\|_{F}^{2}+4\alpha^{2}ms_{B}^{2}\|\Delta_{y}^{(k)}\|_{F}^{2}+8\alpha^{2}mns_{B}^{2}\sigma^{2}+4m\|\Delta_{x}^{(k)}\|_{F}^{2}$ are positive numbers. To proceed on, we assume $\alpha\leq\frac{1}{6s_{B}L}$, so that the inversion of the $2\times 2$ coefficient matrix satisfies
$\displaystyle\begin{pmatrix}1&0\\ 0&1\end{pmatrix}\leq{\begin{pmatrix}1&-3\alpha^{2}s_{B}^{2}L^{2}\\ 0&1-4\alpha^{2}s_{B}^{2}L^{2}\end{pmatrix}}^{-1}\leq\begin{pmatrix}1&0.1\\ 0&1.5\end{pmatrix}.$
Therefore, we have
$\displaystyle{\begin{pmatrix}\frac{L}{16m}\\ \frac{L}{16m}\end{pmatrix}}^{\top}\begin{pmatrix}\mathbb{E}_{k}[\Delta_{1}^{(k)}]\\ \mathbb{E}_{k}[\Delta_{2}^{(k)}]\end{pmatrix}\leq{\begin{pmatrix}\frac{L}{16m}\\ \frac{L}{16m}\end{pmatrix}}^{\top}\begin{pmatrix}1&0.1\\ 0&1.5\end{pmatrix}\begin{pmatrix}v_{1}\\ v_{2}\end{pmatrix}\leq{\begin{pmatrix}\frac{L}{10m}\\ \frac{L}{10m}\end{pmatrix}}^{\top}\begin{pmatrix}v_{1}\\ v_{2}\end{pmatrix}.$
This gives that
$\displaystyle\quad\frac{L}{16m}\mathbb{E}_{k}[\Delta_{1}^{(k)}+\Delta_{2}^{(k)}]\leq\frac{s_{A}^{2}L}{m}\|\Delta_{x}^{(k)}\|_{F}^{2}+\alpha^{2}Ls_{B}^{2}\|\Delta_{y}^{(k)}\|_{F}^{2}+2\alpha^{2}Lns_{B}^{2}\sigma^{2},$
which finishes the proof of Lemma 5.
B.4 Proof of Lemma 6
In this section, we provide some inequalities for the consensus error $\Delta_{x}$ and $\Delta_{y}$. These inequalities will only be used in the final part of proof of our main theorem.
Lemma 8 (Consensus of x).
Suppose Assumption 43 and Proposition 1 hold. For any $T\geq 0$, we have
$\displaystyle\sum_{k=0}^{T}\mathbb{E}[\|\Delta_{x}^{(k+1)}\|^{2}]$ $\displaystyle\leq 2\alpha^{2}s_{A}^{2}\sum_{k=0}^{T}\mathbb{E}[\|\Delta_{y}^{(k)}\|^{2}]$
$\displaystyle\quad+2\alpha^{2}n^{2}s_{A}^{2}\sum_{k=0}^{T}\mathbb{E}[\|\overline{\nabla f}^{(k)}\|^{2}]+2n\alpha^{2}s_{A}^{2}(T+1)\sigma^{2}$
Proof.
First note that
$\displaystyle\Delta_{x}^{(k+1)}$ $\displaystyle=(I-A_{\infty}){\mathbf{x}}^{(k+1)}=(A-A_{\infty}){\mathbf{x}}^{(k)}-(I-A_{\infty}){\mathbf{y}}^{(k)}$
$\displaystyle=\ldots=\sum_{i=0}^{k}(A^{k-i}-A_{\infty}){\mathbf{y}}^{(i)}$
$\displaystyle=\sum_{i=0}^{k}(A^{k-i}-A_{\infty})\Delta_{y}^{(i)}+\sum_{i=0}^{k}(A^{k-i}-A_{\infty})n\pi_{B}\bar{g}^{(i)}$
Using the rolling sum lemma, we have
$\displaystyle\sum_{k=0}^{T}\|\Delta_{x}^{(k+1)}\|_{F}^{2}\leq 2s_{A}^{2}\sum_{k=0}^{T}\|\Delta_{y}^{(k+1)}\|_{F}^{2}+2n^{2}s_{A}^{2}\|\pi_{B}\|^{2}\sum_{k=0}^{T}\|\bar{g}^{(k)}\|_{F}^{2}.$
Note that $\mathbb{E}[\|\bar{g}^{(k)}\|_{F}^{2}]\leq n^{-1}\sigma^{2}+\mathbb{E}[\|\overline{\nabla f}^{(k)}\|_{F}^{2}]$ and $\|\pi_{B}\|\leq 1$, we finish the proof. ∎
Lemma 9 (Consensus of y).
Suppose Assumption 43 and Proposition 1 hold. For any $T\geq 0$, when $\alpha\leq\frac{1}{6\sqrt{n}s_{A}s_{B}L}$, we have
$\displaystyle\sum_{k=0}^{T}\|\Delta_{y}^{(k+1)}\|_{F}^{2}\leq 30ns_{B}^{2}(T+1)\sigma^{2}+8ns_{B}^{2}L\Delta+108\alpha^{2}L^{2}n^{3}s_{A}^{2}s_{B}^{2}\sum_{k=0}^{T}\mathbb{E}[\|\overline{\nabla f}^{(k)}\|_{F}^{2}]$
Proof.
$\displaystyle\Delta_{y}^{(k+1)}$ $\displaystyle=(B-B_{\infty})\Delta_{y}^{(k)}+(I-B_{\infty})({\mathbf{g}}^{(k+1)}-{\mathbf{g}}^{(k)})$
$\displaystyle=\ldots=(B^{k+1}-B_{\infty}){\mathbf{g}}^{(0)}+\sum_{i=0}^{k}(B^{k-i}-B_{\infty})({\mathbf{g}}^{(i+1)}-{\mathbf{g}}^{(i)})$
Using the rolling sum lemma, we have
$\displaystyle\sum_{k=0}^{T}\|\Delta_{y}^{(k+1)}\|_{F}^{2}\leq s_{B}^{2}\|{\mathbf{g}}^{(0)}\|+s_{B}^{2}\sum_{k=0}^{T}\|{\mathbf{g}}^{(k+1)}-{\mathbf{g}}^{(k)}\|_{F}^{2}$
By Cauchy-Schwarz Inequality and our assumption for bounded noise, we have
$\displaystyle\mathbb{E}[\|{\mathbf{g}}^{(k+1)}-{\mathbf{g}}^{(k)}\|_{F}^{2}]$ $\displaystyle\leq 6n\sigma^{2}+3\mathbb{E}[\|\nabla f({\mathbf{x}}^{(k+1)})-\nabla f({\mathbf{x}}^{(k)})\|_{F}^{2}]$
$\displaystyle\leq 6n\sigma^{2}+3L^{2}\mathbb{E}[\|{\mathbf{x}}^{(k+1)}-{\mathbf{x}}^{(k)}\|_{F}^{2}]$
Besides, $\mathbb{E}[\|{\mathbf{g}}^{(0)}\|_{F}^{2}]\leq n\sigma^{2}+\|\nabla f({\mathbf{x}}^{(0)})\|_{F}^{2}\leq n\sigma^{2}+2nL\Delta$. Therefore, we have
$\displaystyle\sum_{k=0}^{T}\|\Delta_{y}^{(k+1)}\|_{F}^{2}\leq ns_{B}^{2}(6T+7)\sigma^{2}+2ns_{B}^{2}L\Delta+3s_{B}^{2}L^{2}\sum_{k=0}^{T}\mathbb{E}[\|{\mathbf{x}}^{(k+1)}-{\mathbf{x}}^{(k)}\|_{F}^{2}].$ (29)
Since
$\displaystyle\quad\mathbb{E}[\|{\mathbf{x}}^{(k+1)}-{\mathbf{x}}^{(k)}\|_{F}^{2}]=\mathbb{E}[\|(A-I){\mathbf{x}}^{(k)}-\alpha{\mathbf{y}}^{(k)}\|_{F}^{2}]$ (30)
$\displaystyle=\mathbb{E}[\|(A-I)\Delta_{x}^{(k)}-\alpha\Delta_{y}^{(k)}+n\alpha\pi_{B}\bar{g}^{(k)}\|_{F}^{2}]$
$\displaystyle\leq 3\|A-I\|_{2}^{2}\mathbb{E}[\|\Delta_{x}^{(k)}\|_{F}^{2}]+3\alpha^{2}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]+3n\alpha^{2}\sigma^{2}+3n^{2}\alpha^{2}\mathbb{E}[\|\overline{\nabla f}^{(k)}\|_{F}^{2}],$
and $\|A-I\|_{2}^{2}\leq 2n$, we can apply (30) to substitute the last term in (29). This produces
$\displaystyle(1-9\alpha^{2}L^{2}s_{B}^{2})\sum_{k=0}^{T}\|\Delta_{y}^{(k+1)}\|_{F}^{2}\leq ns_{B}^{2}(6T+7+9\alpha^{2}L^{2}T)\sigma^{2}$
$\displaystyle\quad+9ns_{B}^{2}L^{2}\sum_{k=0}^{T}\mathbb{E}[\|\Delta_{x}^{(k)}\|_{F}^{2}]+9\alpha^{2}L^{2}n^{2}s_{B}^{2}\sum_{k=0}^{T}\mathbb{E}[\|\overline{\nabla f}^{(k)}\|_{F}^{2}]+2ns_{B}^{2}L\Delta$
Using the conclusion of Lemma 8, this can be further transformed to
$\displaystyle\quad(1-9\alpha^{2}L^{2}s_{B}^{2}-18\alpha^{2}L^{2}ns_{A}^{2}s_{B}^{2})\sum_{k=0}^{T}\|\Delta_{y}^{(k+1)}\|_{F}^{2}$
$\displaystyle\leq ns_{B}^{2}(6T+7+9\alpha^{2}L^{2}T+18\alpha^{2}L^{2}ns_{A}^{2}T)\sigma^{2}+2ns_{B}^{2}L\Delta$
$\displaystyle\quad+(9\alpha^{2}L^{2}n^{2}s_{B}^{2}+18\alpha^{2}L^{2}n^{3}s_{A}^{2}s_{B}^{2})\sum_{k=0}^{T}\mathbb{E}[\|\overline{\nabla f}^{(k)}\|_{F}^{2}]$
When $\alpha\leq\frac{1}{6\sqrt{n}s_{A}s_{B}L}$, this can be simplified to
$\displaystyle\sum_{k=0}^{T}\|\Delta_{y}^{(k+1)}\|_{F}^{2}\leq 30ns_{B}^{2}(T+1)\sigma^{2}+8ns_{B}^{2}L\Delta+108\alpha^{2}L^{2}n^{3}s_{A}^{2}s_{B}^{2}\sum_{k=0}^{T}\mathbb{E}[\|\overline{\nabla f}^{(k)}\|_{F}^{2}]$
B.5 Proof of Theorem 1
Plug Lemma 5 into Lemma 4, we obtain that
$\displaystyle\quad\frac{\alpha cm}{4}\|\nabla f(\hat{x}^{(k)})\|^{2}+\frac{\alpha c(1-4\alpha mcL)}{2}\sum_{i=0}^{m-1}\mathbb{E}_{k}[\|\overline{\nabla f}^{(k+i)}\|^{2}]$ (31)
$\displaystyle\leq f(\hat{x}^{(k)})-\mathbb{E}_{k}[f(\hat{x}^{(k+m)})]+\left(\frac{2\alpha^{2}c^{2}mL}{n}+4\alpha^{2}Lns_{B}^{2}\right)\sigma^{2}$
$\displaystyle\quad+\frac{s_{A}^{2}L}{m}\|\Delta_{x}^{(k)}\|_{F}^{2}+4\alpha^{2}Ls_{B}^{2}\|\Delta_{y}^{(k)}\|_{F}^{2}.$
When $\alpha\leq\frac{1}{8cmL}$, $\frac{\alpha c(1-4\alpha mcL)}{2}$ can be replaced with $\frac{c\alpha}{4}$. Sum up (31) for $k\in[m^{*}]$, take full expectation and divide both sides by $\frac{c\alpha T}{4}$ we obtain that
$\displaystyle\quad\frac{1}{K}\sum_{k\in[m^{*}]}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]+\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|^{2}]$ (32)
$\displaystyle\leq\frac{4\Delta}{c\alpha T}+\left(\frac{2c\alpha L}{n}+\frac{4\alpha Lns_{B}^{2}}{cm}\right)\sigma^{2}$
$\displaystyle\quad+\frac{4s_{A}^{2}L}{cm\alpha T}\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{x}^{(k)}\|_{F}^{2}]+\frac{16\alpha Ls_{B}^{2}}{cmT}\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}].$
Since $\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{x}^{(k)}\|_{F}^{2}]\leq\sum_{k=0}^{T-1}\|\Delta_{x}^{(k+1)}\|_{F}^{2}$, we can plug in Lemma 8 into (32), and obtain that
$\displaystyle\quad\frac{1}{K}\sum_{k\in[m^{*}]}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]+\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|^{2}]$ (33)
$\displaystyle\leq\frac{4\Delta}{c\alpha T}+\left(\frac{2c\alpha L}{n}+\frac{8\alpha Ln(s_{A}^{4}+s_{B}^{2})}{cm}\right)\sigma^{2}$
$\displaystyle\quad+\frac{8\alpha Ln^{2}s_{A}^{4}}{cmT}\sum_{t=0}^{T-1}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|^{2}]+\frac{16\alpha L(s_{A}^{4}+s_{B}^{2})}{cmT}\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}].$
Similarly, since $\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]\leq\sum_{k=0}^{T-1}\|\Delta_{y}^{(k+1)}\|_{F}^{2}$, we further plug in Lemma 9 into (33), and obtain that
$\displaystyle\quad\frac{1}{K}\sum_{k\in[m^{*}]}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]+\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|^{2}]$ (34)
$\displaystyle\leq\frac{4\Delta}{c\alpha T}+C_{\sigma}\sigma^{2}+\frac{128ns_{B}^{2}\alpha L^{2}(s_{A}^{4}+s_{B}^{2})\Delta}{cmT}+C_{f}\sum_{t=0}^{T-1}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|^{2}].$
where
$\displaystyle C_{\sigma}$ $\displaystyle=\frac{2c\alpha L}{n}+\frac{8\alpha Ln(s_{A}^{4}+s_{B}^{2})}{cm}+\frac{480\alpha Lns_{B}^{2}(s_{A}^{4}+s_{B}^{2})}{cm},$
$\displaystyle C_{f}$ $\displaystyle=\frac{8\alpha Ln^{2}s_{A}^{4}}{cmT}+\frac{2000\alpha^{3}L^{3}n^{3}s_{A}^{2}s_{B}^{2}(s_{A}^{4}+s_{B}^{2})}{cmT}.$
When $m\geq 20n^{2}c^{-2}s_{B}^{2}(s_{A}^{4}+s_{B}^{2})$, we have $C_{\sigma}\leq\frac{30c\alpha L}{n}$ and $C_{f}\leq\frac{2c\alpha L}{5T}+\frac{100c\alpha^{3}L^{3}n^{3}s_{A}^{2}}{T}$. When we further set $\alpha\leq\min\{\frac{1}{8cL},\frac{1}{4ns_{A}L}\}$, we have $C_{f}\leq\frac{1}{T}$. Under these conditions, (34) can be simplified to
$\displaystyle\quad\frac{1}{K}\sum_{k\in[m^{*}]}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]\leq\frac{4\Delta}{c\alpha T}+\frac{30c\alpha L}{n}\sigma^{2}+\frac{64n\alpha L^{2}(s_{A}^{4}+s_{B}^{2})\Delta}{cmT}.$ (35)
Finally, to satisfy all the constraints on $m$ and $\alpha$, we can set $m=\lceil 20n^{2}c^{-2}s_{B}^{2}(s_{A}^{4}+s_{B}^{2})\rceil$, and
$\alpha=\min\left\{\frac{c}{21n^{2}s_{B}^{2}(s_{A}^{4}+s_{B}^{2})L},\frac{1}{8cL},\frac{\sqrt{n\Delta}\sigma}{3c\sqrt{LT}}\right\}.$
Plugging this into (35) gives that
$\displaystyle\min_{t\in\{0,1,\ldots,T-1\}}\mathbb{E}[\|\nabla f(\hat{x}^{(t)}\|^{2}]\leq\frac{1}{K}\sum_{k\in[m^{*}]}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]\leq\frac{20\sigma\sqrt{L\Delta}}{\sqrt{nT}}+C_{\Delta}\frac{L\Delta}{T},$
where
$C_{\Delta}=32+32ncs_{A}+84n^{2}c^{-2}s_{B}^{2}(s_{A}^{4}+s_{B}^{2})$
is a constant decided by $A$ and $B$.
Appendix C Experiment Details
In this section, we supplement experiment setup details.
C.1 Weight Matrix Design
We use the following procedure to generate a row/column-stochastic weight matrix. For a fixed number of nodes $n$, we begin by initializing a 0-1 adjacency matrix $W$ as the identity matrix $I$ to account for self-loops on each node. Subsequently, based on the specific network topology definition, we set $W_{i,j}=1$ if there is a directed edge from node $i$ to node $j$. After constructing this initial 0-1 matrix $W$, we then randomly assign integer weights from 11 11 to 99 99 to all entries where $W_{i,j}=1$. Finally, we perform either row normalization or column normalization on this weighted matrix to obtain a row-stochastic matrix $A$ or a column-stochastic matrix $B$, respectively.
C.2 Synthetic data
For experiments conducted on the synthetic dataset, the objective function is given by:
$\displaystyle f(x)=\frac{1}{n}\sum_{i=1}^{n}f_{i}(x),\quad x\in\mathbb{R}^{d},$ (36)
where
$f_{i}(x)=\underbrace{\frac{1}{L_{i}}\sum_{l=1}^{L_{i}}\ln\left(1+\exp(-y_{i,l}h_{i,l}^{\top}x)\right)}_{\text{Logistic Loss}}+\underbrace{\rho\sum_{j=1}^{d}\frac{[x]_{j}^{2}}{1+[x]_{j}^{2}}}_{\text{Non-convex Regularization}},\quad L_{i}=L_{\text{Local}}/n.$
Data Synthesis Process:
  1. 1.
    Global Data Generation:
    • Generate optimal parameters: $x_{\text{opt}}\sim\mathcal{N}(0,I_{d})$
    • Create global feature matrix $H\in\mathbb{R}^{L_{\text{total}}\times d}$ with $h_{l,j}\sim\mathcal{N}(0,1)$
    • Compute labels $Y\in\mathbb{R}^{L_{\text{total}}}$ with $y_{l}\in\{-1,+1\}$ via randomized thresholding:
      $y_{l}=\begin{cases}1&\text{if }1/z_{l}>1+\exp(-h_{l}^{\top}x_{\text{opt}})\\ -1&\text{otherwise}\end{cases},\quad z_{l}\sim\mathcal{U}(0,1)$
  2. 2.
    Data Distribution:
    • Partition $H$ and $L$ across $n$ nodes using:
      $H^{(i)}=H[iM:(i+1)M,\;:\;],\;\quad Y^{(i)}=Y[iM:(i+1)M],\;\quad M=L_{\text{total}}/n$
      where $L_{\text{total}}$ must be divisible by $n$
  3. 3.
    Local Initial Points:
    $x_{i}^{\star}=x_{\text{opt}}+\varepsilon_{i},\quad\varepsilon_{i}\sim\mathcal{N}(0,\sigma_{h}^{2}I_{d}),\quad\sigma_{h}=10$
Gradient Computation: At each iteration, each node $i$ independently computes its stochastic gradient by randomly sampling a minibatch of size $B$ from its local dataset of size $L_{i}=L_{\text{total}}/n$. The gradient computation consists of two components:
$\nabla f_{i,B}(x)=\underbrace{-\frac{1}{B}\sum_{b=1}^{B}\frac{y_{i,b}h_{i,b}}{1+\exp(y_{i,b}h_{i,b}^{\top}x)}}_{\text{Logistic Loss Gradient}}+\underbrace{\rho\sum_{j=1}^{d}\frac{2[x]_{j}}{(1+[x]_{j}^{2})^{2}}}_{\text{Non-convex Regularization}}$
where the minibatch $\{h_{i,b},y_{i,b}\}_{b=1}^{B}$ is drawn uniformly from the $L_{i}$ local samples without replacement. Notice that the gradient on each node $i$ is computed on the local parameter $x_{i}\in\mathbb{R}^{d}$.
Implementation Details:
  • Global dataset size $L_{\text{total}}=204800$, batch size $B=200$
  • Dimension $d=10$, regularization $\rho=0.01$
  • Node configurations: $n\in\{4,6,8,9,12,16,18,24,25\}$, and the least common multiple (LCM) is $3,600$.
Notice that $L_{\text{total}}=1,440,000=3,600\cdot 200\cdot 2=\text{LCM}\cdot B\cdot 2$.
Evaluation Metric: We track $\|n^{-1}\mathds{1}^{\top}\nabla f({\mathbf{x}}^{(k)})\|$ as gradient norm, where
$\displaystyle n^{-1}\mathds{1}^{\top}\nabla f({\mathbf{x}}^{(k)})=\frac{1}{n}\sum_{i=1}^{n}\nabla f_{i,M}(\bm{x}_{i}^{(k)}),\;M=L_{\text{Total}}/n.$
The experimental results shown in our plots represent the average performance across 2020 2020 independent repetitions, where each trial is conducted with the fixed random $\text{seed}=42$ for reproducibility.
C.3 Neural Network Experiment
We first employ three-layer fully connected neural networks for MNIST classification.
Data Distribution: For each node $k\in\{1,\dots,n\}$, we sample a class proportion vector
$\mathbf{q}^{k}=(q_{k,0},q_{k,1},\dots,q_{k,9})\sim\text{Dirichlet}(\alpha\cdot\mathbf{1}^{10})$
where $\mathbf{1}^{10}$ is a 1010 1010-dimensional vector of ones. Each $q_{k,c}$ represents the target proportion of samples from class $c$ that node $k$ should ideally hold. By definition of the Dirichlet distribution, $q_{k,c}\geq 0,\forall c$ and $\sum_{c=0}^{9}q_{k,c}=1,\forall k$.
Denote $N_{k,c}^{\text{target}}=q_{k,c}\cdot\frac{M}{n}$ as the target number of samples of class $c$ to be assigned to node $k$. Ignoring the slight deviations caused by non-divisibility, the number of samples assigned to each node $k$ is approximately $60,000/n$. Here we set $\alpha=0.9$ to get the heterogeneous data and the Figure 5 shows the distribution of samples with 1616 1616 nodes. The Figure 2 shows that the Push Pull achieves linear speedup on synthetic data.
Figure 5: Comparison of uniform (left) and heterogeneous (right) data distributions across 1616 1616 nodes ($\alpha=0.9$).
Evaluation Metric: After each batch computation, we first average the gradients obtained from the $n$ neural networks. Then we compute the norm of this averaged gradient. This norm is subsequently normalized by the number of parameters P of a single neural network. In our experiments, $\textbf{P}=1,863,690$.
$\displaystyle\left\|\frac{1}{n}\sum_{i=1}^{n}\mathbf{g}_{i,B_{ij}}\right\|_{2}/\sqrt{\textbf{P}},\text{ the gradient }\mathbf{g}_{i,B_{i}}\text{ is computed on the $\displaystyle j$-th batch of node $\displaystyle i$}$ (37)
The Figure 3 shows that the Push Pull achieves linear speedup on both directed and undirected topologies.
Finally, we trained Resnet18 models on the CIFAR10 dataset to evaluate the performance of our Push-Pull method on a benchmark computer vision task. This involved conducting training across different numbers of nodes ($n>1$) utilizing the three previously described network topologies: directed exponential, grid graphs, and undirected random graphs. The performance was compared to single-node training ($n=1$) without distributed optimization. (see Figure 4).
Implementation Details:
  • Learning rate $lr=0.005$ on each topology and each number of node.
  • Batch size $bs=128$ on each topology and each number of node.