### v3

# On the Linear Speedup of the Push-Pull Method for Decentralized Optimization over Digraphs

Abstract
The linear speedup property is essential for demonstrating the advantage of distributed algorithms over their single-node counterparts. This paper studies the stochastic Push-Pull method, a widely adopted decentralized optimization algorithm for directed graphs (digraphs). Unlike methods that rely on Push-Only or Pull-Only communications, Push-Pull avoids nonlinear correction and has shown superior empirical performance across various settings. Despite its success, whether Push-Pull achieves linear speedup has remained an open question, revealing a significant gap between its empirical performance and theoretical understanding. To bridge this gap, we propose a novel multi-step descent analysis framework and prove that Push-Pull indeed achieves linear speedup over arbitrary strongly connected digraphs. Our results provide a comprehensive theoretical understanding of the stochastic Push-Pull method, finally aligning its theoretical convergence with empirical performance. Code: https://github.com/pkumelon/PushPull.
1 Introduction
Scaling modern machine learning tasks to massive datasets and increasingly large models requires efficient distributed computation across multiple nodes. In distributed optimization, a network of $n$ nodes cooperatively solves
$\displaystyle\min_{x\in\mathbb{R}^{d}}\quad f(x):=\frac{1}{n}\sum_{i=1}^{n}f_{i}(x)\quad\mbox{where}\quad f_{i}(x)=\mathbb{E}_{\xi_{i}\sim\mathcal{D}_{i}}[F_{i}(x;\xi_{i})].$ (1)
where each local function $f_{i}$ is defined by a data distribution $\mathcal{D}_{i}$. Each node has access only to stochastic gradients $\nabla F_{i}(x;\xi_{i})$ from its local data, while communication is required to aggregate information across the network. In practice, the distributions $\{\mathcal{D}_{i}\}_{i=1}^{n}$ are heterogeneous, so the local losses $f_{i}$ generally differ. This heterogeneity poses significant challenges to both the design and analysis of distributed algorithms.
A natural baseline for solving (1) is centralized stochastic gradient descent (SGD), where gradients are aggregated globally either through a central server [1] or collective communication primitives such as Ring All-Reduce [2]. While centralized SGD enjoys strong theoretical guarantees—including the celebrated linear speedup property in iteration complexity—its reliance on global averaging leads to significant communication overhead and creates a single point of failure at the server. These drawbacks limit scalability in large, heterogeneous systems.
Decentralized SGD has emerged as a compelling alternative [3, 4]. In this setting, each node maintains its own local model and communicates only with its neighbors. This localized communication substantially reduces bandwidth usage and alleviates latency bottlenecks. Moreover, decentralized methods are inherently robust to failures: as long as the communication graph remains connected, the network can continue to train without dependence on a single master node.
In this paper, we consider decentralized optimization over directed graphs (digraphs). Directed communication naturally models many real-world systems, including heterogeneous robotic swarms with asymmetric links, sensor networks with one-sided message passing, and distributed deep learning environments where communication is constrained by bandwidth asymmetry. Directed networks are also attractive because certain classes of sparse digraphs, such as exponential graphs [5], EquiTopo [6], and B-ary trees [7], achieve both scalability and strong connectivity.
1.1 Decentralized algorithms over digraphs
In a connected network, the topology can be characterized by a mixing matrix. This matrix reflects how nodes are connected and is important for understanding how the network affects algorithm performance. For undirected networks, it is straightforward to construct symmetric and doubly stochastic mixing matrices. For directed networks, achieving a doubly stochastic matrix is generally infeasible. Consequently, mixing matrices in such cases are typically either column-stochastic [8, 9] or row-stochastic [10, 11], but not both.
Building on the connectivity pattern of the underlying network topology, decentralized optimization algorithms over directed graphs can be broadly categorized into three families in the literature. The first family relies on the Push-Sum technique [12, 13], which employs column-stochastic mixing matrices for all communications. Accordingly, we refer to this class as Push-Only algorithms in this paper. When exact gradients $\nabla f_{i}$ are available, the well-known Subgradient-Push algorithm [8, 13] converges to the desired solution, albeit with relatively slow sublinear rates even under strong convexity. More advanced methods such as EXTRA-Push [14], DEXTRA [15], ADD-OPT [16], and Push-DIGing [9] improve convergence rate by mitigating the effects of data heterogeneity. Stochastic variants, including SGD-Push [5] and S-ADDOPT [17], extend these approaches to settings with stochastic gradients.
The second family consists of algorithms that use only row-stochastic matrices for communication, referred to as the Pull-Only methods. Just as Push-Sum underpins the design of Push-Only algorithms, the Pull-Diag gossip protocol [11] serves as the foundation for Pull-Only methods. Reference [11] adapted the distributed gradient descent algorithm to the Pull-Only setting. Subsequently, gradient tracking techniques were extended to this framework by [18, 19], while momentum-based Pull-Only gradient tracking methods were developed in [20, 21].
The third family comprises algorithms that alternate between row-stochastic and column-stochastic mixing matrices, commonly referred to as the Push-Pull method [22] or the AB method [23]. In this framework, a row-stochastic matrix is used to exchange model parameters, while a column-stochastic matrix is employed to share gradient information. As demonstrated in [23, 22, 24], the Push-Pull family typically achieves significantly faster convergence rates than algorithms relying solely on column- or row-stochastic matrices. Reference [25, 26] extended the Push-Pull method to stochastic optimization settings, while [7] improved its convergence rate on B-ary tree network topologies. Accelerated or constrained variant of the Push-Pull method have also been explored in [27, 28].
1.2 Linear speedup in iteration complexity
The linear speedup is a central property for distributed optimization. It states that the number of iterations required to reach an $\epsilon$-accurate solution decreases linearly with the number of nodes $n$. Centralized SGD achieves this scaling, converging in $\mathcal{O}(1/(n\epsilon^{2}))$ iterations. Without this property, the benefit of parallelization is limited, as increasing the number of nodes would not yield proportional gains in iteration complexity.
For decentralized algorithms on undirected networks, linear speedup is well established. The first rigorous analysis for decentralized SGD was given in [29]. Building on this work, a range of advanced algorithms—including Exact Diffusion/NIDS [30, 31] and gradient tracking methods [32, 33, 34, 9]—have been proven to achieve linear speedup [35]. Time-varying network extensions are also addressed in [36].
For digraphs, theoretical understandings are less explored. For Push-Only algorithms with column-stochastic mixing, linear speedup has been established in [5, 37], with subsequent refinements achieving optimal iteration complexity in [38]. For Pull-Only algorithms with row-stochastic mixing, convergence has been extensively analyzed under various settings [11, 18, 19, 20, 21], but these studies did not establish linear speedup because row-stochastic matrices inherently introduce a deviation from the globally averaged gradient. Only recently, linear speedup for Pull-Only algorithms was proven in [39]. These advances reveal the effectiveness of both Push-Only and Pull-Only methods over digraphs.
Despite this progress, establishing the linear speedup property for Push-Pull algorithms—those combining both row- and column-stochastic matrices—has remained an open problem. While Push-Pull methods consistently outperform Push-Only and Pull-Only variants in practice, their theoretical understanding remains limited. Existing analyses focus mostly on strongly convex settings [23, 25, 22, 7], and provide no guarantees for nonconvex or stochastic optimization. The most relevant result to date is [7], which demonstrates linear speedup for Push-Pull under a specialized B-ary tree topology, but the analysis cannot be extended to general network structures. A recent work [40] establishes guarantees for Push-Pull in a specific non-convex federated learning setting, but does not prove the linear speedup property. Table 1 provides a summary of existing results on decentralized algorithms over digraphs.
1.3 Motivation and main results
The above discussion highlights a striking gap between theory and practice: Push-Pull algorithms consistently outperform Push-Only and Pull-Only methods in empirical studies, yet their scalability theory is far less developed. This gap motivates two fundamental questions:
  • Q1.
    What is the intrinsic difference between the Push-Pull method and the other two families, i.e., Push-Only and Pull-Only? Why is the theoretical analysis of Push-Pull more challenging?
  • Q2.
    Can push–pull achieve linear speedup on general strongly connected digraphs under standard assumptions?
In this work, we address both questions. Regarding Question Q1, we show that Push-Only and Pull-Only can be reformulated to align closely with centralized SGD, making their linear speedup analysis relatively straightforward. In contrast, Push-Pull exhibits a persistent, non-vanishing deviation from centralized SGD dynamics, rendering the conventional single-step descent analysis framework ineffective.
For Question Q2, we introduce a novel multi-step descent analysis framework that aggregates errors over blocks of iterations. This framework uncovers a telescoping structure in the gradient-tracking recursion: intermediate gradient noise cancels out, while boundary terms are exponentially attenuated by mixing dynamics. Consequently, the residual error remains bounded at a constant scale, enabling us to establish the desired $\mathcal{O}(1/(n\epsilon^{2}))$ iteration complexity for Push-Pull.
To the best of our knowledge, this paper presents the first general proof that Push-Pull achieves linear speedup on arbitrary strongly connected digraphs in the stochastic nonconvex setting. In particular, our multi-step descent analysis framework not only bridges the gap between theory and practice for Push-Pull but also introduces techniques that may be of independent interest for the broader study of decentralized algorithms. A summary of existing results is provided in Table 1.
Table 1: Convergence rate comparison of representative decentralized algorithms in non-convex and stochastic settings over directed graphs. “Conv. Rate” refers to the asymptotic rate as $T\to\infty$; “Iter. Complex.” denotes the number of iteration requried to achieve $\epsilon$-accuracy; “S.C.” denotes “Strongly Connected”; “R-Sto.”, “C-Sto.” and “R/C Alt.” refer to “Row-Stochastic”, “Column-Stochastic” and “Row/Column Alternate”, respectively. Note that References [25, 26] study the Push-Pull algorithm under strongly convex objectives, rather than the non-convex setting. $\sigma$ denotes gradient noise.
Algorithms
Conv.
Rate
Iter.
Complex.
Topology
Mixing
Matrix
Centralized SGD $\mathcal{O}(\frac{\sigma}{\sqrt{nT}})$ $\mathcal{O}(\frac{\sigma^{2}}{n\epsilon^{2}})$ Star graph -
Pull-Diag-GT [39] $\mathcal{O}(\frac{\sigma}{\sqrt{nT}})$ $\mathcal{O}(\frac{\sigma^{2}}{n\epsilon^{2}})$ S.C. digraph R-Sto.
Push-Diging [38] $\mathcal{O}(\frac{\sigma}{\sqrt{nT}})$ $\mathcal{O}(\frac{\sigma^{2}}{n\epsilon^{2}})$ S.C. digraph C-Sto.
Push-Pull [25] - $\;\;\;$-\;\; S.C. digraph R/C Alt.
Push-Pull [26] - $\;\;\;$-\;\; S.C. digraph R/C Alt.
Push-Pull [40] $\mathcal{O}(\frac{\sigma}{\sqrt{T}})$ $\mathcal{O}(\frac{\sigma^{2}}{\epsilon^{2}})$ S.C. digraph R/C Alt.
Push-Pull [7] $\mathcal{O}(\frac{\sigma}{\sqrt{nT}})$ $\mathcal{O}(\frac{\sigma^{2}}{n\epsilon^{2}})$ B-ary Tree R/C Alt.
Push-Pull (Ours) $\mathcal{O}(\frac{\sigma}{\sqrt{nT}})$ $\mathcal{O}(\frac{\sigma^{2}}{n\epsilon^{2}})$ S.C. digraph R/C Alt.
The remainder of the paper is organized as follows. Section 2 introduces the notations, assumptions, and performance metrics used throughout the paper. Section 3 reviews the Push-Only, Pull-Only, and Push-Pull methods and provides insights into the derivation of their linear speedup properties. Section 4 presents our novel multi-step descent analysis framework and explains the intuition behind its role in enabling linear speedup analysis. Section 5 formally develops the supporting lemmas and states the main convergence theorem. Section 6 reports extensive numerical experiments that validate the ability of the Push-Pull method to achieve linear speedup. Finally, Section 7 concludes the paper.
2 Notations, Algorithms, Assumptions and Metrics
This section introduces notations, assumptions, and metrics that will be used throughout the paper.
Notations. The directed network is denoted by $\mathcal{G}=(\mathcal{N},\mathcal{E})$, where $\mathcal{N}=\{1,2,\ldots,n\}$ is the set of nodes and $\mathcal{E}$ the set of directed edges. An edge $(j,i)\in\mathcal{E}$ indicates that node $j$ can send information to node $i$. The graph is strongly connected if a directed path exists between every pair of nodes. We say $\mathcal{G}$ is the underlying topology of a nonnegative weight matrix $A=[a_{ij}]$ if $a_{ij}>0$ precisely when $(j,i)\in\mathcal{E}$. Let $\mathds{1}\in\mathbb{R}^{n}$ denote the all-ones vector and $[n]:=\{1,2,\ldots,n\}$. For a vector $v$, $\mathrm{Diag}(v)$ denotes the diagonal matrix with $v$ on its diagonal, while $\min v$ and $\max v$ denote its smallest and largest entries, respectively.
We let $A$ denote a row-stochastic matrix ($A\mathds{1}=\mathds{1}$) and $B$ a column-stochastic matrix ($B^{\top}\mathds{1}=\mathds{1}$), with Perron vectors [41] $\pi_{A}$ and $\pi_{B}$, respectively. Define $A_{\infty}:=\mathds{1}\pi_{A}^{\top}$ and $B_{\infty}:=\pi_{B}\mathds{1}^{\top}$. Additionally, we define $n\times d$ matrices
$\displaystyle{\mathbf{x}}^{(t)}$ $\displaystyle\hskip-1.42262pt:=\hskip-1.42262pt[({x}_{1}^{(t)})^{\top}\hskip-1.42262pt;({x}_{2}^{(t)})^{\top}\hskip-1.42262pt;\cdots;({x}_{n}^{(t)})^{\top}\hskip-1.42262pt]$
$\displaystyle\hskip-1.42262pt{\mathbf{g}}^{(t)}$ $\displaystyle\hskip-1.42262pt:=\hskip-1.42262pt[\nabla F_{1}({x}_{1}^{(t)};\hskip-1.42262pt\xi_{1}^{(t)})^{\top}\hskip-1.42262pt;\hskip-1.42262pt\cdots\hskip-1.42262pt;\hskip-1.42262pt\nabla F_{n}({x}_{n}^{(t)};\xi_{n}^{(t)})^{\top}\hskip-1.42262pt]$
$\displaystyle\nabla f({\mathbf{x}}^{(t)})$ $\displaystyle\hskip-1.42262pt:=\hskip-1.42262pt[\nabla f_{1}({x}^{(t)}_{1})^{\top};\cdots;\nabla f_{n}({x}^{(t)}_{n})^{\top}$]
where subscripts index nodes and superscripts index iterations. Let ${\mathcal{F}}_{t}$ be the filtration containing all information up to ${\mathbf{x}}^{(t)}$, excluding the stochastic gradients at iteration $t$. We use $\mathbb{E}_{t}[\cdot]=\mathbb{E}[\cdot\mid{\mathcal{F}}_{t}]$ for conditional expectation. Norms are denoted as follows: $\|\cdot\|$ for the vector $\ell_{2}$-norm, $\|\cdot\|_{F}$ for the Frobenius norm, and $\|\cdot\|_{2}:=\max_{\|v\|=1}\|Av\|$ for the matrix spectral norm.
Push-Pull algorithm. The Push-Pull algorithm, initially proposed for deterministic decentralized optimization [22, 23], can be naturally extended to the stochastic gradient setting [25]:
$\displaystyle{\mathbf{x}}^{(t+1)}$ $\displaystyle=A{\mathbf{x}}^{(t)}-\alpha{\mathbf{y}}^{(t)},$ (2a)
$\displaystyle{\mathbf{y}}^{(t+1)}$ $\displaystyle=B{\mathbf{y}}^{(t)}+{\mathbf{g}}^{(t+1)}-{\mathbf{g}}^{(t)},$ (2b)
where $A$, $B$ are row-stochastic and column-stochastic matrices determined by the underlying digraph. For initialization, we let $x_{1}^{(0)}=x_{2}^{(0)}=\cdots=x_{n}^{(0)}=x^{(0)}$ and ${\mathbf{y}}^{(0)}={\mathbf{g}}^{(0)}$. The row-stochastic mixing matrix $A$ in (2a) facilitates consensus among the local variables maintained by each node, whereas the column-stochastic mixing matrix $B$ in (2b) enables the computation of a globally averaged gradient. In the absence of gradient noise, i.e., when ${\mathbf{g}}^{(t)}=\nabla f({\mathbf{x}}^{(t)})$, it has been shown in [22, 23] that ${\mathbf{x}}^{(t+1)}$ converges to ${\mathbf{x}}^{\star}=[(x^{\star})^{\top};\cdots;(x^{\star})^{\top}]\in\mathbb{R}^{n\times d}$ for strongly convex problems, where $x^{\star}$ is the solution to problem (1).
Assumptions. We first specify the underlying graphs ($\mathcal{G}_{A}$, $\mathcal{G}_{B}$) and the corresponding weight matrices ($A$, $B$) used in the stochastic Push-Pull method.
Assumption 1 (Communication Graph).
Assume ${\mathcal{G}}_{A}$ and ${\mathcal{G}}_{B}$ to be strongly connected digraphs (see the definition in notations) with $n$ nodes.
Assumption 2 (Weight Matrix).
The weight matrix $A$ is row-stochastic with underlying digraph $\mathcal{G}_{A}$, and the weight matrix $B$ is column-stochastic with underlying digraph $\mathcal{G}_{B}$. Both $A$ and $B$ have positive trace.
Assumptions 1 and 2 are standard in the analysis of the Push-Pull method [24, 25, 27]. In our paper, their primary role is to facilitate the derivation of Proposition 1. Strictly speaking, these assumptions could be replaced as long as Proposition 1 remains valid, without affecting the main convergence results.
Proposition 1 (Exponential Decay).
For $A$ and $B$ satisfying Assumptions 1 and 2, there exist entrywise positive Perron vectors $\pi_{A},\pi_{B}\in\mathbb{R}^{n}$ such that
$\pi_{A}^{\top}A=\pi_{A}^{\top},\quad B\pi_{B}=\pi_{B},\quad\|\pi_{A}\|_{1}=\|\pi_{B}\|_{1}=1,\quad\pi_{A}^{\top}\pi_{B}>0.$
Besides, there exist constants $p_{A},p_{B}>0$, $\lambda_{A},\lambda_{B}\in[0,1)$ such that
$\|A^{t}-A_{\infty}\|_{2}\leq p_{A}\lambda_{A}^{t},\quad\|B^{t}-B_{\infty}\|_{2}\leq p_{B}\lambda_{B}^{t},\quad\forall k\geq 0.$ (3)
Proof.
The proof follows directly from the Perron-Frobenius theorem [41]. ∎
We next state assumptions regarding stochastic gradients and function smoothness.
Assumption 3 (Gradient Oracle).
There exists a constant $\sigma\geq 0$ such that
$\displaystyle\mathbb{E}_{t}[\nabla F_{i}(x_{i}^{(t)};\xi_{i}^{(t)})]$ $\displaystyle=\nabla f_{i}(x_{i}^{(t)})\quad\mbox{and}\quad\mathbb{E}_{t}[\|\nabla F_{i}(x_{i}^{(t)};\xi_{i}^{(t)})-\nabla f_{i}(x_{i}^{(t)})\|_{F}^{2}]\leq\sigma^{2}$
for any $i\in[n]$ and $t\geq 0$. In addition, we assume that stochastic gradients at different nodes are independent for any $t\geq 0$ and $1\leq i<j\leq n$.
Assumption 4 (smoothness).
There exist positive constants $L,\Delta$ such that
$\displaystyle\|\nabla f_{i}(x)-\nabla f_{i}(y)\|\leq L\|x-y\|\quad\mbox{and}\quad f_{i}(x)-\inf_{x}f_{i}(x)\leq\Delta,\quad\forall i\in[n].$
Metrics. Various effective metrics have been proposed in literautre to capture the impact of digraphs on decentralized stochastic optimization. Let $\Pi_{A}:=\mathrm{Diag}(\pi_{A})$ and $\Pi_{B}:=\mathrm{Diag}(\pi_{B})$ where $\pi_{A}$ and $\pi_{B}$ are defined in Proposition 1. Given a column-stochastic mixing matrix $B$, references [25, 38] introduce $\beta_{B}:=\|\Pi_{B}^{-1/2}(B-B_{\infty})\Pi_{B}^{1/2}\|_{2}\in[0,1)$ as a measure of graph connectivity, and define $\kappa_{B}:=\max\pi_{B}/\min\pi_{B}$ to reflect the deviation of the equilibrium distribution $\pi_{B}$ from the uniform distribution $n^{-1}\mathds{1}$. Similarly, reference [38] introduces $\beta_{A}:=\|\Pi_{A}^{1/2}(A-A_{\infty})\Pi_{A}^{-1/2}\|_{2}\in[0,1)$ and $\kappa_{A}:=\max\pi_{A}/\min\pi_{A}$ to characterize the influence of the row-stochastic mixing matrix $A$ on decentralized algorithms.
This paper introduces a new set of metrics to facilitate the convergence analysis of the Push-Pull method. When Proposition 1 holds, we define:
$\displaystyle s_{A}$ $\displaystyle:=\max\big\{\sum_{i=0}^{\infty}\|A^{i}-A_{\infty}\|_{2},\sum_{i=0}^{\infty}\|A^{i}-A_{\infty}\|_{2}^{2}\big\},$ (4)
$\displaystyle s_{B}$ $\displaystyle:=\max\big\{\sum_{i=0}^{\infty}\|B^{i}-B_{\infty}\|_{2},\sum_{i=0}^{\infty}\|B^{i}-B_{\infty}\|_{2}^{2}\big\},\vskip 5.69054pt$ (5)
$\displaystyle c$ $\displaystyle:=n\,\pi_{A}^{\top}\pi_{B}.$ (6)
By the exponential decay property (3), both $s_{A}$ and $s_{B}$ are finite. In this paper, the influence of the mixing matrices is described solely through $s_{A}$, $s_{B}$, $c$, and $n$. Their connection to previously introduced metrics is given in the next proposition.
Proposition 2 (Relation with existing metrics).
Suppose Assumption 1 and 2 hold. The following inequalities on $s_{A}$, $s_{B}$ and $c$ hold:
$\displaystyle 1\leq s_{A}\leq\frac{M_{A}^{2}(1+\frac{1}{2}\ln(\kappa_{A}))}{1-\beta_{A}},$ $\displaystyle\quad 1\leq s_{B}\leq\frac{M_{B}^{2}(1+\frac{1}{2}\ln(\kappa_{B}))}{1-\beta_{B}},$
$\displaystyle n\max\{\min\pi_{A},\min\pi_{B}\}$ $\displaystyle\leq\,c\leq n\min\{\max\pi_{A},\max\pi_{B}\}.$
where $M_{A}:=\max_{t\geq 0}\|A^{t}-A_{\infty}\|_{2}$ and $M_{B}:=\max_{t\geq 0}\|B^{t}-B_{\infty}\|_{2}$ are finite numbers which can be proved to be no larger than $\sqrt{n}$.
Proof.
See Appendix A. ∎
3 Intrinsic Differences between Push-Only, Pull-Only, and Push-Pull
This section provides a concise overview of the linear speedup analysis for the Push-Only and Pull-Only algorithms, clarifying their key differences from Push-Pull. We also highlight the limitations of existing analyses and explain why they cannot establish the linear speedup property of Push-Pull.
3.1 Push-Only and Pull-Only align with centralized SGD
Many decentralized algorithms—such as decentralized SGD [36], gradient tracking [32, 33, 34], Push-Diging [9], Gradient-Pull [11], and FROST [19]—are designed to emulate centralized SGD in a decentralized setting. These algorithms can be reformulated as
$\displaystyle\hat{x}^{(t+1)}=\underbrace{\hat{x}^{(t)}-\alpha\bar{g}^{(t)}}_{\text{centralized SGD}}+\epsilon^{(t)},$ (7)
where $\hat{x}^{(t)}=\sum_{i=1}^{n}v_{i}^{(t)}(x_{i}^{(t)})^{\top}\in\mathbb{R}^{1\times d}$ is a convex combination of the local decision variables $\{x_{i}^{(t)}\}_{i=1}^{n}$, with $v_{i}^{(t)}$ denoting the combination weights. The term $\bar{g}^{(t)}=\tfrac{1}{n}\mathds{1}^{\top}{\mathbf{g}}^{(t)}$ is the globally averaged stochastic gradient, and $\epsilon^{(t)}$ captures the residual error due to decentralization. When $\epsilon^{(t)}=0$, the recursion in (7) reduces to centralized SGD, whose linear speedup property is well established (see Table 1).
For Push-Only and Pull-Only algorithms, the residual $\epsilon^{(t)}$ is either exactly zero or decays exponentially to zero. This ensures that the trajectory of the weighted average of local iterates closely tracks that of centralized SGD. Once the residual is sufficiently small, the linear speedup property follows directly from the standard analysis of centralized SGD. We illustrate this alignment with two examples.
Example 1 (Push-DiGing [9]).
Push-DiGing is a Push-Only method, whose iterates are given by
$\displaystyle{\mathbf{x}}^{(t+1)}$ $\displaystyle=A_{t}{\mathbf{x}}^{(t)}-\alpha V_{t+1}^{-1}{\mathbf{y}}^{(t)},$ (8a)
$\displaystyle{\mathbf{y}}^{(t+1)}$ $\displaystyle=B{\mathbf{y}}^{(t)}+{\mathbf{g}}^{(t+1)}-{\mathbf{g}}^{(t)},$ (8b)
where $B$ is column-stochastic, $v_{t}=B^{t}\mathds{1}$, $V_{t}=\mathrm{Diag}(v_{t})$, $A_{t}=V_{t+1}^{-1}BV_{t}$, and ${\mathbf{g}}^{(t)}$ stacks local stochastic gradients. Left-multiplying (8a) by $n^{-1}v_{t+1}^{\top}$ yields
$n^{-1}v_{t+1}^{\top}{\mathbf{x}}^{(t+1)}=n^{-1}v_{t}^{\top}{\mathbf{x}}^{(t)}-\alpha\bar{g}^{(t)}.$
Defining $\hat{x}^{(t)}=n^{-1}v_{t}^{\top}{\mathbf{x}}^{(t)}$, the recursion matches (7) with $\epsilon^{(t)}=0$.
Example 2 (FROST [18, 19]).
FROST/PullDiag-GT is a Pull-Only method based on a row-stochastic matrix $A$. Its updates are
$\displaystyle{\mathbf{x}}^{(t+1)}$ $\displaystyle=A{\mathbf{x}}^{(t)}-\alpha{\mathbf{y}}^{(t)},$ (9a)
$\displaystyle{\mathbf{y}}^{(t+1)}$ $\displaystyle=A{\mathbf{y}}^{(t)}+D_{t+1}^{-1}{\mathbf{g}}^{(t+1)}-D_{t}^{-1}{\mathbf{g}}^{(t)},$ (9b)
where $D_{t}=\mathrm{Diag}(A^{t})$, ${\mathbf{g}}^{(t)}$ is the stacked stochastic gradient, and ${\mathbf{y}}^{(0)}={\mathbf{g}}^{(0)}$. Left-multiplying (9a) by $n^{-1}\pi_{A}^{\top}$ gives
$n^{-1}\pi_{A}^{\top}{\mathbf{x}}^{(t+1)}=n^{-1}\pi_{A}^{\top}{\mathbf{x}}^{(t)}-\alpha\bar{g}^{(t)}+\epsilon^{(t)},$
where $\epsilon^{(t)}=\alpha n^{-1}(\mathds{1}^{\top}-\pi_{A}^{\top}D_{t}^{-1}){\mathbf{g}}^{(t)}$ decays exponentially fast [39, Lemma 12]. Setting $\hat{x}^{(t)}=n^{-1}\pi_{A}^{\top}{\mathbf{x}}^{(t)}$ recovers the form (7).
3.2 Push–Pull does not explicitly align with centralized SGD
Unlike Push-Only and Pull-Only, the Push-Pull method cannot be perfectly aligned with centralized SGD. Although it can be written in the form of (7), it suffers from a non-vanishing error term. Recall the Push-Pull recursion (2). Left-multiplying (2b) by $\mathds{1}^{\top}$ yields
$\displaystyle\mathds{1}^{\top}{\mathbf{y}}^{(t+1)}=\mathds{1}^{\top}{\mathbf{y}}^{(t)}+\mathds{1}^{\top}{\mathbf{g}}^{(t+1)}-\mathds{1}^{\top}{\mathbf{g}}^{(t)}=\mathds{1}^{\top}{\mathbf{g}}^{(t+1)},$ (10)
where the last equality follows from the initialization ${\mathbf{y}}^{(0)}={\mathbf{g}}^{(0)}$. Left-multiplying step (2a) by $\pi_{A}^{\top}$, we obtain
$\displaystyle\pi_{A}^{\top}{\mathbf{x}}^{(t+1)}$ $\displaystyle=\pi_{A}^{\top}A{\mathbf{x}}^{(t)}-\alpha\pi_{A}^{\top}B_{\infty}{\mathbf{y}}^{(t)}-\alpha\pi_{A}^{\top}(I-B_{\infty}){\mathbf{y}}^{(t)}$
$\displaystyle=\pi_{A}^{\top}{\mathbf{x}}^{(t)}-\alpha\pi_{A}^{\top}\pi_{B}\mathds{1}^{\top}{\mathbf{y}}^{(t)}-\alpha\pi_{A}^{\top}(I-B_{\infty}){\mathbf{y}}^{(t)}.$
Letting $\hat{x}^{(t)}=\pi_{A}^{\top}{\mathbf{x}}^{(t)}$ and using (10), we obtain
$\displaystyle\hat{x}^{(t+1)}=\hat{x}^{(t)}-c\alpha\bar{g}^{(t)}+\epsilon^{(t)},\quad\text{where}\quad\epsilon^{(t)}=-\alpha\pi_{A}^{\top}(I-B_{\infty}){\mathbf{y}}^{(t)},$ (11)
with $c=n\pi_{A}^{\top}\pi_{B}$ (see Definition (6)) and $\bar{g}^{(t)}=n^{-1}\mathds{1}^{\top}{\mathbf{g}}^{(t)}$. Unlike Push-DiGing (Example 1), where $\epsilon^{(t)}=0$, or FROST (Example 2), where $\epsilon^{(t)}$ decays exponentially, the error term $\pi_{A}^{\top}(I-B_{\infty}){\mathbf{y}}^{(t)}$ in Push-Pull does not vanish but instead stabilizes at a non-negligible constant. As shown in Figure 1, this persistent error remains regardless of network size or topology, making the linear speedup analysis of Push-Pull substantially more challenging.
Figure 1: With a fixed stepsize $\alpha$, the error term $\epsilon^{(t)}$ in Push-Pull does not vanish but rapidly stabilizes at a non-negligible constant. Results are shown for logistic regression on a synthetic dataset.
4 A Novel Multiple-Step Descent Analysis Framework
Establishing linear speedup for Push-Pull is challenging due to the presence of a non-vanishing error in (11). We begin by revisiting the classical single-step descent framework, highlighting why it suffices for Push-Only and Pull-Only methods but fails for Push-Pull, thereby motivating the development of a new multi-step descent analysis.
4.1 Limitations of single-step descent analysis
Following conventional single-step descent analysis, we establish the descent inequality based on update (11):
$\displaystyle\|\nabla f(\hat{x}^{(t)})\|^{2}\hskip-2.84526pt\lesssim\hskip-2.84526pt\frac{f(\hat{x}^{(t)})\hskip-2.84526pt-\hskip-2.84526pt\mathbb{E}_{t}[f(\hat{x}^{(t+1)})]}{c\alpha}\hskip-2.27621pt+\hskip-2.27621pt\frac{c\alpha L}{n}\sigma^{2}\hskip-2.27621pt+\hskip-2.27621pt\alpha L\|\pi_{A}^{\top}(I\hskip-2.84526pt-\hskip-2.84526ptB_{\infty}){\mathbf{y}}^{(t)}\|^{2}\hskip-2.84526pt+\hskip-2.84526pt\Delta_{\text{consent}}^{(t)}$ (12)
where $\Delta_{\text{consent}}^{(t)}$ denotes consensus errors among local variables, typically of order $\mathcal{O}(\alpha^{2})$ and hence negligible when $\alpha$ is sufficiently small. We use $\lesssim$ to denote inequalities that hold up to absolute numerical constants. Taking full expectation and summing over $t=0,\dots,T-1$ yields
$\displaystyle\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla f(\hat{x}^{(t)})\|^{2}]$ $\displaystyle\lesssim\frac{f(\hat{x}^{(0)})\hskip-1.42262pt-\hskip-1.42262ptf(\hat{x}^{(T)})}{c\alpha T}\hskip-1.42262pt+\hskip-1.42262pt\frac{c\alpha L}{n}\sigma^{2}\hskip-1.42262pt\hskip-1.42262pt$ (13)
$\displaystyle\quad+\frac{\alpha L}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\pi_{A}^{\top}(I\hskip-1.42262pt-\hskip-1.42262ptB_{\infty}){\mathbf{y}}^{(t)}\|^{2}]+\frac{1}{T}\sum_{t=1}^{T-1}\mathbb{E}[\Delta^{(t)}_{\rm consent}$]. (14)
Since $\tfrac{1}{T}\sum_{t=1}^{T-1}\mathbb{E}[\Delta^{(t)}_{\rm consent}$] is negligible for sufficiently small $\alpha$, the convergence rate is dominated by the two main terms in (13), provided the first term in (14) is either zero (Example 1) or decays exponentially (Example 2). With $\alpha=\sqrt{n/(c^{2}\sigma^{2}T)}$, we achieve $\tfrac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla f(\hat{x}^{(t)})\|^{2}]=\mathcal{O}(\sigma/\sqrt{nT})$, implying $\mathcal{O}(1/(n\epsilon^{2}))$ iteration complexity and hence establish the linear speedup property.
However, unlike Push-Only and Pull-Only, the term $\mathbb{E}[\|\pi_{A}^{\top}(I-B_{\infty}){\mathbf{y}}^{(t)}\|^{2}]$ in Push-Pull persists (see Figure 1). The central challenge in establishing linear speedup is therefore to prove the following inequality:
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\pi_{A}^{\top}(I-B_{\infty})\mathbf{y}^{(t)}\|^{2}]\lesssim\frac{c}{n}\sigma^{2}+\text{negligible terms}.$ (15)
If this inequality holds, the first term in (14) can be absorbed into the second term of (13). Choosing $\alpha=\sqrt{n/(c^{2}\sigma^{2}T)}$ will yield the same linear speedup rate. Unfortunately, this cannot be proved with standard techniques. Recall that $B_{\infty}=\pi_{B}\mathds{1}_{n}^{\top}$ and define $\Delta_{y}^{(t)}:=(I-B_{\infty})\mathbf{y}^{(t)}$. Left-multiplying (2b) by $(I-B_{\infty})$ gives
$\displaystyle\Delta_{y}^{(t+1)}=(B-B_{\infty})\Delta_{y}^{(t)}+(I-B_{\infty})\big({\mathbf{g}}^{(t+1)}-{\mathbf{g}}^{(t)}\big).$ (16)
where ${\mathbf{g}}^{(t)}$ stacks the $n$ stochastic gradients $\nabla F_{i}(x_{i}^{(t)};\xi_{i}^{(t)})$ (each with variance $\sigma^{2}$), we can therefore derive that $\mathbb{E}[\|\pi_{A}^{\top}\Delta_{y}^{(t)}\|^{2}]=\Theta(n\sigma^{2})$, which is far larger than the desired $c\sigma^{2}/n$. This motivates the need for a fundamentally different approach.
4.2 A novel multi-step descent analysis
In contrast to analyzing each iteration in isolation, we now introduce a novel $m$-step recursion that aggregates information across a block of $m$ iterations. This blockwise perspective reveals cancellation effects in the gradient-tracking recursion and enables the gradient noise to be controlled at the desired level. Specifically, for a positive integer $m$ and any index $k$ that is a multiple of $m$, the recursion (11) can be written as
$\hat{x}^{(k+m)}=\hat{x}^{(k)}-c\alpha\sum_{i=0}^{m-1}\bar{g}^{(k+i)}-\alpha\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)},$ (17)
where $\Delta_{y}^{(t)}=(I-B_{\infty})\mathbf{y}^{(t)}$. Analogous to the derivation of the single-step descent inequality (12), we can establish the following $m$-step descent inequality:
$\displaystyle\|\nabla f(\hat{x}^{(k)})\|^{2}\hskip-2.27621pt\lesssim\hskip-2.27621pt\frac{f(\hat{x}^{(k)})\hskip-2.27621pt-\hskip-2.27621pt\mathbb{E}_{k}[f(\hat{x}^{(k+m)})]}{c\alpha m}\hskip-2.27621pt+\hskip-2.27621pt\frac{c\alpha L}{n}\sigma^{2}\hskip-2.27621pt+\hskip-2.27621pt\frac{\alpha L}{m}\|\pi_{A}^{\top}\hskip-2.27621pt\sum_{i=0}^{m-1}\hskip-2.84526pt\Delta_{y}^{(k+i)}\hskip-0.56905pt\|^{2}\hskip-2.27621pt+\hskip-2.27621pt\Delta_{\text{consent}}^{(k,m)}$ (18)
where $\Delta_{\text{consent}}^{(k,m)}$ denotes consensus errors within the $m$-step block, which are typically $\mathcal{O}(\alpha^{2})$ and negligible for small $\alpha$. A formal statement of (18) is provided in Lemma 3. Let $T=mK$ denote the total number of iterations and define $[m^{*}]:=\{0,m,2m,\ldots,(K-1)m\}$. Taking the full expectation of (18) and summing over all $k\in[m^{*}]$ yields
$\displaystyle\quad\min_{0\leq t\leq T-1}\mathbb{E}[\|\nabla f(\hat{x}^{(t)})\|^{2}]$ $\displaystyle\leq\frac{1}{K}\sum_{k\in[m^{*}]}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]\lesssim\frac{f(\hat{x}^{(0)})-f(\hat{x}^{(T)})}{c\alpha T}+\frac{c\alpha L}{n}\sigma^{2}$
$\displaystyle\hskip-22.76228pt+\frac{\alpha L}{T}\sum_{k\in[m^{*}]}\mathbb{E}[\|\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}\|^{2}]+\frac{\alpha L}{T}\sum_{k\in[m^{*}]}\mathbb{E}[\Delta_{\text{consent}}^{(k,m)}].$ (19)
Similar to the single-step descent inequality (14), the central component in the $m$-step analysis is to control $\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}$. Summing the single-step recursion (16) from $i=0$ to $m-1$ yields
$\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}=\Delta_{y}^{(k)}+(B-B_{\infty})\sum_{i=0}^{m-2}\Delta_{y}^{(k+i)}+(I-B_{\infty})({\mathbf{g}}^{(k+m-1)}-{\mathbf{g}}^{(k)}).$
This telescoping structure cancels all intermediate stochastic gradients ${\mathbf{g}}^{(k+i)}$ for $1\leq i\leq m-2$, leaving only boundary terms $\mathbf{g}^{(k)}$ and $\mathbf{g}^{(k+m-1)}$. Repeatedly applying this argument yields the equality
$\displaystyle\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}=\Big(\sum_{j=0}^{m-1}(B^{j}-B_{\infty})\Big)\Delta_{y}^{(k)}+\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})({\mathbf{g}}^{(k+j)}-{\mathbf{g}}^{(k)}).$ (20)
Unlike the single-step recursion (15), the $m$-step formulation cancels all interior noise, leaving only boundary differences ${\mathbf{g}}^{(k+j)}-{\mathbf{g}}^{(k)}$, each attenuated by exponentially decaying factors $B^{m-j-1}-B_{\infty}$ (see Proposition 2). As a result, the block-level noise remains bounded at a constant level, independent of $m$.
In particular, each boundary difference contributes gradient noise with variance $n\sigma^{2}$, but its effect is scaled by $\lambda_{B}^{m-j-1}$, yielding $\sum_{j=1}^{m-1}\lambda_{B}^{m-j-1}n\sigma^{2}\leq(n\sigma^{2})/(1-\lambda_{B})$ which is constant in $m$. Therefore,
$\displaystyle\ \frac{1}{T}\sum_{k\in[m^{*}]}\mathbb{E}[\|\pi_{A}^{\top}\Big(\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}\Big)\|^{2}]$ (21)
$\displaystyle\lesssim$ $\displaystyle\ \frac{n\sigma^{2}}{m}+\text{negligible terms}\lesssim\frac{\sigma^{2}}{n}+\text{negligible terms}.$
The second inequality holds when $m$ is chosen sufficiently large [^1] In rigorous proofs, the choice of $m$ is constrained by additional factors and is therefore more involved than the simplified condition stated here, see more details in Section 5., e.g., $m\geq\text{Const}\cdot n^{2}$. Recalling that $T=mK$, substituting (21) into (18) and choosing $\alpha=\sqrt{n/(c^{2}\sigma^{2}T)}$ yield the desired linear speedup rate. This highlights the unique advantage of the multiple-step descent analysis: by aggregating errors across $m$ steps within a block, it reveals a telescoping cancellation in the gradient-tracking recursion, leaving only boundary terms that are exponentially decayed during iterations. Consequently, the effective gradient noise is controlled at the desired $\mathcal{O}(\sigma^{2}/n)$ level, enabling the first proof of linear speedup for Push-Pull on strongly connected digraphs.
5 Achieving Linear Speedup for Push-Pull
In this section, we formalize the aforementioned arguments through supporting lemmas and rigorously establish the linear speedup property for Push-Pull.
5.1 Supporting lemmas
Lemma 3 (m-step Descent Lemma).
Suppose Assumption 34 and Proposition 1 hold. For any integers $k\geq 0$, we define $\hat{x}^{(k)}=\pi_{A}^{\top}{\mathbf{x}}^{(k)}$. When $\alpha\leq\frac{1}{6cmL}$, the following inequality holds for integers $k\geq 0,m\geq 6c^{-2}s_{B}^{2}$ (Proof is in Appendix B.1):
$\displaystyle\quad\|\nabla f(\hat{x}^{(k)})\|^{2}+\frac{1}{m}\sum_{i=0}^{m-1}\mathbb{E}_{k}[\|\overline{\nabla f}^{(k+i)}\|^{2}]$ (22)
$\displaystyle\leq\frac{4(f(\hat{x}^{(k)})-\mathbb{E}_{k}[f(\hat{x}^{(k+m)})])}{\alpha cm}+\frac{8\alpha cL}{n}\sigma^{2}+\frac{8\alpha Lns_{B}^{2}}{cm}\sigma^{2}$ (23)
$\displaystyle\quad+\frac{2L^{2}}{m}\Delta_{1}^{(k)}+\frac{2L^{2}}{m}\Delta_{2}^{(k)}+\frac{12s_{B}^{2}}{c^{2}m^{2}}\|\Delta_{y}^{(k)}\|_{F}^{2},$ (24)
where $\overline{\nabla f}^{(k+i)}:=n^{-1}\mathbf{1}^{\top}\nabla f({\mathbf{x}}^{(k+i)})$ denotes the average of the local deterministic gradients evaluated at ${\mathbf{x}}^{(k+i)}$, and
$\displaystyle\Delta_{1}^{(k)}:=\sum_{j=0}^{m-1}\mathbb{E}_{k}[\|{\mathbf{x}}^{(k+j)}-\hat{{\mathbf{x}}}^{(k)}\|_{F}^{2}],\quad\Delta_{2}^{(k)}:=\sum_{j=1}^{m}\mathbb{E}_{k}[\|{\mathbf{x}}^{(k+j)}-{\mathbf{x}}^{(k)}\|_{F}^{2}].$
In (23), when $m\geq c^{-2}n^{2}s_{B}^{2}$, the noise terms associated with $\sigma^{2}$ are dominated solely by the linear speedup term $\tfrac{8\alpha cL}{n}\sigma^{2}$, which is consistent with our analysis in (21). The remaining error terms appearing in (24) are estimated in the next lemma.
Lemma 4.
Suppose Assumption 34 and Proposition 1 hold. With $\Delta_{1}^{(k)}$, $\Delta_{2}^{(k)}$ defined above, when $\alpha\leq\min\{\frac{1}{10cm\sqrt{s_{B}s_{B^{m}}^{2}}L},\frac{1}{6L}\}$ and $m\geq C_{m}$, we have
$\displaystyle\quad\frac{2L^{2}}{m}\sum_{k\in[m^{*}]}\mathbb{E}[\Delta_{1}^{(k)}+\Delta_{2}^{(k)}]+\frac{12s_{B}^{2}}{c^{2}m^{2}}\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]$
$\displaystyle\leq\frac{1}{m}\sum_{t=0}^{T-1}\mathbb{E}_{k}[\|\overline{\nabla f}^{(t)}\|^{2}]+(\frac{73c\alpha L}{n}+\frac{384s_{B}^{3}s_{B^{m}}^{2}}{m^{2}c^{2}})\sigma^{2}+C_{\Delta,1}\frac{L\Delta}{4m^{2}}.$
Here, $C_{m}$ and $C_{\Delta,1}$ denote matrix-related constants (see (B.2)). The quantity $s_{B^{m}}$ represents the matrix measure of $B^{m}$, as defined in (5). In general, we have $s_{B^{m}}\leq s_{B}$, and moreover, $s_{B^{m}}$ converges to $\lVert I-B_{\infty}\rVert_{2}^{2}$ as $m\to\infty$ (Proof is in Appendix B.2).
With the above lemmas, we are ready to state the main result of this paper.
Theorem 1.
Suppose Assumption 34 and Proposition 1 hold. For given positive integers $m$ and $K$, we define $T:=mK$. When $\alpha\leq\frac{1}{10cm\sqrt{s_{B}s_{B^{m}}^{2}}L}$ and $m\geq C_{m}$, we have
$\displaystyle\quad\frac{1}{K}\sum_{k\in[m^{*}]}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]\leq\frac{4\Delta}{c\alpha T}+\frac{100c\alpha L}{n}\sigma^{2}+\frac{384s_{B}^{3}s_{B^{m}}^{2}}{m^{2}c^{2}}\sigma^{2}+C_{\Delta,1}\frac{L\Delta}{mT}.$ (25)
where $C_{m}$ and $C_{\Delta,1}$ are matrix-related constants defined in (B.2) and (53). Furthermore, if we choose
$\displaystyle m$ $\displaystyle=\lceil\max\{4c^{-1}s_{B}^{1.5}s_{B^{m}}\left(\frac{nT\sigma^{2}}{L\Delta}\right)^{0.25},C_{m}\}\rceil,$
$\displaystyle\alpha$ $\displaystyle=\min\left\{\frac{\sqrt{n\Delta}\sigma}{10c\sqrt{LT}},\frac{1}{10cm\sqrt{s_{B}s_{B^{m}}^{2}}L},\frac{1}{6L}\right\},$
it holds that
$\displaystyle\min_{t\in\{0,1,\ldots,T-1\}}\mathbb{E}[\|\nabla f(\hat{x}^{(t)}\|^{2}]$ $\displaystyle\leq\frac{44\sigma\sqrt{L\Delta}}{\sqrt{nT}}+\frac{L\Delta(m^{-1}C_{\Delta,1}+10m\sqrt{s_{B}s_{B^{m}}^{2}}+c^{-1})}{T}$ (26)
$\displaystyle=\frac{44\sigma\sqrt{L\Delta}}{\sqrt{nT}}+{\mathcal{O}}\left(\frac{L\Delta}{T}\right)^{\frac{3}{4}}+{\mathcal{O}}\left(\frac{L\Delta}{T}\right).$
Proof.
See Appendix B.3. ∎
Comparison with existing analyses. Theorem 1 provides the first convergence guarantee for stochastic Push-Pull on nonconvex problems. In contrast, prior analyses focus either on the strongly convex setting [42, 25] or on Push-Pull variants restricted to specific network topologies [7]. The leading term $\tfrac{44\sigma\sqrt{L\Delta}}{\sqrt{nT}}$ demonstrates a linear speedup with respect to the number of nodes $n$ for any topology satisfying Proposition 2. Consequently, Push-Pull requires $\mathcal{O}(1/(n\epsilon^{2}))$ iterations to achieve an $\epsilon$-accurate solution, with the complexity decreasing linearly as $n$ increases.
Impact of $m$: In our analysis, we consider $m$-step gradient descent, where the parameter $m$ governs the noise-related terms, as reflected by the $\mathcal{O}(m^{-2}\sigma^{2})$ term in (25). However, $m$ cannot be taken arbitrarily large: increasing $m$ reduces the admissible range of the learning rate, thereby slowing convergence in the analysis. Overall, $m$ balances noise reduction against convergence rate. With an appropriate choice of $m$, one can suppress noise, achieve linear speedup, and still guarantee $\mathcal{O}(1/\sqrt{nT})$.
6 Experiment
In this section, we empirically examine the linear speedup properties of the Push-Pull algorithm across a variety of network topologies. Each node in the stochastic gradient oracle operates on a local dataset, computing gradients from randomly sampled data points at every iteration. The experimental study consists of two phases: (i) nonconvex regression on synthetic data, and (ii) training a fully connected neural network on the heterogeneously distributed MNIST dataset.
To further evaluate practical performance on standard benchmarks, we extend our experiments to training ResNet-18 on CIFAR-10. The distributed training is performed across multiple nodes ($n>1$) under different network topologies, including directed exponential graphs, grid graphs, and random graphs. For reference, all results are systematically compared against the single-node baseline ($n=1$) to quantify the speedup obtained through distributed optimization.
Complete implementation details are provided in Appendix C, and the source code with experimental scripts is publicly available at our GitHub repository.
6.1 Nonconvex Logistic Regression on Synthetic Dataset
In the first experiment, the loss function at each node is logistic regression with nonconvex regularization terms. We evaluate six network topologies: exponential, grid, and ring graphs (directed), and random, geometric, and nearest-neighbor graphs (undirected). The mixing matrices $A$ and $B$ are constructed using a specialized procedure, which yields row-stochastic or column-stochastic matrices, but not doubly stochastic ones. Details are provided in Appendix C.1 and Appendix C.2.
As shown in Figure 2, the results exhibit clear linear speedup. Across all topologies, the gradient noise level decreases proportionally to the square root of the number of nodes, given the same number of communication rounds. The learning rate is fixed at $lr=0.005$ for all experiments.
Figure 2: Linear speedup performance over various topologies on synthetic data, the learning rates are all $0.005$.
6.2 Training Neural Network for MNIST Classification
In the second experiment, we trained a three-layer fully connected neural network for MNIST classification. The same network topologies as in the synthetic dataset experiment were used. The MNIST data are distributed heterogeneously according to a Dirichlet distribution (see Appendix C.3). As shown in Figure 3, Push-Pull achieves linear speedup on both directed and undirected topologies.
Figure 3: Linear speedup performance over various topologies, the learning rates are the same over the same topology.
6.3 Training Neural Network for CIFAR10 Image Classification
In the third experiment, we trained ResNet-18 for CIFAR-10 classification across varying numbers of nodes and network topologies. We considered three representative topologies: directed exponential graphs, grid graphs, and undirected random graphs. For reference, $n=1$ corresponds to training ResNet-18 on a single node without distributed optimization.
Figure 4 illustrates the performance of the Push-Pull method. We observe that its accuracy is initially lower than that of single-node training, due to the overhead introduced by network topology. However, after a sufficient number of epochs, Push-Pull surpasses the single-node baseline, benefiting from its inherent speedup advantages.
Figure 4: The test accuracy across different numbers of nodes and various topologies. All experiments were conducted using a learning rate of $0.005$. The lower three figures specifically depict the accuracy after $100$ epochs.
7 Conclusion
In this paper, we introduce a novel analysis technique and demonstrate linear speedup convergence for the Push-Pull method across arbitrary strongly connected topologies. This discovery enhances the community’s comprehension of the Push-Pull method and validates its effectiveness. Numerical experiments have corroborated our theoretical findings.
Appendix A Linear Algebra Inequalities
Lemma 5 (Rolling Sum Lemma).
For a row-stochastic weight matrix satisfying Proposition 1, the following inequality holds for any $T\geq 0$:
$\displaystyle\sum_{k=0}^{T}\|\sum_{i=0}^{k}(A^{k-i}-A_{\infty})\Delta^{(i)}\|_{F}^{2}\leq s_{A}^{2}\sum_{i=0}^{T}\|\Delta^{(i)}\|_{F}^{2},$ (27)
where $\Delta^{(i)}\in\mathbb{R}^{n\times d}$ are arbitrary matrices, $s_{A}$ is defined as in (4). Furthermore, if Assumption 1 and 2 hold, we have
$\displaystyle s_{A}:=\max\{\sum_{i=0}^{\infty}\|A^{i}-A_{\infty}\|_{2},\sum_{i=0}^{\infty}\|A^{i}-A_{\infty}\|_{2}^{2}\}\leq\frac{M_{A}^{2}(1+\frac{1}{2}\ln(\kappa_{A}))}{1-\beta_{A}}.$
where $\beta_{A}$, $\kappa_{A},M_{A}$ are defined in Proposition 2. The row-stochastic matrix $A$ and the subscript $A$ can be replaced by a column-stochastic matrix $B$ without changing the lemma.
Proof.
First, we prove that
$\displaystyle\|A^{i}-A_{\infty}\|_{2}\leq\sqrt{\kappa_{A}}\beta_{A}^{i},\forall i\geq 0.$ (28)
Notice that $\beta_{A}:=\|A-A_{\infty}\|_{\pi_{A}}$ and
$\|A^{i}-A_{\infty}\|_{\pi_{A}}=\|(A-A_{\infty})^{i}\|_{\pi_{A}}\leq\|A-A_{\infty}\|^{i}_{\pi_{A}}=\beta_{A}^{i},$
we have
$\|(A^{i}-A_{\infty})v\|=\|\Pi_{A}^{-1/2}(A^{i}-A_{\infty})v\|_{\pi_{A}}\leq\sqrt{\underline{\pi_{A}}}\beta_{A}^{i}\|v\|_{\pi_{A}}\leq\sqrt{\kappa_{A}}\beta_{A}^{i}\|v\|,$
where $\underline{\pi_{A}}:=\min\pi_{A}$, $\overline{\pi_{A}}:=\max\pi_{A}$. The last inequality comes from $\|v\|_{\pi_{A}}\leq\overline{\pi_{A}}\|v\|$. Therefore, (28) holds.
Second, we want to prove that
$\displaystyle s_{A}\leq\frac{M_{A}^{2}(1+\frac{1}{2}\ln(\kappa_{A}))}{1-\beta_{A}}.$ (29)
Towards this end, we define $M_{A}:=\max\{1,\max_{k\geq 0}\|A^{k}-A_{\infty}\|_{2}\}$. According to (28), $M_{A}$ is well-defined. We also define $p=\max\left\{\frac{\ln(\sqrt{\kappa_{A}})-\ln(M_{A})}{-\ln(\beta_{A})},0\right\}$, then we can verify that $\|A^{i}-A_{\infty}\|_{2}\leq\min\{M_{A},M_{A}\beta_{A}^{i-p}\},\forall i\geq 0.$ With this inequality, we can bound $\sum_{i=0}^{k}\|A^{k+1-i}-A_{\infty}\|_{2}$ as follows:
$\displaystyle\sum_{i=0}^{k}\|A^{k-i}-A_{\infty}\|_{2}=\sum_{i=0}^{\min\{\lfloor p\rfloor,k\}}\|A^{i}-A_{\infty}\|_{2}+\sum_{i=\min\{\lfloor p\rfloor,k\}+1}^{k}\|A^{i}-A_{\infty}\|_{2}$ (30)
$\displaystyle\leq$ $\displaystyle\sum_{i=0}^{\min\{\lfloor p\rfloor,k\}}M_{A}+\sum_{i=\min\{\lfloor p\rfloor,k\}+1}^{k}M_{A}\beta_{A}^{i-p}$
$\displaystyle\leq$ $\displaystyle M_{A}\cdot(1+\min\{\lfloor p\rfloor,k\})+M_{A}\cdot\frac{1}{1-\beta_{A}}\beta_{A}^{\min\{\lfloor p\rfloor,k\}+1-p}.$
If $p=0$, (30) is simplified to $\sum_{i=0}^{k}\|A^{k-i}-A_{\infty}\|_{2}\leq M_{A}\cdot\frac{1}{1-\beta_{A}}$ and (29) is naturally satisfied. If $p>0$, let $x=\min\{\lfloor p\rfloor,k\}+1-p\in[0,1)$, (29) is simplified to
$\sum_{i=0}^{k}\|A^{k-i}-A_{\infty}\|_{2}\leq M_{A}(x+p+\frac{\beta_{A}^{x}}{1-\beta_{A}})\leq M_{A}(p+\frac{1}{1-\beta_{A}})=\frac{M_{A}(1+\frac{1}{2}\ln(\kappa_{A}))}{1-\beta_{A}}.$
By the definition of $p$, we have $p\leq\frac{\frac{1}{2}\ln(\kappa_{A})}{1-\beta_{A}}$. This completes the estimation of $\sum_{i=0}^{\infty}\|A^{k-i}-A_{\infty}\|_{2}$. For the summation of the squared norm, we can use the same $p$ and obtain the upper bound without modifying the previous steps:
$\sum_{i=0}^{k}\|A^{k-i}-A_{\infty}\|_{2}^{2}\leq M_{A}^{2}(p+\frac{1}{1-\beta_{A}})=\frac{M_{A}^{2}(1+\frac{1}{2}\ln(\kappa_{A}))}{1-\beta_{A}}.$
This finishes the proof of (29).
Finally, to obtain (27), we use Jensen’s inequality. For any positive numbers $a_{0},a_{1},\ldots,a_{k}$ satisfying $\sum_{i=0}^{k}a_{i}=1$, we have
$\displaystyle\|\sum_{i=0}^{k}(A^{k-i}-A_{\infty})\Delta^{(i)}\|_{F}^{2}=\|\sum_{i=0}^{k}a_{k-i}\cdot a_{k-i}^{-1}(A^{k-i}-A_{\infty})\Delta^{(i)}\|_{F}^{2}$ (31)
$\displaystyle\leq$ $\displaystyle\sum_{i=0}^{k}a_{k-i}\|a_{k-i}^{-1}(A^{k-i}-A_{\infty})\Delta^{(i)}\|_{F}^{2}\leq\sum_{i=0}^{k}a_{k-i}^{-1}\|A^{k-i}-A_{\infty}\|_{2}^{2}\|\Delta^{(i)}\|_{F}^{2}.$
By choosing $a_{k-i}=(\sum_{i=0}^{k}\|A^{k-i}-A_{\infty}\|_{2})^{-1}\|A^{k-i}-A_{\infty}\|_{2}$ in (31), we obtain that
$\displaystyle\|\sum_{i=0}^{k}(A^{k-i}-A_{\infty})\Delta^{(i)}\|_{F}^{2}\leq\sum_{i=0}^{k}\|A^{k-i}-A_{\infty}\|_{2}\cdot\sum_{i=0}^{k}\|A^{k-i}-A_{\infty}\|_{2}\|\Delta^{(i)}\|_{F}^{2}.$ (32)
By summing up (32) from $k=0$ to $T$, we obtain that
$\displaystyle\sum_{k=0}^{T}\|\sum_{i=0}^{k}(A^{k-i}-A_{\infty})\Delta^{(i)}\|_{F}^{2}\leq s_{A}\sum_{k=0}^{T}\sum_{i=0}^{k}\|A^{k-i}-A_{\infty}\|_{2}\|\Delta^{(i)}\|_{F}^{2}$
$\displaystyle\leq$ $\displaystyle s_{A}\sum_{i=0}^{T}(\sum_{k=i}^{T}\|A^{k-i}-A_{\infty}\|_{2})\|\Delta^{(i)}\|_{F}^{2}\leq s_{A}^{2}\sum_{i=0}^{T}\|\Delta^{(i)}\|_{F}^{2},$
which finishes the proof of this lemma. ∎
Lemma 6.
Given $l\geq 0$ and matrices $M^{(0)},M^{(1)},\ldots,M^{(l)}\in\mathbb{R}^{n\times n}$. For any $k\geq 0$, If $H=\sum_{i=0}^{l}M^{(i)}{\mathbf{g}}^{(k+i)}$ is a linear combination of stochastic gradients, then
$\displaystyle\mathbb{E}_{k}[\|H\|_{F}^{2}]\leq n\sigma^{2}\sum_{i=0}^{l}\|M^{(i)}\|_{2}^{2}+\mathbb{E}_{k}[\|\sum_{i=0}^{l}M^{(i)}\nabla f({\mathbf{x}}^{(k+i)})\|_{F}^{2}].$
Proof.
This is a direct corollary from the unbiasedness and linearly independent gradients in Assumption 3. ∎
Appendix B Proofs of Main Theorem
B.1 Proof of Lemma 3
For simplicity, we define $\Delta_{y}^{(k)}:=(I-B_{\infty}){\mathbf{y}}^{(k)}$ and $\overline{\nabla f}^{(k)}:=n^{-1}\mathds{1}^{\top}\nabla f({\mathbf{x}}^{(k)})$. According to Assumption 3, we have
$\mathbb{E}_{k}[\bar{g}^{(k+i)}]=\mathbb{E}_{k}[\mathbb{E}_{k+i}[\bar{g}^{(k+i)}]]=\mathbb{E}_{k}[\overline{\nabla f}^{(k+i)}],\quad\forall i\geq 0.$
Note that
$\pi_{A}^{\top}\sum_{i=0}^{m-1}{\mathbf{y}}^{(k+i)}=c\sum_{i=0}^{m-1}\bar{g}^{(k+i)}+\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)},$
the descent lemma on $\hat{x}^{(k+m)}=\hat{x}^{(k)}-\alpha\pi_{A}^{\top}\sum_{i=k}^{k+m-1}{\mathbf{y}}^{(i)}$ can be written as follows:
$\displaystyle f(\hat{x}^{(k+m)})-f(\hat{x}^{(k)})$ $\displaystyle\leq-c\alpha\left\langle\sum_{i=0}^{m-1}\bar{g}^{(k+i)},\nabla f(\hat{x}^{(k)})\right\rangle-\alpha\left\langle\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)},\nabla f(\hat{x}^{(k)})\right\rangle$
$\displaystyle\quad+c^{2}\alpha^{2}L\|\sum_{i=0}^{m-1}\bar{g}^{(k+i)}\|^{2}+\alpha^{2}L\|\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}\|^{2}.$
Take expectation conditioned on ${\mathcal{F}}_{k}$ and we obtain that
$\displaystyle\mathbb{E}_{k}[f(\hat{x}^{(k+m)})]-f(\hat{x}^{(k)})\leq-c\alpha\sum_{i=0}^{m-1}\mathbb{E}_{k}\left[\left\langle\overline{\nabla f}^{(k+i)},\nabla f(\hat{x}^{(k)})\right\rangle\right]$ (33)
$\displaystyle\quad-\alpha\mathbb{E}_{k}\left[\left\langle\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)},\nabla f(\hat{x}^{(k)})\right\rangle\right]+c^{2}\alpha^{2}L\mathbb{E}_{k}[\|\sum_{i=0}^{m-1}\bar{g}^{(k+i)}\|^{2}]$
$\displaystyle\quad+\alpha^{2}L\mathbb{E}_{k}[\|\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}\|^{2}].$
The first inner product term in (33) can be bounded by:
$\displaystyle-\left\langle\overline{\nabla f}^{(k+i)},\nabla f(\hat{x}^{(k)})\right\rangle$ $\displaystyle=-\frac{1}{2}\|\overline{\nabla f}^{(k+i)}\|^{2}-\frac{1}{2}\|\nabla f(\hat{x}^{(k)})\|^{2}+\frac{1}{2}\|\overline{\nabla f}^{(k+i)}-\nabla f(\hat{x}^{(k)})\|^{2}$
$\displaystyle\leq-\frac{1}{2}\|\overline{\nabla f}^{(k+i)}\|^{2}-\frac{1}{2}\|\nabla f(\hat{x}^{(k)})\|^{2}+\frac{L^{2}}{2}\|{\mathbf{x}}^{(k+i)}-\hat{{\mathbf{x}}}^{(k)}\|_{F}^{2},$
where the last inequality comes from Cauchy-Schwarz inequality and $L$-smoothness. The first quartic term in (33) can be bounded by:
$\mathbb{E}_{k}[\|\sum_{i=0}^{m-1}\bar{g}^{(k+i)}\|^{2}]\leq n^{-1}m\sigma^{2}+\mathbb{E}_{k}[\|\sum_{i=0}^{m-1}\overline{\nabla f}^{(k+i)}\|^{2}],$
where we use the fact that $\bar{g}^{(k+i)}$ are conditionally unbiased estimates and
$\mathbb{E}_{k}\left[\left\langle\bar{g}^{(k+i)}-\overline{\nabla f}^{(k+i)},\bar{g}^{(k+j)}-\overline{\nabla f}^{(k+j)}\right\rangle\right]=0.$
Plug the estimation for the inner product into (33), we obtain that
$\displaystyle\frac{\alpha cm}{2}\mathbb{E}_{k}[\|\nabla f(\hat{x}^{(k)})\|^{2}]+\frac{\alpha c}{2}\sum_{i=0}^{m-1}\mathbb{E}_{k}[\|\overline{\nabla f}^{(k+i)}\|^{2}]\leq\mathbb{E}_{k}[f(\hat{x}^{(k)})-f(\hat{x}^{(k+m)})]$ (34)
$\displaystyle+\frac{c\alpha L^{2}}{2}\sum_{i=0}^{m-1}\mathbb{E}_{k}[\|{\mathbf{x}}^{(k+i)}-\hat{{\mathbf{x}}}^{(k)}\|_{F}^{2}]+\frac{\alpha^{2}c^{2}mL}{n}\sigma^{2}+c^{2}\alpha^{2}L\mathbb{E}_{k}[\|\sum_{i=0}^{m-1}\overline{\nabla f}^{(k+i)}\|^{2}]$
$\displaystyle-\alpha\mathbb{E}\left[\left\langle\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)},\nabla f(\hat{x}^{(k)})\right\rangle\right]+\alpha^{2}L\mathbb{E}_{k}[\|\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}\|^{2}].$
Next, we try to simplify the last quadratic term in (34) . Notice that
$\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}=\left(\sum_{j=0}^{m-1}(B^{j}-B_{\infty})\right)\Delta_{y}^{(k)}+\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})({\mathbf{g}}^{(k+j)}-{\mathbf{g}}^{(k)}),$
the quadratic term can be bounded as:
$\displaystyle\mathbb{E}_{k}[\|\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}\|^{2}]\leq\mathbb{E}_{k}[\|\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}\|_{F}^{2}]$ (35)
$\displaystyle\leq 2\|\sum_{j=0}^{m-1}(B^{j}-B_{\infty})\|_{2}^{2}\mathbb{E}_{k}[\|\Delta_{y}^{(k)}\|_{F}^{2}]+2\mathbb{E}_{k}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})({\mathbf{g}}^{(k+j)}-{\mathbf{g}}^{(k)})\|_{F}^{2}].$
The first term here can be estimated using the fact that $\|\sum_{j=0}^{m-1}(B^{j}-B_{\infty})\|_{2}^{2}\leq s_{B}^{2}$. For the second term in (35), according to Assumption 3, we have
$\displaystyle\quad\mathbb{E}_{k}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})({\mathbf{g}}^{(k+j)}-{\mathbf{g}}^{(k)})\|_{F}^{2}]$ (36)
$\displaystyle=\sum_{j=1}^{m-1}\mathbb{E}_{k}[\|(B^{m-j-1}-B_{\infty})({\mathbf{g}}^{(k+j)}-\nabla f({\mathbf{x}}^{(k+j)}))\|_{F}^{2}]$
$\displaystyle\quad+\mathbb{E}_{k}[\|(\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty}))({\mathbf{g}}^{(k)}-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
$\displaystyle\quad+\mathbb{E}_{k}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
$\displaystyle\leq n\sigma^{2}\sum_{j=1}^{m-1}\|B^{m-j-1}-B_{\infty}\|_{2}^{2}+n\sigma^{2}\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})\|_{2}^{2}$
$\displaystyle\quad+\mathbb{E}_{k}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
$\displaystyle\leq 2ns_{B}^{2}\sigma^{2}+\mathbb{E}_{k}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
Similarly, using the fact that $\hat{x}^{(k)}$ is known under filtration ${\mathcal{F}}_{k}$, we have
$\displaystyle\quad-\alpha\mathbb{E}_{k}\left[\left\langle\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)},\nabla f(\hat{x}^{(k)})\right\rangle\right]$ (37)
$\displaystyle=-\alpha\left\langle\mathbb{E}_{k}[\pi_{A}^{\top}\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))],\nabla f(\hat{x}^{(k)})\right\rangle$
$\displaystyle\quad+\alpha\left\langle\pi_{A}^{\top}\left(\sum_{i=0}^{m-1}(B^{m-i-1}-B_{\infty})\right)\Delta_{y}^{(k)},\nabla f(\hat{x}^{(k)})\right\rangle$
$\displaystyle\overset{\text{Cauchy}}{\leq}\frac{2\alpha}{cm}\mathbb{E}_{k}[\|\pi_{A}^{\top}\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|^{2}]$ (38)
$\displaystyle\quad+\frac{\alpha cm}{8}\|\nabla f(\hat{x}^{(k)})\|^{2}+\frac{2\alpha}{cm}\|\pi_{A}^{\top}\left(\sum_{i=0}^{m-1}(B^{m-i-1}-B_{\infty})\right)\Delta_{y}^{(k)}\|_{F}^{2}+\frac{\alpha cm}{8}\|\nabla f(\hat{x}^{(k)})\|^{2}$
$\displaystyle\leq\frac{2\alpha}{cm}\mathbb{E}_{k}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]+\frac{\alpha cm}{4}\|\nabla f(\hat{x}^{(k)})\|^{2}$
$\displaystyle\quad+\frac{2\alpha s_{B}^{2}}{cm}\|\Delta_{y}^{(k)}\|_{F}^{2},$
where the last inequality uses $\|\pi_{A}\|\leq 1$ and $\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})\|_{2}\leq s_{B}$. Plug (35) and (37) into (34), we obtain that
$\displaystyle\quad\frac{\alpha cm}{4}\mathbb{E}_{k}[\|\nabla f(\hat{x}^{(k)})\|^{2}]+\frac{\alpha c}{2}\sum_{i=0}^{m-1}\mathbb{E}_{k}[\|\overline{\nabla f}^{(k+i)}\|^{2}]-c^{2}\alpha^{2}L\mathbb{E}_{k}[\|\sum_{i=0}^{m-1}\overline{\nabla f}^{(k+i)}\|^{2}]$ (39)
$\displaystyle\leq\mathbb{E}_{k}[f(\hat{x}^{(k)})-f(\hat{x}^{(k+m)})]+\frac{c\alpha L^{2}}{2}\sum_{i=0}^{m-1}\mathbb{E}[\|{\mathbf{x}}^{(k+i)}-\hat{{\mathbf{x}}}^{(k)}\|_{F}^{2}]$
$\displaystyle+\frac{2\alpha^{2}c^{2}mL}{n}\sigma^{2}+2\alpha^{2}Lns_{B}^{2}\sigma^{2}+(2\alpha^{2}Ls_{B}^{2}+\frac{2\alpha s_{B}^{2}}{cm})\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]$
$\displaystyle+(2\alpha^{2}L+\frac{2\alpha}{cm})\mathbb{E}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
For $c^{2}\alpha^{2}L\mathbb{E}[\|\sum_{i=0}^{m-1}\overline{\nabla f}^{(k+i)}\|^{2}]$, we have the following estimate:
$\displaystyle\quad\mathbb{E}[\|\sum_{i=0}^{m-1}\overline{\nabla f}^{(k+i)}\|^{2}]\leq m\sum_{i=0}^{m-1}\mathbb{E}[\|\overline{\nabla f}^{(k+i)}\|^{2}].$ (40)
Take (40) into (39), and suppose $\alpha\leq\frac{1}{4cmL}$, we obtain that
$\displaystyle\quad\frac{\alpha cm}{4}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]+\frac{\alpha c}{4}\sum_{i=0}^{m-1}\mathbb{E}[\|\overline{\nabla f}^{(k+i)}\|^{2}]$ (41)
$\displaystyle\leq\mathbb{E}[f(\hat{x}^{(k)})-f(\hat{x}^{(k+m)})]+\frac{c\alpha L^{2}}{2}\sum_{i=0}^{m-1}\mathbb{E}[\|{\mathbf{x}}^{(k+i)}-\hat{{\mathbf{x}}}^{(k)}\|_{F}^{2}]$
$\displaystyle+\frac{2\alpha^{2}c^{2}mL}{n}\sigma^{2}+2\alpha^{2}Lns_{B}^{2}\sigma^{2}+\frac{3\alpha s_{B}^{2}}{cm}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]$
$\displaystyle+\frac{3\alpha}{cm}\mathbb{E}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}].$
We can further apply Cauchy’s inequality on the last term in (41), i.e.,
$\displaystyle\quad\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}$
$\displaystyle\leq\sum_{j=1}^{m-1}\|B^{m-j-1}-B_{\infty}\|_{2}^{2}\sum_{j=1}^{m-1}\|\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)})\|_{F}^{2}$
$\displaystyle\leq s_{B}\sum_{j=1}^{m-1}\|\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)})\|_{F}^{2}\leq s_{B}L^{2}\sum_{j=1}^{m-1}\|{\mathbf{x}}^{(k+j)}-{\mathbf{x}}^{(k)}\|_{F}^{2}.$
Finally, when $m\geq 6c^{-2}s_{B}$, we have $\frac{3\alpha s_{B}L^{2}}{cm}\leq\frac{c\alpha L^{2}}{2}$, and (41) can be simplified to
$\displaystyle\quad\frac{\alpha cm}{4}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]+\frac{\alpha c}{4}\sum_{i=0}^{m-1}\mathbb{E}[\|\overline{\nabla f}^{(k+i)}\|^{2}]$
$\displaystyle\leq\mathbb{E}[f(\hat{x}^{(k)})-f(\hat{x}^{(k+m)})]+\frac{c\alpha L^{2}}{2}\sum_{i=0}^{m-1}\mathbb{E}[\|{\mathbf{x}}^{(k+i)}-\hat{{\mathbf{x}}}^{(k)}\|_{F}^{2}]$
$\displaystyle+\frac{2\alpha^{2}c^{2}mL}{n}\sigma^{2}+2\alpha^{2}Lns_{B}^{2}\sigma^{2}+\frac{3\alpha s_{B}^{2}}{cm}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]$
$\displaystyle+\frac{c\alpha L^{2}}{2}\sum_{j=1}^{k}\mathbb{E}_{k}[\|{\mathbf{x}}^{(k+j)}-{\mathbf{x}}^{(k)}\|_{F}^{2}],$
By dividing $\frac{\alpha cm}{4}$ on both sides of this inequality, we finish the proof.
B.2 Proof of Lemma 4
We first provide the complete definition of the constant $C_{m}$ used in the statement. We define $C_{m}$ as the maximum of all lower bounds imposed on $m$ to ensure the validity of subsequent inequalities and simplifications:
$\displaystyle C_{m}=\max\Big\{$ $\displaystyle n^{-1}s_{A}^{2},20c^{-2}ns_{A}s_{A^{m}},\;c^{-2}n^{2}s_{B}^{2},30c^{-2}s_{B}^{2}s_{B^{m}}$
$\displaystyle c^{-2}n^{2}(8s_{A}^{2}s_{B}+2ns_{B}+2s_{A}s_{B}+\sqrt{n}s_{B}),20c^{-2}n^{3}s_{A}s_{A^{m}}s_{B}s_{B^{m}}\Big\}.$ (42)
We define $\delta_{B}^{(0)}=I-B_{\infty}$, $\delta_{B}^{(j)}=B^{j}-B^{j-1},j\geq 1$ and the following constants:
$\displaystyle N_{7}$ $\displaystyle:=\sum_{t=0}^{\infty}\|\sum_{j=0}^{t}A^{t-j}\delta_{B}^{(j)}\|_{2}^{2}\leq 8s_{A}^{2}s_{B}+2n\|\pi_{A}\|^{2}s_{B}\leq 8s_{A}^{2}s_{B}+2ns_{B},$
$\displaystyle N_{8}$ $\displaystyle:=\sum_{t=0}^{\infty}\|\sum_{j=0}^{t}A^{t-j}\delta_{B}^{(j)}\|_{2}\leq 2s_{A}s_{B}+\sqrt{n}\|\pi_{A}\|s_{B}\leq 2s_{A}s_{B}+\sqrt{n}s_{B}.$
For all arguments used in this section, we suppose Proposition 2 and Assumption 34 hold. We first declare that, when $\alpha\leq\frac{1}{6s_{B}L}$, $m\geq\max\{c^{-2}n^{2}(N_{7}+N_{8}^{2}),n^{-1}s_{A}^{2},n^{-1}s_{A},c^{-2}n^{2}s_{A}^{2}\}$, the following inequalities hold $\forall k,m\geq 1$.
$\displaystyle\Delta_{1}^{(k)}$ $\displaystyle\leq 40mn\|\Delta_{x}^{(k)}\|_{F}^{2}+40\alpha^{2}ms_{B}\|\Delta_{y}^{(k)}\|_{F}^{2}+20c^{2}\alpha^{2}m^{2}n^{-1}\sigma^{2}$ (43)
$\displaystyle\quad+20c^{2}\alpha^{2}m^{2}\sum_{j=0}^{m-1}\mathbb{E}_{k}[\|\overline{\nabla f}^{(k+j)}\|^{2}],$
$\displaystyle\Delta_{2}^{(k)}$ $\displaystyle\leq 40mn\|\Delta_{x}^{(k)}\|_{F}^{2}+40\alpha^{2}ms_{B}\|\Delta_{y}^{(k)}\|_{F}^{2}+20c^{2}\alpha^{2}m^{2}n^{-1}\sigma^{2}$ (44)
$\displaystyle\quad+20c^{2}\alpha^{2}m^{2}\sum_{j=0}^{m-1}\mathbb{E}_{k}[\|\overline{\nabla f}^{(k+j)}\|^{2}],$
Note that
$\displaystyle\quad{\mathbf{x}}^{(k+i)}-\hat{{\mathbf{x}}}^{(k)}-(A^{i}-A_{\infty})\Delta_{x}^{(k)}=-\alpha\sum_{j=0}^{i-1}A^{i-j-1}\Delta_{y}^{(k+j)}-\alpha\sum_{j=0}^{i-1}A^{i-j-1}B_{\infty}{\mathbf{y}}^{(k+j)}$
$\displaystyle=-\alpha\sum_{j=1}^{i-1}A^{i-j-1}(B^{j}-B_{\infty})\Delta_{y}^{(k)}-\alpha\sum_{l=1}^{i-1}\sum_{j=0}^{i-l-1}A^{i-j-l-1}\delta_{B}^{(j)}({\mathbf{g}}^{(k+l)}-{\mathbf{g}}^{(k)})$
$\displaystyle\quad\quad-\alpha\sum_{j=0}^{i-1}(A^{i-j-1}-A_{\infty})n\pi_{B}\bar{g}^{(k+j)}-\alpha c\sum_{j=0}^{i-1}\bar{g}^{(k+j)}.$
Therefore, we can apply Cauchy’s inequality and sum it up from $i=1$ to $m-1$:
$\displaystyle\quad\sum_{i=1}^{m-1}\mathbb{E}_{k}[\|{\mathbf{x}}^{(k+i)}-\hat{{\mathbf{x}}}^{(k)}\|_{F}^{2}]$
$\displaystyle\leq 5\sum_{i=1}^{m-1}\|A^{i}-A_{\infty}\|_{2}^{2}\|\Delta_{x}^{(k)}\|_{F}^{2}+5\alpha^{2}\sum_{i=1}^{m-1}\|\sum_{j=1}^{i-1}A^{i-j-1}(B^{j}-B_{\infty})\|_{2}^{2}\|\Delta_{y}^{(k)}\|_{F}^{2}$
$\displaystyle\quad+5\alpha^{2}\sum_{i=1}^{m-1}(\sum_{l=1}^{i-1}\|\sum_{j=0}^{i-l-1}A^{i-j-l-1}\delta_{B}^{(j)}\|_{2}^{2}+\|\sum_{l=1}^{i-1}\sum_{j=0}^{i-l-1}A^{i-j-l-1}\delta_{B}^{(j)}\|_{2}^{2})n\sigma^{2}$
$\displaystyle\quad+5\alpha^{2}\sum_{i=1}^{m-1}\mathbb{E}_{k}[\|\sum_{j=l}^{i-1}A^{i-j-1}\delta_{B}^{(j-l)}(\nabla f({\mathbf{x}}^{(k+l)})-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
$\displaystyle\quad+5\alpha^{2}n^{2}s_{A}^{2}\sum_{i=1}^{m-1}\|\overline{\nabla f}^{(k+j)}\|+5\alpha^{2}n\sum_{i=1}^{m-1}\sum_{j=0}^{i-1}\|A^{i-j-1}-A_{\infty}\|_{2}^{2}$
$\displaystyle\quad+5c^{2}\alpha^{2}m^{2}\sum_{j=0}^{m-1}\|\overline{\nabla f}^{(k+j)}\|^{2}+5c^{2}\alpha^{2}m^{2}n^{-1}\sigma^{2}$
$\displaystyle\leq 5s_{A}\|\Delta_{x}^{(k)}\|_{F}^{2}+10\alpha^{2}(s_{A}^{2}s_{B}+mns_{B})\|\Delta_{y}^{(k)}\|_{F}^{2}+5\alpha^{2}m(N_{7}+N_{8}^{2}+c^{2}mn^{-2})n\sigma^{2}$
$\displaystyle\quad+5\alpha^{2}N_{8}^{2}L^{2}\Delta_{2}^{(k)}+5\alpha^{2}(n^{2}s_{A}^{2}+c^{2}m^{2})\|\overline{\nabla f}^{(k+j)}\|^{2}.$
where in the first inequality, the $\sigma^{2}$ terms are separated using Lemma 6, and the second inequality comes from rolling sum. For the decomposition of ${\mathbf{x}}^{(k+i)}-{\mathbf{x}}^{(k)}$, the only difference is that we replace $(A^{i}-A_{\infty})\Delta_{x}^{(k)}$ with $(A^{i}-I)\Delta_{x}^{(k)}$. This gives
$\displaystyle\quad\Delta_{2}^{(k)}\leq$ $\displaystyle 5\sum_{i=1}^{m}\|A^{i}-I\|_{2}^{2}\|\Delta_{x}^{(k)}\|_{F}^{2}+10\alpha^{2}(s_{A}^{2}+mn)s_{B}\|\Delta_{y}^{(k)}\|_{F}^{2}$
$\displaystyle+5\alpha^{2}m(N_{7}+N_{8}^{2}+c^{2}mn^{-2})n\sigma^{2}+5\alpha^{2}N_{8}^{2}L^{2}\Delta_{2}^{(k)}$
$\displaystyle+5\alpha^{2}(n^{2}s_{A}^{2}+c^{2}m^{2})\|\overline{\nabla f}^{(k+j)}\|^{2}.$
Easy to prove that $\sum_{i=1}^{m}\|A^{i}-I\|_{2}^{2}\leq 2s_{A}+2mn$. When $\alpha\leq\frac{1}{5N_{8}L}$, we have
$\displaystyle\Delta_{2}^{(k)}$ $\displaystyle\leq 20(s_{A}+mn)\|\Delta_{x}^{(k)}\|_{F}^{2}+20\alpha^{2}(s_{A}^{2}+mn)s_{B}\|\Delta_{y}^{(k)}\|_{F}^{2}$ (45)
$\displaystyle\quad+10\alpha^{2}m(N_{7}+N_{8}^{2}+c^{2}mn^{-2})n\sigma^{2}+10\alpha^{2}(n^{2}s_{A}^{2}+c^{2}m^{2})\|\overline{\nabla f}^{(k+j)}\|^{2}.$
Plug (45) into the estimate for $\Delta_{1}^{(k)}$, it is not difficult to verify that $\Delta_{1}^{(k)}$ could use the same upper bound as (45). Next, we provide some inequalities for the consensus error $\Delta_{x}^{(k)}$ and $\Delta_{y}^{(k)}$, $k\in[m^{*}]$. For $k\in[m^{*}]$, we have
$\displaystyle\Delta_{y}^{(k+m)}=(B^{m}-B_{\infty})\Delta_{y}^{(k)}+\sum_{i=0}^{m-1}(B^{m-1-i}-B_{\infty})({\mathbf{g}}^{(k+i+1)}-{\mathbf{g}}^{(k+i)})$ (46)
We define $\Delta_{3}^{(k)}=\sum_{i=0}^{m-1}(B^{m-1-i}-B_{\infty})({\mathbf{g}}^{(k+i+1)}-{\mathbf{g}}^{(k+i)})$, which can be written as
$\displaystyle\Delta_{3}^{(k)}$ $\displaystyle=(I-B_{\infty}){\mathbf{g}}^{(k+m)}+(B-I){\mathbf{g}}^{(k+m-1}+(B^{2}-B){\mathbf{g}}^{(k+m-2)}$
$\displaystyle\quad+\ldots+(B^{m-1}-B^{m-2}){\mathbf{g}}^{(k+1)}-(B^{m-1}-B_{\infty}){\mathbf{g}}^{(k)},$
according to Lemma 6, we have $\mathbb{E}_{k}[\|\Delta_{3}^{(k)}\|_{F}^{2}]\leq n\sigma^{2}q_{B}+h_{m,k}$, where
$\displaystyle h_{m,k}$ $\displaystyle:=\sum_{j=1}^{m}\delta_{B}^{(m-j)}\nabla f({\mathbf{x}}^{(k+j)})-(B^{m-1}-B_{\infty})\nabla f({\mathbf{x}}^{(k)}),$
$\displaystyle q_{B}$ $\displaystyle=\sum_{j=1}^{m}\|\delta_{B}^{(m-j)}\|_{2}^{2}+\|B^{m-1}-B_{\infty}\|_{2}^{2}\leq 2s_{B}.$
For $h_{m,k}$, we can rearrange it as
$\displaystyle h_{m,k}$ $\displaystyle=\sum_{j=1}^{m}\delta_{B}^{(m-j)}(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)})).$
By applying the Jensen’s inequality, we have
$\displaystyle\mathbb{E}_{k}[\|h_{m,k}\|_{F}^{2}]$ $\displaystyle\leq\sum_{t=0}^{m}\|\delta_{B}^{(t)}\|_{2}^{2}\sum_{j=1}^{m}\mathbb{E}_{k}[\|\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)})\|_{F}^{2}]$
$\displaystyle\leq 2s_{B}L^{2}\sum_{j=1}^{m}\mathbb{E}_{k}[\|{\mathbf{x}}^{(k+j)}-{\mathbf{x}}^{(k)}\|_{F}^{2}]=2s_{B}L^{2}\Delta_{2}^{(k)}.$
Combining the estimates above, we reach that
$\mathbb{E}_{k}[\|\Delta_{3}^{(k)}\|_{F}^{2}]\leq 2ns_{B}\sigma^{2}+2s_{B}L^{2}\Delta_{2}^{(k)}.$
To proceed on, we iterate (46) and obtain that
$\displaystyle\Delta_{y}^{(k+m)}=(B^{m+k}-B_{\infty}){\mathbf{g}}^{(0)}+\sum_{j=0}^{k/m}(B^{k-jm}-B_{\infty})\Delta_{3}^{(jm)}.$
Using the rolling sum lemma, we have
$\displaystyle\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{y}^{(k+m)}\|_{F}^{2}]\leq 2\sum_{k\in[m^{*}]}\mathbb{E}[\|(B^{k}-B_{\infty}){\mathbf{g}}^{(0)}\|_{F}^{2}]+2s_{B^{m}}^{2}\sum_{k\in[m^{*}]}\mathbb{E}[\mathbb{E}_{k}[\|\Delta_{3}^{(k)}\|_{F}^{2}]]$
$\displaystyle\quad\leq s_{B^{m}}(n\sigma^{2}+2nL\Delta)+4ns_{B}s_{B^{m}}^{2}K\sigma^{2}+4s_{B}s_{B^{m}}^{2}L^{2}\sum_{k\in[m^{*}]}\mathbb{E}[\Delta_{2}^{(k)}].$
Note that
$\mathbb{E}[\|\Delta_{y}^{(0)}\|_{F}^{2}]=\mathbb{E}[\|(I-B_{\infty}){\mathbf{g}}^{(0)}\|_{F}^{2}]\leq 2(n\sigma^{2}+2nL\Delta),$
we combine coefficients and reach that
$\displaystyle\sum_{k\in[m^{*}]}\|\Delta_{y}^{(k)}\|_{F}^{2}$ $\displaystyle\leq 7s_{B}s_{B^{m}}^{2}Kn\sigma^{2}+6s_{B^{m}}nL\Delta+4s_{B}s_{B^{m}}^{2}L^{2}\sum_{k\in[m^{*}]}\Delta_{2}^{(k)}.$
Replace the $\Delta_{2}$ term with the estimate in (44), we reach that
$\displaystyle(1-160\alpha^{2}mL^{2}s_{B}^{2}s_{B^{m}}^{2})\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]$
$\displaystyle\leq$ $\displaystyle(7s_{B}s_{B^{m}}^{2}+80c^{2}m^{2}n^{-1}\alpha^{2}L^{2}s_{B}s_{B^{m}}^{2})Kn\sigma^{2}+6s_{B^{m}}nL\Delta$
$\displaystyle+160s_{B}s_{B^{m}}^{2}L^{2}mn\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{x}^{(k)}\|_{F}^{2}]+80c^{2}m^{2}s_{B}s_{B^{m}}^{2}\alpha^{2}L^{2}\sum_{t=0}^{T-1}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|^{2}]$
When $\alpha\leq\frac{1}{10cm\sqrt{s_{B}s_{B^{m}}^{2}}L}$ and $m\geq 10c^{-2}s_{B}s_{B^{m}}$, we have $1-160\alpha^{2}mL^{2}s_{B}^{2}s_{B^{m}}^{2}\geq 0.5$. and the coefficient of $Kn\sigma^{2}$ becomes smaller than $8s_{B}s_{B^{m}}^{2}$. Thus,
$\displaystyle\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]\leq 16s_{B}s_{B^{m}}^{2}Kn\sigma^{2}+12s_{B^{m}}nL\Delta$ (47)
$\displaystyle\quad+320s_{B}s_{B^{m}}^{2}L^{2}mn\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{x}^{(k)}\|_{F}^{2}]+160c^{2}m^{2}s_{B}s_{B^{m}}^{2}\alpha^{2}L^{2}\sum_{t=0}^{T-1}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|^{2}].$
By taking this back to (44), we have
$\displaystyle\sum_{k\in[m^{*}]}\mathbb{E}[\Delta_{2}^{(k)}]$ $\displaystyle\leq 40mn(1+320\alpha^{2}L^{2}ms_{B}^{2}s_{B^{m}}^{2})\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{x}^{(k)}\|_{F}^{2}]$ (48)
$\displaystyle\quad+480\alpha^{2}ms_{B^{m}}nL\Delta+(640ms_{B}^{2}s_{B^{m}}^{2}+20c^{2}m^{2}n^{-2})\alpha^{2}Kn\sigma^{2}$
$\displaystyle\quad+20c^{2}\alpha^{2}m^{2}(1+160\alpha^{2}L^{2}ms_{B}^{2}s_{B^{m}}^{2})\sum_{t=0}^{T-1}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|^{2}].$
$\displaystyle\leq 50mn\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{x}^{(k)}\|_{F}^{2}]+480\alpha^{2}ms_{B^{m}}nL\Delta$
$\displaystyle\quad+21c^{2}m^{2}n^{-1}\alpha^{2}K\sigma^{2}+21c^{2}\alpha^{2}m^{2}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|^{2}].$
where the second inequality uses $\alpha\leq\frac{1}{10cmL}$ and $m\geq 4c^{-2}s_{B}^{2}s_{B^{m}}^{2}n^{2}$.
For $k\in[m^{*}]$, we have
$\displaystyle\Delta_{x}^{(k+m)}=(A^{m}-A_{\infty})\Delta_{x}^{(k)}-\alpha\sum_{i=0}^{m-1}(A^{m-i}-A_{\infty}){\mathbf{y}}^{(k+i)}.$ (49)
For $k\in[m^{*}]$, define $\Delta_{4}^{(k)}:=\sum_{i=0}^{m-1}(A^{m-i}-A_{\infty}){\mathbf{y}}^{(k+i)}$, then we can iterate (49):
$\Delta_{x}^{(k+m)}=-\alpha\sum_{j=0}^{k/m}(A^{k-jm}-A_{\infty})\Delta_{4}^{(jm)},\quad k\in[m^{*}].$
Using the rolling sum lemma, we know that for $k\in[m^{*}]$,
$\quad\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{x}^{(k+m)}\|_{F}^{2}]\leq\alpha^{2}s_{A^{m}}^{2}\sum_{k\in[m^{*}]}\mathbb{E}[\mathbb{E}_{k}[\|\Delta_{4}^{(k)}\|_{F}^{2}]].$
To further analyze $\mathbb{E}_{k}[\Delta_{4}^{(k)}]$, $k\in[m^{*}]$, we expand it as
$\displaystyle\Delta_{4}^{(k)}$ $\displaystyle=\sum_{i=0}^{m-1}(A^{m-i}-A_{\infty})\Delta_{y}^{(k+i)}+n\sum_{i=0}^{m-1}(A^{m-i}-A_{\infty})\pi_{B}\bar{g}^{(k+i)}$
$\displaystyle=\sum_{i=0}^{m-1}(A^{m-i}-A_{\infty})(B^{i}-B_{\infty})\Delta_{y}^{(k)}+n\sum_{i=0}^{m-1}(A^{m-i}-A_{\infty})\pi_{B}\bar{g}^{(k+i)}$
$\displaystyle\quad+\sum_{j=i}^{m-1}(A^{m-i}-A_{\infty})\delta_{B}^{(j-i)}({\mathbf{g}}^{(k+j)}-{\mathbf{g}}^{(k)}).$
Note that
$\displaystyle\|\sum_{i=0}^{m-1}(A^{m-i}-A_{\infty})(B^{i}-B_{\infty})\|_{2}^{2}$ $\displaystyle\leq s_{A}s_{B},\,\,\sum_{i=0}^{m-1}\|(A^{m-i}-A_{\infty})\delta_{B}^{(j-i)}\|_{2}\leq 2s_{A}s_{B},$
$\displaystyle\sum_{i=0}^{m-1}\|(A^{m-i}-A_{\infty})\delta_{B}^{(j-i)}\|_{2}^{2}$ $\displaystyle\leq 2s_{A}s_{B},\,\,\|\sum_{i=0}^{m-1}(A^{m-i}-A_{\infty})\delta_{B}^{(j-i)}\|_{2}^{2}\leq 2s_{A}s_{B}.$
Therefore, according to Lemma 6, we can separate noiseless and noisy terms in $\sum_{i=0}^{m-1}\pi_{B}\bar{g}^{(k+i)}$ and $\sum_{j=i}^{m-1}(A^{m-i}-A_{\infty})(B^{j-i}-B^{j-i+1})({\mathbf{g}}^{(k+j)}-{\mathbf{g}}^{(k)})$, then apply Jensen’s inequality. This gives
$\displaystyle\mathbb{E}_{k}[\|\Delta_{4}^{(k)}\|_{F}^{2}]\leq 3s_{A}s_{B}\|\Delta_{y}^{(k)}\|_{F}^{2}+3n^{2}s_{A}\sum_{i=0}^{m-1}\mathbb{E}_{k}[\|\overline{\nabla f}^{(k+i)}\|_{F}^{2}]+3ns_{A}\sigma^{2}$
$\displaystyle\quad+6s_{A}s_{B}\sum_{j=i}^{m-1}\mathbb{E}_{k}[\|\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)})\|_{F}^{2}]+12s_{A}s_{B}n\sigma^{2}$
$\displaystyle\leq 3s_{A}s_{B}\|\Delta_{y}^{(k)}\|_{F}^{2}+3n^{2}s_{A}\sum_{i=0}^{m-1}\mathbb{E}_{k}[\|\overline{\nabla f}^{(k+i)}\|_{F}^{2}]+6s_{A}s_{B}L^{2}\Delta_{2}^{(k)}+15s_{A}s_{B}n\sigma^{2}.$
Plug in our estimate on $\Delta_{4}^{(k)}$, we have
$\displaystyle\quad\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{x}^{(k+m)}\|_{F}^{2}]\leq\alpha^{2}s_{A^{m}}^{2}\sum_{k\in[m^{*}]}\mathbb{E}[\mathbb{E}_{k}[\|\Delta_{4}^{(k)}\|_{F}^{2}]]$
$\displaystyle\leq 15\alpha^{2}ns_{A}s_{B}s_{A^{m}}^{2}K\sigma^{2}+3\alpha^{2}n^{2}s_{A}s_{A^{m}}^{2}\sum_{t=0}^{T-1}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|_{F}^{2}]$
$\displaystyle\quad+6\alpha^{2}L^{2}s_{A}s_{B}s_{A^{m}}^{2}\sum_{k\in[m^{*}]}\mathbb{E}[\Delta_{2}^{(k)}]+3\alpha^{2}s_{A}s_{B}s_{A^{m}}^{2}\sum_{k\in[m^{*}]}\mathbb{E}\|\Delta_{y}^{(k)}\|_{F}^{2}$
Note that we start from consensual $x$, so $\Delta_{x}^{(0)}=0$. Plug in (48) and (47) we have
$\displaystyle\quad(1-480\alpha^{2}mL^{2}s_{A}s_{B}s_{A^{m}}^{2}n-960\alpha^{2}L^{2}ms_{A}s_{B}^{2}s_{A^{m}}^{2}s_{B^{m}}^{2}n)\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{x}^{(k)}\|_{F}^{2}]$
$\displaystyle\leq(15+126c^{2}\alpha^{2}L^{2}m^{2}n^{-2}+48s_{B}s_{B^{m}}^{2})\alpha^{2}s_{A}s_{B}s_{A^{m}}^{2}nK\sigma^{2}$
$\displaystyle\quad+(3+(126+480s_{B}s_{B^{m}}^{2})c^{2}\alpha^{2}L^{2}m^{2})\alpha^{2}n^{2}s_{A}s_{A^{m}}^{2}\sum_{t=0}^{T-1}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|_{F}^{2}]$
$\displaystyle\quad+(36+2880\alpha^{2}L^{2})s_{A}s_{B}s_{A^{m}}^{2}s_{B^{m}}\alpha^{2}nL\Delta$
Using $\alpha\leq\min\{\frac{1}{10\sqrt{s_{B}s_{B^{m}}^{2}}cmL},\frac{1}{6L}\}$ and $m\geq 20c^{-2}ns_{A}s_{B}s_{A^{m}}^{2}$, the coefficients can be simplified to
$\displaystyle\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{x}^{(k)}\|_{F}^{2}]\leq 64\alpha^{2}s_{A}s_{B}^{2}s_{B^{m}}^{2}s_{A^{m}}^{2}nK\sigma^{2}$ (50)
$\displaystyle\quad+7\alpha^{2}n^{2}s_{A}s_{A^{m}}^{2}\sum_{t=0}^{T-1}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|_{F}^{2}]+416\alpha^{2}s_{A}s_{B}s_{A^{m}}^{2}s_{B^{m}}nL\Delta$
Taking (50) back to (47), we obtain that
$\displaystyle\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]\leq(1+1280\alpha^{2}L^{2}mn^{2}s_{A}s_{B}^{2}s_{B^{m}}^{2}s_{A^{m}}^{2})16s_{B}s_{B^{m}}^{2}Kn\sigma^{2}$
$\displaystyle\quad+(12+133120\alpha^{2}L^{2}s_{A}s_{B}^{2}s_{B^{m}}s_{A^{m}}^{2})s_{B^{m}}nL\Delta$
$\displaystyle\quad+160(c^{2}m^{2}s_{B}s_{B^{m}}^{2}+140\alpha^{2}L^{2}n^{2}s_{A}s_{A^{m}}^{2}s_{B}s_{B^{m}}^{2})\alpha^{2}L^{2}\sum_{t=0}^{T-1}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|_{F}^{2}].$
When $\alpha\leq\frac{1}{10cm\sqrt{s_{B}s_{B^{m}}^{2}}L}$ and $m\geq 10c^{-2}n^{2}s_{A}s_{A^{m}}^{2}s_{B}$, we have
$\displaystyle\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]$ $\displaystyle\leq 32s_{B}s_{B^{m}}^{2}Kn\sigma^{2}+26s_{B^{m}}nL\Delta$ (51)
$\displaystyle\quad+170c^{2}m^{2}\alpha^{2}L^{2}s_{B}s_{B^{m}}^{2}\sum_{t=0}^{T-1}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|_{F}^{2}].$
Plug (50) into (48), we obtain that
$\displaystyle\sum_{k\in[m^{*}]}\mathbb{E}[\Delta_{2}^{(k)}]\leq(21c^{2}m^{2}n^{-2}+3200mns_{A}s_{B}^{2}s_{A^{m}}^{2}s_{B^{m}}^{2})\alpha^{2}nK\sigma^{2}$ (52)
$\displaystyle\quad+21280s_{A}s_{B}s_{A^{m}}^{2}s_{B^{m}}\alpha^{2}mn^{2}L\Delta+(21c^{2}m+350n^{3}s_{A}s_{A^{m}}^{2})\alpha^{2}m\sum_{t=0}^{T-1}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|_{F}^{2}].$
$\displaystyle\leq 181c^{2}m^{2}n^{-1}\alpha^{2}K\sigma^{2}+21280s_{A}s_{B}s_{A^{m}}^{2}s_{B^{m}}\alpha^{2}mn^{2}L\Delta+38\alpha^{2}c^{2}m^{2}\sum_{t=0}^{T-1}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|_{F}^{2}],$
where the second inequality uses $m\geq 20c^{-2}n^{3}s_{A}s_{A^{m}}^{2}s_{B}^{2}s_{B^{m}}^{2}$. Now we are ready to estimate the summation of error terms and prove Lemma 4. Remember that $\Delta_{1}^{(k)}$ and $\Delta_{2}^{(k)}$ share the same upper bound, we utilize (52), (51) and obtain that
$\displaystyle\quad\frac{2L^{2}}{m}\sum_{k\in[m^{*}]}\mathbb{E}[\Delta_{1}^{(k)}+\Delta_{2}^{(k)}]+\frac{12s_{B}^{2}}{c^{2}m^{2}}\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]$
$\displaystyle\leq C_{\sigma,1}K\sigma^{2}+\frac{1}{m}C_{f,1}\sum_{t=0}^{T-1}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|_{F}^{2}]+C_{\Delta,1}\frac{L\Delta}{m^{2}},$
where
$\displaystyle C_{\sigma,1}$ $\displaystyle=724c^{2}mn^{-1}\alpha^{2}L^{2}+\frac{384s_{B}^{3}s_{B^{m}}^{2}}{m^{2}c^{2}}$
$\displaystyle C_{f,1}$ $\displaystyle=152c^{2}\alpha^{2}L^{2}m^{2}+2040m\alpha^{2}L^{2}s_{B}^{3}s_{B^{m}}^{2}$
$\displaystyle C_{\Delta,1}$ $\displaystyle=85120s_{A}s_{B}s_{A^{m}}^{2}s_{B^{m}}\alpha^{2}L^{2}m^{2}n^{2}+\frac{312s_{B}^{2}s_{B^{m}}n}{c^{2}}.$
When $\alpha\leq\frac{1}{10cm\sqrt{s_{B}s_{B^{m}}^{2}}L}$ and $m\geq c^{-2}30s_{B}^{1.5}$, we have
$C_{\sigma,1}\leq\frac{73c\alpha L}{n}+\frac{384s_{B}^{3}s_{B^{m}}^{2}}{m^{2}c^{2}},\quad C_{f,1}\leq 1,\quad C_{\Delta,1}\leq\frac{10s_{A}s_{A^{m}}^{2}n^{2}+312s_{B}^{2}s_{B^{m}}n}{c^{2}}.$ (53)
This finishes the proof of Lemma 4. Finally, we leave a brief discussion on $s_{B^{m}}$. Recall that
$s_{B^{m}}=\max\{\sum_{i=0}^{\infty}\|B^{im}-B_{\infty}\|_{2},\sum_{i=0}^{\infty}\|B^{im}-B_{\infty}\|_{2}^{2}\},$
when $m$ increases, both $\sum_{i=1}^{\infty}\|B^{im}-B_{\infty}\|_{2}$ and $\sum_{i=1}^{\infty}\|B^{im}-B_{\infty}\|_{2}^{2}$ converge to zero exponentially fast, and $\|I-B_{\infty}\|_{2}\geq 1$. Thus, $s_{B^{m}}$ converges to $\|I-B_{\infty}\|_{2}^{2}$ exponentially fast as $m$ grows.
B.3 Proof of Theorem 1
Sum up Lemma 3 for $k\in[m^{*}]$, then plug in Lemma 4, we obtain that
$\displaystyle\quad\frac{\alpha cm}{4}\sum_{k\in[m^{*}]}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]+\frac{\alpha c}{4}(1-C_{f,1})\sum_{t=0}^{T-1}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|^{2}]$ (54)
$\displaystyle\leq f(x^{(0)})-f(\hat{x}^{T})+\frac{\alpha cm}{4}\left(\frac{8\alpha cL}{n}+\frac{8\alpha Lns_{B}^{2}}{cm}+C_{\sigma,1}\right)K\sigma^{2}+C_{\Delta,1}\frac{c\alpha L\Delta}{4m}.$
Divide both sides by $\frac{\alpha cmK}{4}$ and utilize $C_{f,1}\leq 1$, $f(x^{(0)})-f(\hat{x}^{T})\leq\Delta$, $m\geq c^{-2}n^{2}s_{B}^{2}$ and $C_{\sigma,1}\leq\frac{73c\alpha L}{n}+\frac{384s_{B}^{3}s_{B^{m}}^{2}}{m^{2}c^{2}}$, $T=mK$, we have
$\displaystyle\quad\frac{1}{K}\sum_{k\in[m^{*}]}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]\leq\frac{\Delta}{c\alpha T}+\frac{100c\alpha L}{n}\sigma^{2}+\frac{384s_{B}^{3}s_{B^{m}}^{2}}{m^{2}c^{2}}\sigma^{2}+C_{\Delta,1}\frac{L\Delta}{mT}.$ (55)
Inequality (55) holds when $\alpha\leq\min\{\frac{1}{10cm\sqrt{s_{B}s_{B^{m}}^{2}}},\frac{1}{6L}\}$ and $m$ is larger than some matrix-related constants.
To obtain a standard linear speedup result, we set
$\displaystyle m$ $\displaystyle=\lceil\max\{4c^{-1}s_{B}^{1.5}s_{B^{m}}\left(\frac{nT\sigma^{2}}{L\Delta}\right)^{0.25},C_{m}\}\rceil,$
$\displaystyle\alpha$ $\displaystyle=\min\left\{\frac{\sqrt{n\Delta}\sigma}{10c\sqrt{LT}},\frac{1}{10cm\sqrt{s_{B}s_{B^{m}}^{2}}L},\frac{1}{6L}\right\},$
giving that
$\displaystyle\quad\min_{t\in\{0,1,\ldots,T-1\}}\mathbb{E}[\|\nabla f(\hat{x}^{(t)}\|^{2}]\leq\frac{1}{K}\sum_{k\in[m^{*}]}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]$
$\displaystyle\leq\frac{44\sigma\sqrt{L\Delta}}{\sqrt{nT}}+\frac{L\Delta(m^{-1}C_{\Delta,1}+10m\sqrt{s_{B}s_{B^{m}}^{2}}+c^{-1})}{T}$
$\displaystyle=\frac{44\sigma\sqrt{L\Delta}}{\sqrt{nT}}+{\mathcal{O}}\left(\frac{L\Delta}{T}\right)^{\frac{3}{4}}+{\mathcal{O}}\left(\frac{L\Delta}{T}\right).$
Appendix C Experiment Details
In this section, we supplement experiment setup details.
C.1 Weight Matrix Design
We use the following procedure to generate a row/column-stochastic weight matrix. For a fixed number of nodes $n$, we begin by initializing a 0-1 adjacency matrix $W$ as the identity matrix $I$ to account for self-loops on each node. Subsequently, based on the specific network topology definition, we set $W_{i,j}=1$ if there is a directed edge from node $i$ to node $j$. After constructing this initial 0-1 matrix $W$, we then randomly assign integer weights from $1$ to $9$ to all entries where $W_{i,j}=1$. Finally, we perform either row normalization or column normalization on this weighted matrix to obtain a row-stochastic matrix $A$ or a column-stochastic matrix $B$, respectively.
C.2 Synthetic data
For experiments conducted on the synthetic dataset, we minimize the objective:
$\displaystyle f(x)=\frac{1}{n}\sum_{i=1}^{n}f_{i}(x),\quad x\in\mathbb{R}^{d},$ (56)
where
$f_{i}(x)=\underbrace{\frac{1}{L_{i}}\sum_{l=1}^{L_{i}}\ln\left(1+\exp(-y_{i,l}h_{i,l}^{\top}x)\right)}_{\text{Logistic Loss}}+\underbrace{\rho\sum_{j=1}^{d}\frac{[x]_{j}^{2}}{1+[x]_{j}^{2}}}_{\text{Non-convex Regularization}},\quad L_{i}=L_{\text{Local}}/n.$
Data Synthesis Process. We generate the synthetic dataset through the following three-step procedure. (1) Global Data Generation: We first generate optimal parameters $x_{\text{opt}}\sim\mathcal{N}(0,I_{d})$ and create a global feature matrix $H\in\mathbb{R}^{L_{\text{total}}\times d}$ with entries $h_{l,j}\sim\mathcal{N}(0,1)$. Labels $Y\in\mathbb{R}^{L_{\text{total}}}$ with $y_{l}\in\{-1,+1\}$ are computed via randomized thresholding:
$y_{l}=\begin{cases}1&\text{if }1/z_{l}>1+\exp(-h_{l}^{\top}x_{\text{opt}})\\ -1&\text{otherwise}\end{cases},\quad z_{l}\sim\mathcal{U}(0,1)$
(2) Data Distribution: We partition $H$ and $Y$ equally across $n$ nodes using $H^{(i)}=H[iM:(i+1)M,:]$ and $Y^{(i)}=Y[iM:(i+1)M]$ where $M=L_{\text{total}}/n$, requiring $L_{\text{total}}$ to be divisible by $n$. (3) Local Initial Points: Each node $i$ starts from $x_{i}^{\star}=x_{\text{opt}}+\varepsilon_{i}$ where $\varepsilon_{i}\sim\mathcal{N}(0,\sigma_{h}^{2}I_{d})$ with $\sigma_{h}=10$.
Gradient Computation. At each iteration, node $i$ independently computes its stochastic gradient by randomly sampling a minibatch of size $B$ from its local dataset of size $L_{i}=L_{\text{total}}/n$:
$\nabla f_{i,B}(x)=\underbrace{-\frac{1}{B}\sum_{b=1}^{B}\frac{y_{i,b}h_{i,b}}{1+\exp(y_{i,b}h_{i,b}^{\top}x)}}_{\text{Logistic Loss Gradient}}+\underbrace{\rho\sum_{j=1}^{d}\frac{2[x]_{j}}{(1+[x]_{j}^{2})^{2}}}_{\text{Non-convex Regularization}}$
where the minibatch $\{h_{i,b},y_{i,b}\}_{b=1}^{B}$ is drawn uniformly without replacement from the $L_{i}$ local samples. The gradient at each node $i$ is computed on its local parameter $x_{i}\in\mathbb{R}^{d}$.
Implementation Details. We use global dataset size $L_{\text{total}}=204800$ with batch size $B=200$, dimension $d=10$, and regularization parameter $\rho=0.01$. Node configurations include $n\in\{4,6,8,9,12,16,18,24,25\}$ with least common multiple (LCM) of $3,600$. Note that $L_{\text{total}}=1,440,000=3,600\cdot 200\cdot 2=\text{LCM}\cdot B\cdot 2$.
Evaluation Metric. We track the gradient norm $\|n^{-1}\mathds{1}^{\top}\nabla f({\mathbf{x}}^{(k)})\|$ where
$\displaystyle n^{-1}\mathds{1}^{\top}\nabla f({\mathbf{x}}^{(k)})=\frac{1}{n}\sum_{i=1}^{n}\nabla f_{i,M}(\bm{x}_{i}^{(k)}),\quad M=L_{\text{Total}}/n.$
Results represent averages over 20 independent trials with fixed random seed 42 for reproducibility.
C.3 Neural Network Experiments
We evaluate our Push-Pull method on two benchmark tasks: MNIST classification using three-layer fully connected networks and CIFAR-10 classification using ResNet-18.
Data Distribution. To simulate heterogeneous data distributions across nodes, we employ Dirichlet sampling. For each node $k\in\{1,\dots,n\}$, we sample a class proportion vector $\mathbf{q}^{k}=(q_{k,0},q_{k,1},\dots,q_{k,9})\sim\text{Dirichlet}(\alpha\cdot\mathbf{1}^{10})$, where $\mathbf{1}^{10}$ is a 10-dimensional vector of ones. Each $q_{k,c}$ represents the proportion of samples from class $c$ assigned to node $k$, satisfying $q_{k,c}\geq 0$ and $\sum_{c=0}^{9}q_{k,c}=1$. The target number of samples of class $c$ for node $k$ is $N_{k,c}^{\text{target}}=q_{k,c}\cdot M/n$, resulting in approximately $60,000/n$ samples per node for MNIST and $50,000/n$ for CIFAR-10. We set $\alpha=0.9$ to create heterogeneous distributions, as illustrated in Figure 5 for 16 nodes.
Evaluation Metrics. We measure convergence using the normalized gradient norm computed after each batch. Specifically, we average the gradients from all $n$ nodes and normalize by the square root of the total parameter count:
$\displaystyle\left\|\frac{1}{n}\sum_{i=1}^{n}\mathbf{g}_{i,B_{j}}\right\|_{2}\Big/\sqrt{P},$ (57)
where $\mathbf{g}_{i,B_{j}}$ is the gradient computed on the $j$-th batch of node $i$, and $P$ is the number of parameters. For MNIST with three-layer FCNs, $P=1,863,690$; for CIFAR-10 with ResNet-18, $P=11,173,962$.
Experimental Setup. We test our method on various network topologies including directed exponential and grid graphs, as well as undirected random graphs. Training uses learning rate $lr=0.005$ and batch size $B=128$ for MNIST, $B=128$ for CIFAR-10, consistent across all topologies and node configurations. Performance is compared against single-node baseline training ($n=1$). Figures 2 and 3 demonstrate that Push-Pull achieves linear speedup on both synthetic data and real datasets across directed and undirected topologies. Figure 4 shows the validation accuracy comparison for CIFAR-10.
Figure 5: Comparison of uniform (left) and heterogeneous (right) data distributions across $16$ nodes ($\alpha=0.9$).